### v1

# Rectified Flows for Fast Multiscale Fluid Flow Modeling

Abstract
The statistical modeling of fluid flows is very challenging due to their multiscale dynamics and extreme sensitivity to initial conditions. While recently proposed conditional diffusion models achieve high fidelity, they typically require hundreds of stochastic sampling steps at inference. We introduce a rectified flow framework that learns a time-dependent velocity field, transporting input to output distributions along nearly straight trajectories. By casting sampling as solving an ordinary differential equation (ODE) along this straighter flow field, our method makes each integration step much more effective, using as few as eight steps versus (more than) 128 steps in standard score-based diffusion, without sacrificing predictive fidelity. Experiments on challenging multiscale flow benchmarks show that rectified flows recover the same posterior distributions as diffusion models, preserve fine-scale features that MSE-trained baselines miss, and deliver high-resolution samples in a fraction of inference time.
1 Introduction
Partial differential equations (PDEs) [4] are the backbone of modeling physical phenomena, from atmospheric dynamics to aerodynamics and magnetohydrodynamics. Currently, solutions of PDEs are simulated with high-fidelity numerical solvers such as finite difference, finite element, and spectral methods [24]. These are often prohibitively expensive for many–query tasks [25] such as uncertainty quantification, optimization, and real-time prediction. In particular, these solvers encounter a significant challenge for classes of non-linear PDEs, whose solutions exhibit extreme sensitivity to initial and boundary data, leading to chaotic behavior and complex multi-scale features. In such settings, one seeks, not a deterministic trajectory, but rather the statistical solution of the PDE, i.e., the push-forward of the underlying probability measure on inputs (initial and boundary conditions etc) by the PDE solution operator [5]. Currently, these statistical solutions are sampled with ensemble Monte Carlo method by repeatedly calling high-fidelity PDE solvers [13]. Given the computational cost of the PDE solver, ensemble methods become infeasible as ensemble sizes grow.
Machine learning and AI based algorithms are increasingly being explored for the data-driven simulation of PDEs [20]. In particular, neural operators [12, 2], which directly seek to learn the PDE solution operator from data, are already widely used in a PDE context [20]. Examples of neural operators have been proposed in [14, 15, 18, 26, 11, 22, 31, 10, 16, 1] and references therein.
However, these neural operators have been found to be unsuitable for simulating PDE with chaotic, multiscale solutions such as governing equations of fluid dynamics. For instance, recently in [21], it was shown both theoretically and empirically that neural operators, trained to minimize mean absolute (square) errors, will always collapse to the ensemble mean when used for simulating statistics of PDEs with chaotic multiscale solutions.
Given this context, generative AI models—especially score-based and diffusion models—have recently shown remarkable success at learning complex, high-dimensional distributions by progressive denoising score matching  [29]. In the context of chaotic PDEs, GenCFD [21] introduced a conditional diffusion framework that preserves spectral content across fine scales, mitigating the “spectral collapse” of naive MSE-trained surrogates, see also [23, 7, 8, 9]. However, sampling a score-based Diffusion model requires simulating a reverse-time stochastic differential equation (SDE) over 100+ discretization steps, each invoking a large ML model on a moderate-to-high resolution spatial grid, resulting in high runtimes that undermine real-time or resource-constrained applications.
To overcome this limitation of diffusion models, we adapt rectified flows [17]: a deterministic ODE-based analogue of diffusion that “straightens” transport paths in probability space. Rather than small random increments, rectified flows solve an ODE whose velocity field is trained to match the diffusion model’s instantaneous score, but whose trajectories remain nearly linear. This structure permits large time integration, reducing the number of solver steps by an order of magnitude or more, while provably recovering the same target distribution.
Hence, the main contributions of our paper are,
  • Rectified flow framework: We propose a deterministic transport approach for conditional PDE data, leveraging “straightened” trajectories that can be integrated in far fewer steps than typical diffusion-based methods. This reduced order model perspective challenges current wide-spread practices which rely on overtly-complicated interpolations between base and target measures.
  • Empirical gains in multi-scale problems: Across diverse 2D benchmarks for the Incompressible Navier–Stokes and Compressible Euler equations governing fluid dynamics, our method generates samples up to 22x faster than diffusion baselines (e.g. GenCFD), while retaining multi-scale accuracy. Furthermore, we observe a significant gain of efficiency in training, where rectified-flows converge reliably faster across all experimental testbeds, while preserving their strong performance.
2 Approach
Modeling multi-scale flows and their statistical solutions. We focus on time-dependent PDEs of the general form
[…]
subject to boundary conditions $\mathcal{B}(u)=u_{b}$ on $\partial D\times(0,T)$ and initial data $u(x,0)=\bar{u}(x)$. Here, the solution $u:D\mapsto{\mathbb{R}}^{m}$ is evolved in time, with respect to the differential operator $\mathcal{L}$. Concrete examples of (2.1) are the incompressible Navier-Stokes and Compressible Euler equations.
A recurring theme in non-linear, multi-scale fluid flows is extreme sensitivity to initial or boundary data, which can lead to chaotic or near-chaotic behavior. Rather than fix a single $u_{0}$, one may consider an initial measure $\mu_{0}$ to represent uncertain or variable initial states, and then study the statistical solution $\mu_{t}=\mathcal{S}^{t}_{\#}\mu_{0}$ pushed forward by the PDE solution operator $\mathcal{S}^{t}$. In strongly nonlinear regimes, $\mu_{t}$ spreads significantly in function space, creating complex, multi-scale distributions of possible flow fields. Accurately approximating this evolving distribution is essential for uncertainty quantification and design under rough conditions, but doing so with high-fidelity PDE solvers can be prohibitively expensive when large ensembles are required.
High cost of computing statistical solutions. In practice, characterizing $\mu_{t}$ directly often requires performing many forward simulations of the PDE solver, each initialized with a slightly perturbed $u_{0}$. As each simulation might require high-resolution discretizations (e.g. millions of degrees of freedom) and multiple time steps, the overall cost for constructing a sizable ensemble rapidly becomes prohibitive. Even with modern high-performance computing platforms, running hundreds or thousands of high-fidelity statistical simulations can be infeasible. Hence, there is a pressing need for more efficient surrogate models that can sample physically realistic final states without enumerating every realization of the forward PDE solve. Succinctly put, we aim to provide a solution to the following problem
How can we efficiently capture statistical quantities of interest for the pushforward measure $S^{t}_{\#}\mu_{0}$ for highly sensitive solution operators $S^{t}$?
Rectified flows for faster sampling. To mitigate these inefficiencies, we adopt the framework of rectified flows [17]: a deterministic alternative to conditional diffusion designed to reduce the number of required integration steps. In essence, rather than generating samples via small random “diffusion-reversal” increments, rectified flows construct an ordinary differential equation (ODE) with nearly straight trajectories in function space. This straighter transport map can be integrated more aggressively in time, enabling sampling with far fewer solver steps. Importantly, rectified flows retain the desirable property of matching the same final distributions that standard diffusion models learn, but at a fraction of the runtime cost.
Figure 2.1: Overview of ReFlow. (a): ReFlow closely matches high-order statistical moments (in the Wasserstein $W_{1}$-sense) of fluid flows by learning a neural surrogate to approximate the push-forward of the distribution $p(u_{\tau|\tau=0}|u_{i})$, while minimizing inference-time computational costs. (b) Macro-architecture overview: A UViT based network predicts the velocity field $v_{\theta}$, conditioned on Gaussian noise, initial data $u_{i}$ and boundary conditions, and diffusion time $\tau$. Noise and initial / boundary conditions are concatenated as inputs. (c): Results for the cylindrical shear flow dataset. (d): The KDE plots monitor the evolution of a down-projected version of the push-forward measure $p_{\tau}:=X_{\theta}(\tau)_{\#}p(u_{\tau=0}|u_{i})$, where $X_{\theta}$ is the flow induced by the learned vector field $v_{\theta}$. Results for the Richtmyer Meshkov dataset showing flow matching between noise ($p_{\tau=0}$) and target ($p_{\tau=1}$) distributions. ReFlow solves an ODE, learning a constant velocity field that enables straight-line inference paths (visualized with principal component analysis and kernel density estimation at different diffusion times).
3 Related Work
3.1 Diffusion Models in Fluid Dynamics
Conditional diffusion surrogates. GenCFD [21] trains a conditional score-based diffusion model that starts from an initial or low-resolution flow field and iteratively refines it through hundreds of denoising steps, thereby capturing multi-scale turbulent structure. Similarly, the method of [23] conditions a diffusion model on a coarse fluid solution (generated by a learned operator) and “super-resolves” the fine-scale flow. Given these similarities, we refer to such diffusion-based solvers collectively as conditional diffusion PDE solvers.
Rectified flow alternative. Each sample typically needs hundreds of network calls, limiting large-scale or real-time use. Our method replaces the stochastic SDE with a deterministic ODE whose nearly straight trajectories require far fewer steps, achieving similar statistical fidelity at much lower inference cost (see Sections 4.3 and 5). Crucially, we find that rectified flows are actually more efficient from the standpoint of training-time required to reach the same accuracy level with the conditional diffusion approaches; this is also confirmed by earlier work [3] exploring the scalability of rectified flows with transformer-based backbones.
4 Rectified Flows: An ODE-Based Alternative
4.1 Mathematical Formulation
Overview.
Rectified flows provide a deterministic way to transform one probability distribution into another, offering an alternative to stochastic diffusion-based methods. Instead of traversing curved, noise-driven paths, rectified flows attempt to transport mass along nearly “straight” trajectories in state space, which can dramatically reduce sampling costs.
Throughout, our random variables are velocity fields $u:\Omega\!\to\!\mathbb{R}^{3}$ living in a separable, divergence-free space $\mathcal{U}=L^{2}_{\!\mathrm{div}=0}(\Omega;\mathbb{R}^{3})$ (or $H^{s}_{\!\mathrm{div}=0}$, $s>\frac{3}{2}$). All theory below extends verbatim from $\mathbb{R}^{d}$ because $\mathcal{U}$ is Polish; in practice we work with a Galerkin discretisation $\mathbb{R}^{d_{N}}\subset\mathcal{U}$, so every integral is evaluated by Monte-Carlo on its finite grid surrogate.
In all that follows, the diffusion time $\tau\in[0,1]$ is a purely scale-ordering variable: $\tau=0$ denotes the coarsest, lowest–frequency representation of a velocity field, while $\tau=1$ restores full spatial resolution; it is not the physical PDE time but merely an index that governs the progressive refinement of scales along the rectified flow.
Setup via measure coupling. Let $\mu_{0},\mu_{1}\in\mathcal{P}(\mathcal{U})$ be the initial and target field distributions. A coupling $\gamma\in\mathcal{P}(\mathcal{U}\times\mathcal{U})$ satisfies $(\pi_{0})_{\#}\gamma=\mu_{0}$ and $(\pi_{1})_{\#}\gamma=\mu_{1}$, where $\pi_{i}$ is the canonical projection on the $i$’th component of $(u_{0},u_{1})\in\mathcal{U}\times\mathcal{U}$ For $\tau\in[0,1]$ define the linear interpolation
$T_{\tau}(u_{0},u_{1})=(1-\tau)u_{0}+\tau\,u_{1},\qquad d\rho_{\tau}=(T_{\tau})_{\#}\gamma.$
Intuitively, $d\rho_{t}$ is the measure obtained by partially “mixing” points from $u_{0}$ to $u_{1}$.
Definition 1 (Barycentric Velocity Field).
For each $\tau\in[0,1]$ and $d\rho_{\tau}$-almost every $u$, define
$v(u,\tau)\;:=\;\int_{T_{\tau}^{-1}(u)}\bigl{(}u_{1}-u_{0}\bigr{)}\,\gamma^{\tau}_{u}(du_{0},du_{1}),$ (4.1)
where $\gamma^{\tau}_{u}$ is the disintegration of $\gamma$ over the fiber $\;T_{\tau}^{-1}(u)=\{(u_{0},u_{1}):\,(1-\tau)u_{0}+\tau u_{1}=u\}.$
By disintegrating $\gamma$ along these linear interpolations, the optimal velocity field defined above “sweeps” from coarse to fine structure.
Straightness measure. Rectified flows seek a velocity $v$ that does not “bend” unnecessarily over time. One can quantify this via a straightness measure $S(v)$, such as
[…]
When $S(v)$ is small, the velocity field $v(\cdot,\tau)$ remains nearly constant in $\tau$, ensuring trajectories are close to straight lines in state space.
ODE formulation. Sampling from the rectified flow amounts to solving the ODE,
$\frac{du_{\tau}}{d\tau}\;=\;v\bigl{(}u_{\tau},\,\tau\bigr{)},\quad u_{0}\sim\mu_{0}.$ (4.2)
As $\tau$ goes from 00 to $1$, $u_{\tau}$ travels along straighter interpolants between $\mu_{0}$ and $\mu_{1}$. In practice, we learn a $v$ that approximates the barycentric velocity field while promoting small $S(v)$.
Practical learning objective. Given samples $\{(u_{0},u_{1})\}$ from $\gamma$, one can minimize
[…]
which precisely encourages small $S(v)$. This ensures $v$ approximates the conditional expectation $u_{1}-u_{0}$ at each partial interpolation $U_{\tau}$, but remains as “straight” as possible over $\tau$. Furthermore, by integrating over these noise scales, the model sees high frequencies more consistently than an MSE-based approach.
Conditional expectation viewpoint. Equivalently, if $(U_{0},U_{1})\sim\gamma$ and $U_{\tau}=(1-\tau)U_{0}+\tau U_{1}$, then the true minimizer takes the form
$v(u,\tau)\;=\;\mathbb{E}[\,U_{1}-U_{0}\,\mid\,U_{\tau}=u\,].$
by the $L^{2}$ optimality property of conditional expectations. Thus $v(u,\tau)$ is the average instantaneous displacement from $u_{0}$ to $u_{1}$, conditioned on arriving at $u$ at time $\tau$.
4.2 Implementation Details for PDEs
Learning velocity fields in PDE settings. When rectified flows are used to model PDE solutions, we assume access to paired initial/final states $(u_{i},u_{f})$ where $u_{f}=\mathcal{S}(u_{i})$ represents the evolved initial condition $u_{i}$, at some final physical time (e.g. $t=1$). We train a neural network to predict the velocity $v_{\theta}$ that displaces $u_{i}$ towards $u_{f}$, with the additional mention that $v_{\theta}$ shall be conditioned on $u_{i}$ at all diffusion times.
Algorithm 1 Rectified Flow Training (PDE Context)
1:Dataset $\{(u_{i},u_{f})\}$ with $u_{f}=\mathcal{S}(u_{0})$; velocity model $v_{\theta}$; noise schedule $\sigma(\tau)$
2:for each iteration do
3:     Sample a pair $(u_{i},u_{f})$
4:     Sample $\tau\sim\mathcal{U}(0,1)$ and noise $\xi\sim\mathcal{N}(0,I)$
5:     Form $u_{\tau}=\tau\cdot u_{1}+\sigma(1-\tau)\,\xi$
6:     Compute loss:
$\mathcal{L}(\theta)\;=\;\bigl{\|}\bigl{(}u_{1}-u_{0}\bigr{)}\;-\;v_{\theta}\bigl{(}u_{\tau},\,u_{0},\,\tau\bigr{)}\bigr{\|}^{2}$
7:     Update $\theta$ via gradient descent
8:end for
Overview of the model. Our network learns a velocity (or denoising) map
$v_{\theta}(u_{\tau},u_{0},\,\tau)\;=\;\mathrm{UNet}_{\theta}\bigl{(}\,u_{\tau},u_{0},\;\Gamma(\tau)\bigr{)},$
where $u_{\tau}$ is a low-frequency representation of $u_{f}$, $u_{i}$ is the initial datum, and $\Gamma(\cdot)$ is an MLP-based time embedding:
$\Gamma(\tau)\;=\;\mathrm{MLP}\bigl{(}\mathrm{Sinusoid}(\tau)\bigr{)}.$
Within the U-Net (see SM A.1), we alternate residual blocks with multi-head attention layers, downsampling to a bottleneck and then upsampling via skip connections. A final $1\times 1$ projection recovers either the field increment $\hat{u}\approx u_{1}-u_{0}$ or a noise residual. By injecting the rectification time $\Gamma(t)$ into each block, the model performs a fully deterministic trajectory integration instead of stochastic diffusion.
Sampling procedure. Once trained, we generate new samples by numerically integrating the ODE
$\frac{du_{\tau}}{d\tau}=v_{\theta}\bigl{(}u_{\tau},\,u_{i},\,\tau\bigr{)},\quad u_{\tau}|_{\tau=0}=\xi,\quad\xi\sim\mathcal{N}(0,I).$
Any standard ODE solver (e.g. forward Euler, Runge-Kutta) can be used; rectified flows often tolerate relatively large step sizes due to their straighter velocity field.
Efficiency and accuracy. Because rectified flows explicitly encourage a “straighter” velocity field, accurate solutions often require fewer integration steps than their diffusion-based counterparts. This advantage can be crucial in computational fluid dynamics and other PDE applications where each solver step is expensive. See SM D.3 and F for a detailed comparison of step-count vs. accuracy trade-offs.
4.3 Computational Efficiency Comparison
Figure 4.1: Inference cost versus model size. Rectified flows achieve the desired accuracy with only 44448888 ODE steps, while GenCFD variants need 64646464128128128128 reverse-time SDE steps. We vary U-Net depth from $4$ to $8$ residual blocks, growing the parameter count from $5.5$ M to $10.2$ M.
Diffusion surrogates (e.g. GenCFD) generate samples by integrating a reverse-time SDE whose trajectory bends through latent noise; even with accelerated samplers such as DDPM or DPM-solver the procedure still requires tens of score-network evaluations. Rectified flows, in contrast, follow a single deterministic ODE whose nearly straight path tolerates large or adaptive steps, so inference needs only a handful of evaluations. Figure 4.1 quantifies this gap, showing how the runtime of diffusion models scales almost linearly with the number of steps, whereas rectified flows remain fast even as network size grows.
5 Experiments
5.1 Experimental Setup
We benchmark our method, termed henceforth as ReFlow, against four neural surrogate models. GenCFD and GenCFD $\circ$ FNO are conditional diffusion surrogates following the framework introduced earlier in this section, whereas FNO and UViT serve as one-shot deterministic baselines. Detailed descriptions of all baselines are provided in SM C.
Datasets for Training and Evaluation.
We consider three canonical 2D multi-scale PDE problems (see the Supplementary Material for precise mathematical formulations and solver details):
  • SL2D
    Shear Layer in 2D. This setup models the evolution of an incompressible shear layer, initially concentrated near a horizontal interface, which rolls up into intricate vortical structures over time. We generate approximate solutions by numerically solving the 2D Navier–Stokes system (with periodic boundaries) at a resolution of $512^{2}$, then downsample to $128^{2}$ for training and testing. The data has been generated with the spectral viscosity code [30, 27].
  • CS2D
    Cloud-Shock Interaction in 2D. In this popular compressible-flow test, a shock wave moves through a higher-density “cloud” region, triggering complex wave interactions and small-scale turbulent mixing in the wake. We solve the compressible Euler system at an effective resolution of $256^{2}$ using a GPU-optimized finite-volume ALSVINN code of [19, 6], then also downsample to $128^{2}$.
  • RM2D
    Richtmeyer–Meshkov in 2D. In this case, a shock wave interacts with a density interface separating two fluids, depositing vorticity and triggering vigorous mixing. We numerically solve the compressible Euler system on a $256^{2}$ grid, then downsample to $128^{2}$ to create the training and test data. The data has been generated with a high-resolution finite-volume ALSVINN code of [6]
Training and Test Protocol.
We train both our proposed rectified flow-based GenCFD-class algorithm and all baseline models on initial conditions
$u_{0}\,\sim\,\mu,\quad\text{with}\quad\mu$
being the probability distribution specified in the Supplementary Material (SMB.2). For each $u_{0}$, the PDE solver (e.g. a spectral or finite-volume code) is run forward in time to produce high-fidelity snapshots, which serve as training data.
Training-objective. Crucially, during training, the model is only exposed to $(u_{0},u_{1})$ pairs of data as opposed to the conditional measure $p(u(t)|u_{0})$.
Test-time goal: Dirac initial measures. While the training distribution $\mu$ is spread out, we are especially interested in the scenario where the initial condition is fixed and deterministic. Formally, we set
$\widehat{\mu}_{0}=\delta_{\widehat{u}_{0}},$
which is a Dirac measure concentrated at a single function $\widehat{u}_{0}$. Thus, any stochasticity in the resulting flow arises solely from chaotic or multi-scale evolution over time, rather than from large variations in the initial condition itself.
To create a “ground truth” ensemble for comparison under a fixed $\widehat{u}_{0}$, we employ an ensemble perturbation procedure (detailed in SM  B.2), where $\widehat{u}_{0}$ is slightly perturbed to generate multiple nearby initial states. Each perturbed state is then advanced using a high-resolution PDE solver (e.g. the Azeban code for incompressible flows, or a finite-volume code for compressible problems), yielding a set of flow realizations. This set approximates the conditional measure in question, i.e. $p\!\bigl{(}u(t)\mid\widehat{u}_{0}\bigr{)}$.
Distribution settings. We test two main configurations:
  1. 1.
    Matched distribution. We sample $\widehat{u}_{0}$ from the same family $\mu$ used in training. While $\delta_{\widehat{u}_{0}}$ is technically out-of-distribution in the sense that the training distribution was spread out over many initial states, the base function $\widehat{u}_{0}$ still belongs to the same general class of flows.
  2. 2.
    Distribution shift. We consider Dirac initial data $\delta_{\widehat{u}_{0}}$ where $\widehat{u}_{0}\sim\nu$, for a different distribution $\nu\neq\mu$. This scenario probes the ability of our model and baselines to generalize under shifts in the initial condition distribution. Specific details on $\nu$ are provided in SM B.2.
We will show results in both regimes to assess how well rectified flows can capture the spread of chaotic solutions starting from fixed initial states—and how it compares to baselines.
5.2 Evaluation Metrics
We evaluate the generative performance of each method using three main metrics, with details provided in SM D:
  • Mean Error. We measure the $L^{2}$ difference between the model’s predicted mean flow $\mu$ and the ground-truth mean $\mu_{\mathrm{exact}}$. A low mean error indicates accurate large-scale behavior.
  • Standard Deviation Error. The normalized $L^{2}$ distance between the predicted standard deviation $\sigma$ and the reference $\sigma_{\mathrm{exact}}$ captures whether the model reproduces correct spatial variability.
  • One-Point Wasserstein Distance. We compute an Wasserstein-$1$ distance at each spatial point (and average over the domain) to quantify local distributional agreement between ground truth and model outputs.
Table 5.1: Comparison of models across datasets CS2D, SL2D, and RM2D, in terms of mean relative error ($e_{\mu}$), standard deviation relative error ($e_{\sigma}$), and Wasserstein-1 distance ($W_{1}$). Best-performing model is colored in Blue and second-best performing model in Orange.
Model Metric CS2D SL2D RM2D
$\rho$ $m_{x}$ $m_{y}$ $E$ $u_{x}$ $u_{y}$ $\rho$ $m_{x}$ $m_{y}$ $E$
ReFlow $e_{\mu}$ 0.0477 0.0332 0.041 0.015 0.034 0.189 0.018 0.032 0.020 0.021
$e_{\sigma}$ 0.1005 0.072 0.078 0.0584 0.071 0.077 0.032 0.046 0.046 0.085
$W_{1}$ 0.0091 0.0107 0.0124 0.0143 0.0214 0.0164 0.0033 0.0036 0.0028 0.0017
GenCFD $e_{\mu}$ 0.0830 0.0619 0.127 0.0280 0.039 0.267 0.0049 0.0010 0.0025 0.0025
$e_{\sigma}$ 0.195 0.197 0.244 0.169 0.092 0.072 0.0351 0.0254 0.0435 0.0435
$W_{1}$ 0.0151 0.0184 0.0185 0.0247 0.0276 0.0213 0.0092 0.0016 0.0049 0.0049
UViT $e_{\mu}$ 0.111 0.078 0.127 0.0479 0.233 0.623 0.070 0.070 0.071 0.052
$e_{\sigma}$ 0.415 0.343 0.412 0.296 0.28 0.178 0.425 0.570 0.570 0.617
$W_{1}$ 0.087 0.085 0.158 0.172 0.270 0.256 0.060 0.055 0.060 0.036
FNO $e_{\mu}$ 0.0989 0.0748 0.0930 0.0383 0.058 0.408 0.085 0.070 0.063 0.067
$e_{\sigma}$ 0.2573 0.2779 0.3188 0.2777 0.137 0.110 0.366 0.443 0.436 0.450
$W_{1}$ 0.1455 0.1660 0.1879 0.1884 0.184 0.150 0.096 0.051 0.056 0.062
GenCFD $\boldsymbol{\circ}$ FNO $e_{\mu}$ 0.0638 0.0454 0.0603 0.0204 0.055 0.393 0.040 0.027 0.039 0.041
$e_{\sigma}$ 0.1099 0.0969 0.1245 0.0974 0.131 0.100 0.091 0.075 0.091 0.090
$W_{1}$ 0.0467 0.0534 0.0774 0.0743 0.166 0.135 0.0269 0.0158 0.0293 0.0273
Baselines and Evaluation Settings. To stay within reasonable training times for our infrastructure, we restrict the total number of training iterations up to 120,000 for each model and recorded the best achievable results. Under this budget, only a reduced-capacity GenCFD (with two downsampling stages in its U-Net backbone, versus three in Rectified Flow) can match the accuracy reported in Table 5.1. We also evaluated larger GenCFD variants trained for 500,000 iterations, but observed no meaningful improvement over the results shown here. Meanwhile, at inference time Rectified Flow requires only 8 denoising steps to achieve comparable or superior accuracy, whereas diffusion baselines (including GenCFD) need 128 steps, underscoring RF’s efficiency in both training and sampling. FNO and UViT also require smaller models than ReFlow for optimal performance.
Since FNO, and UViT produce deterministic predictions, we approximate their spread by an ensemble perturbation strategy (SM  B.2), where each baseline is fed slightly perturbed inputs and outputs are aggregated to form a sample set. In contrast, both the original GenCFD (baseline) and our ReFlow are inherently probabilistic, requiring no auxiliary procedure to generate diverse samples.
Figure 5.1: Energy-spectrum evolution for a random cloud–shock initial datum. ReFlow (blue) tracks high-wavenumber energy faster and more accurately than GenCFD (red).
5.3 Results.
[…]
Moreover as visualized in the SM, ReFlow generates samples of very high quality, matching the qualitative features of the ground truth and capturing features such as (interacting) vortices as well as propagating shock waves very accurately. Another aspect where ReFlow really shines in being able to generate the correct point pdfs. Although already indicated in the very low 1-point Wasserstein distance errors from Table 5.1, we present 1-point pdfs in the SM to illustrate this observation.
[…]
Figure 5.2: Density trajectories for Sample 1: ReFlow (top) versus GenCFD (bottom).
6 Conclusion and Future Work
We have demonstrated that rectified flows provide an efficient surrogate for sampling multi-scale PDE solutions: on a suite of challenging incompressible and compressible problems, our method matches the statistical fidelity of conditional diffusion models while requiring an order of magnitude fewer solver steps—yielding up to 22× faster inference in practice. Across matched and shifted initial-condition regimes, rectified flows consistently reproduce energy spectra, vorticity distributions, and one-point Wasserstein distances within sampling error, confirming their robustness and multi-scale accuracy.
Key Takeaways.
On account of its straightening bias, ReFlow requires significantly fewer ODE steps at inference-time (8-10) than its conditional-diffusion-based counterparts (more than 128) to achieve the same level of performance on our statistical metrics – consult SM D for more details. Furthermore, ReFlow is much more training efficient, requiring up to 120,000 iterations in reaching the same level of performance, across all experiments, as our other conditional diffusion baselines (which typically take 500-600,000 iterations), see SM D.2. Our results further demonstrate that straightening effects remain evident even when projecting down to a 2D PCA subspace (see SM D.3).
Limitations and Future Work
Looking forward, we aim to scale rectified-flow models to three-dimensional turbulence and coupled multi-physics scenarios, investigate lightweight architectures for real-time deployment, and integrate our generative sampler into uncertainty-quantification pipelines for design and control applications. Continuous benchmarking on higher-resolution datasets and exploration of transfer learning across flow regimes will further validate the practicality of rectified flows as a generative foundation for computational fluid dynamics.
Supplementary Material for:
Rectified Flows for Fast Multiscale Fluid Flow Modeling
A Architecture of the ReFlow Model12
 A.1 Architectural Components12
 A.2 Training Pipeline Details14
 A.3 Summary and Key Takeaways15
B Datasets15
 B.1 Multi-scale Flow Datasets15
 B.2 Preprocessing and Data Organization16
C Models and Baselines17
 C.1 GenCFD Baseline17
 C.2 GenCFD $\circ$ FNO 17
 C.3 FNO 18
 C.4 UViT 18
D Results18
 D.1 Performance Metrics19
 D.2 Training Efficiency Observations 19
 D.3 Experimental Advantages of Straightness20
E Computational Considerations27
F Additional Visualizations28
Appendix A Architecture of the ReFlow Model
Figure A.1: UNet Backbone Architecture Used in ReFlow. This schematic illustrates the core UNet-based architecture used within the ReFlow framework, structured across three resolution levels. For clarity, the number of blocks per level is set to one in this illustration. In the actual GenCFD and ReFlow configurations, each block is repeated four times per level. The bottleneck shows an asymmetry between the encoder and decoder sides: the block on the encoder side includes a Convolution Block, Positional Embedding, and Attention Block, while the corresponding block on the decoder side omits the Positional Embedding.
A.1 Architectural Components
Motivation. Multi-scale PDE data contain wide-ranging spatial scales, so a multi-resolution encoder–decoder is natural. We incorporate time embeddings to condition each layer on the interpolation fraction $t$.
Architecture Layout. We employ a UNet-based architecture augmented with attentional layers, typically termed as UViT by [28]. In addition to the MLP-based time embedding $\Gamma(\tau)$, the input signal is lifted into a higher-dimensional embedding space and later projected back into physical space through convolutional layers. The data flows from high to low resolution across a down-sampling stack composed of convolution blocks into a bottleneck layer that encodes the fine-scale features of the inputs across its channel dimension and mixes them via a multi-head attention mechanism.
On the encoder side, exclusively at the bottleneck level, each attention block preceded by a simple linear positional embedding. This positional encoding helps preserve spatial locality before global mixing through attention. The decoder side does not apply any positional embedding.
The UNet is conditioned at every level through normalized Feature-wise Linear Modulation (FiLM) techniques embedded within the convolution blocks. The conditioning signal modulates the normalization layers via learned scale and shift parameters. This conditioning mechanism is consistently applied across all encoder and decoder stages.
The output is symmetrically reconstructed via an up-sampling pipeline. Downsampling is performed using standard convolutional layers with stride, while upsampling can be implemented either via transposed convolutions or by applying a standard convolution followed by a non-learnable pixel shuffle operation, which rearranges elements of the tensor spatially according to a fixed upsampling factor.
Importantly, the architecture described here resembles the one presented in [21], ensuring consistency with prior state-of-the-art practices in high-fidelity generative modeling for scientific computing.
A.1.1 Convolution Block
Each UNet level employs dedicated convolution blocks that process features while enabling conditioning on the diffusion time $\tau$. These blocks consist of two convolutional layers interleaved with group normalization and Swish activations. Between the convolutions, a normalized FiLM layer adaptively scales features using the conditioning embeddings. A residual connection combines the processed features with a projected skip connection, ensuring stable gradient flow. Crucially, these convolution blocks form the fundamental building units across all encoder and decoder stages, not just the bottleneck, with consistent application of temporal conditioning through the FiLM mechanism.
A.1.2 Normalized FiLM Layer
Each convolution block uses a normalized FiLM layer and applies a convolution layer interleaved with normalization and an activation, then adds a residual skip:
$\mathbf{z}$ $=\;\mathrm{Conv}\bigl{(}\mathrm{Norm}(\mathbf{x})\bigr{)},\qquad\mathbf{z}\;=\;\sigma(\mathbf{z}+\alpha\,\gamma(t)),$ (A.1)
$\mathbf{x}_{\mathrm{out}}$ $=\;\mathbf{x}\;+\;\mathrm{Conv}\!\bigl{(}\mathbf{z}\bigr{)},$
where $\mathrm{Norm}$ is often RMSNorm, GroupNorm or LayerNorm, $\sigma$ is a nonlinearity (e.g. $\mathrm{SiLU}$), and $\gamma(t)$ represents the time embedding (an MLP on $\mathrm{Sinusoid}(t)$) that modulates scale/shift. This temporal conditioning is crucial for guiding the generative process through rectified diffusion trajectories.
A.1.3 Attention Block
The architecture strategically employs multi-head attention exclusively at the bottleneck layer to balance global interaction modeling with computational efficiency. Each attention block processes normalized features through a spatial-channel reformatting, where features are reshaped by a flattening operation. This preserves the channel structure while collapsing the spatial grid into a single flattened dimension, allowing attention to operate across spatial locations. At certain scales, a multi-head attention (MHA) mechanism captures distant flow interactions. We define
$\mathbf{Q}\;=\;W_{Q}\,\mathbf{h},\quad\mathbf{K}\;=\;W_{K}\,\mathbf{h},\quad\mathbf{V}\;=\;W_{V}\,\mathbf{h},\quad\mathcal{A}(\mathbf{h})\;=\;\mathrm{softmax}\!\Bigl{(}\tfrac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\Bigr{)}\,\mathbf{V},$ (A.2)
where $\mathbf{h}\in\mathbb{R}^{(H\cdot W)\times d}$ denotes the flattened input sequence, with each of the $H\cdot W$ spatial positions represented as a $d$-dimensional feature vector. In practice, attention is split into multiple heads for better representational power, then recombined. This global mixing is key for PDE flows dominated by far-field couplings (e.g. vortex merging). The MHA operator is defined as
$\mathrm{MHA}(\mathbf{h})\;=\;W^{O}\;\mathrm{Concat}\left(\mathcal{A}_{1}(\mathbf{h}),\dots,\mathcal{A}_{n}(\mathbf{h})\right),$ (A.3)
where $n\in\mathbb{N}$ denotes the number of attention heads and $W^{O}\in\mathbb{R}^{C\times C}$ is a learned output projection matrix. The attention block applies MHA to a normalized input, then adds a residual connection to preserve the original signal and stabilize training while enabling global context to influence local features:
$\mathbf{z}$ $=\;\mathrm{MHA}\bigl{(}\mathrm{Norm}(\mathbf{x})\bigr{)},$ (A.4)
$\mathbf{x}_{\mathrm{out}}$ $=\;\mathbf{x}\;+\;\mathbf{z}.$
A.2 Training Pipeline Details
The training pipeline follows the algorithm described in the main text faithfully.
Trainer Workflow.
A high-level Trainer class manages:
  1. 1.
    Batching PDE pairs $(u_{0},u_{1})$. These normalized pixel-wise with statistics extracted per each entire dataset.
  2. 2.
    Sampling $\tau\in[0,1]$ and random noise $\xi$ according to a given noise scheduler $\sigma(\cdot)$.
  3. 3.
    Partial interpolation + noise addition: $\tilde{u}_{\tau}:=\tau\,u_{1}+(1-\tau)\,\xi$ (or the flow-based variant). We have experimented with non-linear interpolation methods, but have not noticed meaningful oscillations in performance.
  4. 4.
    Minimizing the squared deviation between model output and target displacement/noise, optionally with a consistency/EMA update.
  5. 5.
    Throughout training, we monitor validation performance on a separate holdout dataset across all main metrics. We explicitly note that different validation splits have been tested out to ensure there is no data leakage that would bias performance metrics.
Table A.1: Representative hyperparameters for 2D fluid-flow tasks. See main text for justification of these choices.
Parameter Value (Optimal empirical performance) Notes
Channels per level (128, 256, 256) Number of channels per level.
Downsample Ratios (2, 2, 2) Downsampling ratio per level.
Attention Heads 8 Each attention block.
Dropout Rate 0.1 In ConvBlock for generalization.
Noise Schedule uniform, log-uniform, cos-map Helps with multi-scale noise.
ODE Steps (Sampling) 4–8 Often >15×>15$\times$> 15 × fewer vs. diffusion.
Batch Size 16 Tied to GPU capacity.
EMA Decay 0.9999 Teacher model for consistency.
Learning Rate $3\times 10^{-4}$ Cosine decay in code.
In practice we noticed that the most robust denoising schedule was just the uniform one. For details on the cos-map schedule, we direct the reader to its description in [3].
A.3 Summary and Key Takeaways
The RectifiedFlow class, combined with a time-aware UNet2D architecture, builds a deterministic transport model well-suited to multi-scale PDE tasks. By interpolating between noise and data along nearly-straight paths, it reduces sampling overhead compared to conventional diffusion PDE solvers while retaining the flexibility to model complex turbulent phenomena. Further experimental and theoretical results appear in SM Section D. Instead of the multi-step stochastic integration in diffusion-based models, RectifiedFlow fits a deterministic ODE from noise to data. This leverages straighter trajectories in frequency space, enabling fewer network evaluations at inference time.
ODE Integration for Sampling.
At inference:
$\frac{d\tilde{u}_{\tau}}{d\tau}\;=\;v_{\theta}\bigl{(}\tilde{u}_{\tau},\,u_{0},\,\tau\bigr{)},\quad\tau\in[0,1],\quad\tilde{u}_{\tau=0}\,=\,\text{random noise}.$
We solve this with odeint. Few steps ($8\!-\!10$) typically suffice, contrasting with the tens or hundreds required by full diffusion samplers.
Appendix B Datasets
B.1 Multi-scale Flow Datasets
Following the dataset generation procedure of Herde et al. [11], the Richtmyer–Meshkov ensemble is created by imposing a randomized sinusoidal perturbation on a two-fluid interface and driving it with a planar shock via prescribed pressure jumps. For the CloudShock and ShearFlow benchmarks, we follow the GenCFD mesh-generation and solver setup of Molinaro et al. [21], initializing CloudShock with concentric density perturbations and ShearFlow with orthogonal shear jets on adaptive Cartesian grids, and employing the same high-order finite-volume scheme and dissipation settings to ensure consistent numerical fidelity across both datasets.
Overview. We experiment on three paired-field benchmarks $(u_{0},u_{1})$:
  • Richtmyer–Meshkov (RM): 64 000 samples
  • Cloud-Shock (CS): 40 000 samples
  • Shear-Layer (SL): 79 200 samples
All datasets track the $u_{0}\to S^{t}(u_{0})|_{t=1}$ solution mapping from inputs to the evolved outputs at $t=1$. The first two datasets monitor the evolution of density ($\rho$), momentum components $(m_{x},m_{y})$, and energy respectively ($E$). The shear-flow dataset monitors the velocity components of the flow $(u_{x},u_{y})$.
Train/Validation Split. For each dataset, we reserve 80% of the samples for training and 20% for validation.
Macro–Micro Ensemble Evaluation (CS & SL). To probe out-of-distribution generalization under small perturbations, we adopt a two-stage ensemble protocol:
  1. 1.
    Macro-sampling: select $M_{\mathrm{macro}}=10$ distinct base initial conditions.
  2. 2.
    Micro-perturbation: for each base, generate $M_{\mathrm{micro}}=1{,}000$ perturbed copies within a small radius.
  3. 3.
    Metrics: compute all performance measures (mean-field error $e_{\mu}$, standard-deviation error $e_{\sigma}$, average $W_{1}$, spectral agreement) by first averaging over each micro-ensemble, then averaging across the $M_{\mathrm{macro}}$ cases.
In-Distribution Evaluation (RM). For the Richtmyer–Meshkov dataset, we additionally evaluate on a held-out test set drawn from the same distribution (no macro–micro perturbations). This allows us to study how deterministic and probabilistic models scale with training set size under standard, in-distribution conditions.
B.2 Preprocessing and Data Organization
Data Processing
Before training, we compute and store global, channel-wise statistics over the entire training dataset. Concretely, for each physical variable $c$ we calculate
$\mu_{c}$ $=\frac{1}{N_{\rm train}\,H\,W\,[D]}\sum_{i=1}^{N_{\rm train}}\sum_{x,y,[z]}u^{(i)}_{c}\bigl{(}x,y,[z],t_{0}\bigr{)},$
$\sigma_{c}$ $=\sqrt{\frac{1}{N_{\rm train}\,H\,W\,[D]}\sum_{i=1}^{N_{\rm train}}\sum_{x,y,[z]}\Bigl{(}u^{(i)}_{c}\bigl{(}x,y,[z],t_{0}\bigr{)}-\mu_{c}\Bigr{)}^{2}}.$
where $N_{\rm train}$ is the number of training samples, $H\times W\,[\times D]$ the spatial grid size, and $u_{c}(x,y,[z],t_{0})$ the initial field for channel $c$. We average over all spatial locations to yield a single mean $\mu_{c}$ and standard deviation $\sigma_{c}$ per channel. The same statistics are also computed over the final-time fields $u_{c}(t_{f})$, producing $\mu_{c}^{\prime}$ and $\sigma_{c}^{\prime}$ for the output channels.
During both training and evaluation, each input field $u_{c}$ is standardized via
$\tilde{u}_{c}=\frac{\,u_{c}\;-\;\mu_{c}\,}{\sigma_{c}+10^{-16}},\qquad\tilde{v}_{c}=\frac{\,v_{c}\;-\;\mu_{c}^{\prime}\,}{\sigma_{c}^{\prime}+10^{-16}},$
where $u_{c}$ and $v_{c}$ denote the initial and target fields, respectively. We then concatenate $\{\tilde{u}_{c}\}$ and $\{\tilde{v}_{c}\}$ along the channel axis before feeding them to the network. Because we always apply the same pre-computed statistics at test time, no information from the evaluation set ever influences these normalization parameters.
Test-time ensemble perturbation (evaluation only).
We adopt the same testing setup proposed in GenCFD [21]. To evaluate how well our model captures the distribution of solutions $u(t)$ arising from a fixed, chaotic initial field $\bar{u}$, we generate a small perturbed ensemble around $\bar{u}$ and evolve each member with a high-fidelity PDE solver—but only at evaluation time. Specifically:
  1. 1.
    Micro-ensemble generation. Around the test initial condition $\bar{u}$, draw $M_{\mathrm{micro}}$ perturbed copies $\{\bar{u}_{j}\}_{j=1}^{M_{\mathrm{micro}}}$ by sampling uniformly in a ball of radius $\varepsilon$ centered at $\bar{u}$.
  2. 2.
    Reference propagation. Integrate each perturbed field $\bar{u}_{j}$ forward to time $t$ using the reference solver, producing end-states $\{u_{j}(t)\}$.
  3. 3.
    Empirical conditional law. Form the empirical measure
    $\hat{P}_{t}(\,\cdot\mid\bar{u})\;=\;\frac{1}{M_{\mathrm{micro}}}\sum_{j=1}^{M_{\mathrm{micro}}}\delta_{u_{j}(t)},$
    which approximates the true chaotic conditional law $P_{t}(\,\cdot\mid\bar{u})$.
When the input distribution itself is non-degenerate, one would first draw $M_{\mathrm{macro}}$ base fields from that distribution and then apply the micro-ensemble procedure to each. For our Dirac initial-condition tests we set $M_{\mathrm{macro}}=1$.
Training vs. evaluation.
During training the model only ever sees independent pairs $(u_{0},u_{1})$—one target per input—and never observes any ensemble. The above ensemble perturbation is used only to construct a ground-truth distribution at test time.
Appendix C Models and baselines
We tested our method against $4$ other baselines, out of which $2$ (namely GenCFD and its FNO-conditioned variant) are diffusion-based algorithms of the same kind. The other two baselines rely on traditional operator learning setup, with the UViT baseline satisfying also the purpose of an ablation study that quantifies the gain obtained by adding diffusion on top.
Shared Backbone. We stress that for GenCFD, GenCFD $\circ$ FNO and UViT we use the exact same UViT backbone architecture outlined in the earlier parameterization of our rectified flow as in Table A.1 for the sake of consistency. The architecture is further illustrated schematically in Figure A.1, which shows the UNet-based design in detail.
GenCFD
GenCFD [21] is an end-to-end conditional score-based diffusion model that directly learns the mapping from an input flow field $u_{0}$ to the target field $u_{1}$. It uses the same UViT backbone described in Table A.1, with noise levels $\sigma_{\tau}$ corresponding to the standard deviation of the added Gaussian perturbation, in contrast to our ReFlow model which evolves over a rectified diffusion time $\tau\in[0,1]$. Common choices for noise schedulers include exponential and tangent formulations. In our case, we consistently adopted an exponential noise schedule, as it yielded the best results across all datasets and metrics reported. For score-based diffusion models, either variance-exploding (VE) or variance-preserving (VP) formulations are typically used. In all GenCFD models, we opted for the VE setting throughout. During training we minimize the denoising score-matching loss over $(u_{0},u_{1})$ pairs sampled uniformly in $\tau\in[0,1]$, using Adam with initial learning rate $3\times 10^{-4}$ (cosine decay), batch size 16, and EMA decay 0.9999. As is common in score-based diffusion models, we applied denoiser preconditioning and used uniform weighting across noise levels. At inference we solve the corresponding SDE (via an Euler-Maruyama scheme with discretization $128$ steps) and incorporate reconstruction guidance to condition on $u_{0}$ so that we recover the conditional posterior measure $p(u_{1}|u_{0})$.
GenCFD $\boldsymbol{\circ}$ FNO
In this hybrid approach, a Fourier Neural Operator (FNO) [14] is first trained under an $\ell_{2}$ loss to predict $u_{1}$ from $u_{0}$. Its output $\hat{u}_{1}^{\rm FNO}$ is concatenated channel-wise with $u_{0}$ and fed into the GenCFD score network at both train and test time. All other architectural and optimization settings (backbone, scheduler, SDE sampler) remain identical to GenCFD, allowing the FNO to provide fast low-frequency priors while the diffusion stage refines fine-scale turbulent features.
Custom implementation.This schematic closely parallels the pipeline of [23], but with two key distinctions. First, we fully replace both the diffusion backbone and its denoising stages with GenCFD. Second, and most important, we condition not only on the FNO’s coarse solution but also on the original high-resolution initial condition. This ensures that our model retains fine-scale structures that would otherwise be lost if we relied solely on the FNO predictions.
FNO
Our FNO baseline follows Li et al. [14] and is implemented in PyTorch. First, each spatial location of the input field $u_{0}$ is lifted from $C_{\rm in}$ to 256 channels via a two-layer MLP with a SiLU activation, then projected down to a 128-channel hidden representation. This is followed by four FnoResBlocks: in each block, the hidden features undergo a Conditional LayerNorm (time embedding size 128), SiLU, and two spectral convolutions (SpectralConv) with $num\_modes=(24,24)$, combined through a learnable “soft-gate” residual skip. Finally, a two-layer projection MLP (128 → 256 → $C_{\rm out}$, with SiLU in between) produces the output field $\hat{u}_{1}$. We train for 100k iterations minimizing $\|u_{1}-\hat{u}_{1}\|_{2}^{2}$ with Adam (learning rate $10^{-3}$, batch size 16), and perform inference in one forward pass—no diffusion or autoregression is used.
UViT
Our UViT surrogate [28] uses the identical UNet2D backbone of Table A.1, injecting time embeddings and employing 8-head multi-head attention at the bottleneck. It is trained deterministically with MSE loss on $(u_{0},u_{1})$, using Adam at $3\times 10^{-4}$, batch size 16, for 100 k iterations (EMA 0.9999). For multi-step forecasting we roll out autoregressively, feeding each prediction back as the next input. This ablates out the diffusion component while matching all other design choices.
Baseline Model Summary and Implementation Setup
Table C.1 summarizes key architectural settings for all baseline models considered in this work. These configurations were selected to reflect the strongest performance achievable within a shared training budget of up to 120,000 iterations, as used in the quantitative comparison reported in Table  5.1.
Padding and boundary conditions.
Each dataset required padding choices tailored to its physical boundary conditions. SL and RM, which exhibit periodic boundaries, use circular padding in all convolutional layers. This significantly improved the performance of the diffusion-based models. In contrast, CS features non-periodic boundaries and is thus trained with standard zero padding. These padding choices apply to all UNet-based architectures except ReFlow, which uniformly uses zero padding across all datasets regardless of boundary conditions.
Table C.1: Summary of architectural depth and approximated parameter count for each model evaluated. All UNet-based models (ReFlow, GenCFD, GenCFD $\circ$ FNO, and UViT) share a consistent backbone design to ensure consistency across diffusion and non-diffusion baselines, while the FNO baseline uses a structurally different architecture. These configurations reflect the best-performing settings under a training budget of up to 120,000 iterations, as used in the statistical comparisons reported in Table 5.1.
Model # Levels # Blocks # Params Noise Schedule Diffusion Scheme
ReFlow 3 4 22M uniform rectified continuous
time variable $t\in[0,1]$
GenCFD 2 4 5M exponential variance-exploding
GenCFD $\circ$ FNO 2 4 5M exponential variance-exploding
UViT 2 4 5M
FNO 4 11M
Appendix D Results
D.1 Performance Metrics
To compare our generated ensembles against the reference Monte Carlo samples, we employ three complementary measures:
  • Mean-field error. Let $\mu_{\rm ref}(x)$ and $\mu_{\rm model}(x)$ denote the pointwise spatial means of the reference and model ensembles, respectively. We measure their discrepancy by the $L^{2}$-norm
    $e_{\mu}\;=\;\bigl{\|}\,\mu_{\rm ref}\;-\;\mu_{\rm model}\bigr{\|}_{2}\,.$
  • Standard-deviation error. Similarly, let $\sigma_{\rm ref}(x)$ and $\sigma_{\rm model}(x)$ be the pointwise standard deviations. We therefore compute
    $e_{\sigma}\;=\;\bigl{\|}\,\sigma_{\rm ref}\;-\;\sigma_{\rm model}\bigr{\|}_{2}\,,$
    Both metrics are reported after normalization by the $L^{2}$ norm of the ground truth.
  • Average 1-Wasserstein distance. For a fixed initial condition $u_{0}$, let $\{u^{(j)}_{\rm ref}(x)\}_{j=1}^{N}$ and $\{u^{(j)}_{\rm model}(x)\}_{j=1}^{N}$ be the ensembles of final-time values produced by the reference solver and our model, respectively. Since these fields are discretized on $M$ spatial points $\{x_{i}\}_{i=1}^{M}$, we obtain two empirical 1D distributions at each $x_{i}$. Denoting their inverse CDFs by $F^{-1}_{\rm ref}(\cdot\,;\,x_{i})$ and $F^{-1}_{\rm model}(\cdot\,;\,x_{i})$, the pointwise Wasserstein-1 distance is
    $W_{1}\bigl{(}p_{\rm ref}(\cdot\mid x_{i}),\,p_{\rm model}(\cdot\mid x_{i})\bigr{)}\;=\;\int_{0}^{1}\bigl{|}\,F^{-1}_{\rm ref}(q;x_{i})\;-\;F^{-1}_{\rm model}(q;x_{i})\bigr{|}\,\mathrm{d}q.$
    We then report the spatial average
    $\overline{W}_{1}\;=\;\frac{1}{M}\sum_{i=1}^{M}W_{1}\bigl{(}p_{\rm ref}(\cdot\mid x_{i}),\,p_{\rm model}(\cdot\mid x_{i})\bigr{)}.$
Energy spectra. Finally, to assess how well different scales are captured, we compute the discrete energy spectrum of each velocity field. Given a sample $u(x)$ on a $d$-dimensional periodic grid with spacing $\Delta$, let $\widehat{u}_{k}$ be its Fourier coefficient at integer wavevector $k\in\mathbb{Z}^{d}$. We bin modes by their $\ell_{1}$ radius $|k|_{1}=k_{1}+\dots+k_{d}$, defining
$E_{r}\;=\;\frac{\Delta^{d}}{2}\,\sum_{\begin{subarray}{c}k\in\mathbb{Z}^{d}\\ |k|_{1}=r\end{subarray}}\bigl{\|}\widehat{u}_{k}\bigr{\|}^{2},\quad r=0,1,2,\dots$
and compare the ensemble-average spectra of reference and model. Our plots show $E_{r}$ versus $r$ on log–log axes to highlight scale-by-scale agreement.
D.2 Training Efficiency Observations
To assess training efficiency, we compare ReFlow with GenCFD. For fairness, we use a GenCFD variant with the same setup and parameter count (22M), matching the three-level UViT architecture of ReFlow instead of its standard lightweight configuration (5M). Both models were trained for up to 160,000 iterations with a batch size of 16.
Figure D.1 summarizes performance over the training trajectory. At approximately 100,000 steps, ReFlow consistently outperforms GenCFD across all key metrics: relative $L^{2}$ error in the mean ($e_{\mu}$), relative $L^{2}$ error in the standard deviation ($e_{\sigma}$), and the average pointwise Wasserstein-1 (W1) distance—evaluated for both $u_{x}$ and $u_{y}$ velocity components. Notably, the $y$-axis uses a logarithmic scale to emphasize differences in error magnitude across training steps.
Across all metrics, ReFlow begins to outperform GenCFD at around 100,000 training steps and continues to improve at a sharper rate. While neither model reaches full convergence within 160,000 iterations, ReFlow consistently exhibits lower errors and more favorable metric trajectories in later stages of training. These results support the observation that ReFlow is more training-efficient within a fixed compute budget and ultimately achieves stronger generalization performance in this setting.
To further probe the long-term capabilities of GenCFD, we extended its training to 600,000 iterations and checked whether it eventually surpasses ReFlow’s performance at its final recorded checkpoint (150,000 steps). Table D.1 summarizes the earliest training steps at which GenCFD overtakes ReFlow for each evaluation metric. While GenCFD does catch up and exceed ReFlow in several cases, it never surpasses ReFlow on the Wasserstein-1 distance for $u_{y}$ within the training horizon considered.
Table D.1: Number of training iterations at which GenCFD surpasses ReFlow’s performance at 150,000 iterations. If not surpassed, indicated by “—”.
Metric GenCFD Iteration GenCFD Value ReFlow Value @ 150,000 iter.
$u_{x}$ Mean Error ($e_{\mu}$) 280,000 3.33 ×$\times 10^{-2}$ 3.45 ×$\times 10^{-2}$
$u_{y}$ Mean Error ($e_{\mu}$) 400,000 1.70 ×$\times 10^{-1}$ 1.88 ×$\times 10^{-1}$
$u_{x}$ Std Error ($e_{\sigma}$) 400,000 6.48 ×$\times 10^{-2}$ 7.24 ×$\times 10^{-2}$
$u_{y}$ Std Error ($e_{\sigma}$) 340,000 7.91 ×$\times 10^{-2}$ 7.97 ×$\times 10^{-2}$
$u_{x}$ Wasserstein-1 ($W_{1}$) 200,000 3.89 ×$\times 10^{-2}$ 3.95 ×$\times 10^{-2}$
$u_{y}$ Wasserstein-1 ($W_{1}$) 4.11 ×$\times 10^{-2}$
Figure D.1: Training efficiency comparison. Performance of ReFlow and GenCFD on the Macro-Micro SL ensemble dataset evaluated at multiple training checkpoints. Metrics match those used in the main text: relative $L^{2}$ error of the mean ($e_{\mu}$) and standard deviation ($e_{\sigma}$), and the average 1-point Wasserstein-1 distance (W1), computed for both velocity components ($u_{x}$, $u_{y}$). The $y$-axis uses a logarithmic scale, while the $x$-axis (training iteration) is linear.
D.3 Experimental Advantages of Straightness
This subsection presents empirical evidence that the “rectified” trajectories generated by our Rectified Flow model (ReFlow) enable faster convergence and yield higher-fidelity reconstructions compared to the baseline GenCFD diffusion approach.
Evaluation settings.
We use the CloudShock dataset throughout for exposition purposes. However, we stress that the same results hold across datasets. For each initial condition, both ReFlow and GenCFD are allotted $T=10$ diffusion steps over normalized diffusion time $\tau\in[0,1]$. We record the model outputs at five evenly spaced timesteps $t=\{0.0,0.25,0.50,0.75,1.0\}$ and analyze:
  • Trajectory inpainting of the density field $\rho$.
  • Latent-Space Trajectory Visualization via PCA via PCA.
  • Per-Sample Average Error Evolution: average MSE across physical channels.
  • Evolution of the Energy Spectrum: radial power-spectrum of energy $E$ as inpaintig progresses.
All other components (backbone, noise schedule, batch size) are held equal for fair comparison.
Trajectory Inpainting
We compare Rectified Flow (ReFlow) and GenCFD reconstructions of the density field $\rho$ over five normalized diffusion timesteps $t\in\{0.00,0.25,0.50,0.75,1.00\}$. ReFlow inpaints fine-scale features in far fewer steps, while GenCFD remains overly diffusive. The leftmost panel corresponds to the initial condition $u_{0}$ upon which both model are conditioned, while the following panels showcase the evolution of the model’s output as diffusion time progresses.
Figure D.2: Density trajectories for Sample 1: ReFlow (top) vs. GenCFD (bottom). ReFlow already recovers sharp shock fronts by diffusion time $t=1.0$.
Figure D.3: Density trajectories for Sample 2.
Figure D.4: Density trajectories for Sample 3.
Latent-Space Trajectory Visualization via PCA
To quantify how directly each model moves through the data manifold, we flatten the $(H\times W\times C)$ field at each timestep into a single feature vector, stack all Rectified Flow and GenCFD vectors, fit a 3-component PCA, and then project each timestep into PC1/PC2/PC3. A straighter path indicates fewer diffusive detours.
(a) Sample 1
(b) Sample 2
(c) Sample 3
(d) Sample 4
(e) Sample 5
(f) Sample 6
Figure D.5: 3D PCA latent-space trajectories (Samples 1–6). We project the flattened snapshots at each of $T$ timesteps into the top three PCA components. Rectified Flow (blue) consistently traces a straighter path than GenCFD (red), confirming its “rectified” evolution. Circles mark $t=0$, squares mark $t=1$, and intermediate timesteps are annotated.
To further illustrate the nature of each model’s generative dynamics, we visualize single-sample generation trajectories from both models in a 2D PCA space for the RM dataset (Figure D.6). For each physical channel ($\rho$, $m_{x}$, $m_{y}$, $p$), we apply PCA to the full set of ground truth snapshots to define a common projection basis. We then overlay the generative trajectory of a single sample in this space. To contextualize the generated dynamics, we add kernel density estimates (KDEs) showing the empirical data distribution in the same projected space.
ReFlow (top) again traces a straighter, more compact path over just 10 rectified steps, moving coherently along the data manifold. GenCFD (bottom), by contrast, uses 128 stochastic denoising steps and produces a more diffuse and nonlinear trajectory, often deviating from high-density regions of the ground truth distribution. While the PCA projection inherently reduces dimensional complexity, the consistent difference in shape and alignment supports the conclusion that ReFlow’s rectified flow better preserves manifold structure and sample efficiency.
Figure D.6: 2D PCA latent-space trajectories. We project the flattened velocity snapshots from the RM dataset into a 2D PCA space, computed independently for each channel ($\rho$, $m_{x}$, $m_{y}$, $p$) across all $T$ time steps. To visualize the underlying distribution of target fields, we apply kernel density estimation (KDE) per channel. The ReFlow model (top) evolves over 10 rectified time steps, producing visibly straighter and more structured trajectories compared to the score-based diffusion model GenCFD (bottom), which uses 128 stochastic steps. Although the projection into two principal components necessarily truncates the high-dimensional dynamics, the qualitative difference remains evident, illustrating the smoother and more directed progression learned by ReFlow.
Per-Sample Average Error Evolution
We compute the mean-squared error at each timestep $t$ and average across all $C$ channels:
$\overline{\mathrm{MSE}}(t)=\frac{1}{C}\sum_{c=1}^{C}\mathrm{MSE}_{c}(t).$
Figure D.7 shows, for Samples 1–6, the log-scaled average MSE (left) and the normalized error $\overline{\mathrm{MSE}}(t)/\overline{\mathrm{MSE}}(0)$ (right). Rectified Flow (blue) consistently reduces error faster and to a lower residual than GenCFD (red).
(a) Sample 1
(b) Sample 2
(c) Sample 3
(d) Sample 4
(e) Sample 5
(f) Sample 6
Figure D.7: Per-sample average MSE evolution. Each subfigure plots, left half: $\log_{10}\overline{\mathrm{MSE}}(t)$; right half: $\overline{\mathrm{MSE}}(t)/\overline{\mathrm{MSE}}(0)$. Rectified Flow (blue) rapidly drives error down and keeps it orders of magnitude below GenCFD (red) across all samples.
Evolution of the Energy Spectrum
We examine the 2D radial power spectrum of the energy field $E$ at five normalized diffusion times $\tau\in\{0.00,0.25,0.50,0.75,1.00\}$. Each figure below shows log–log power vs. wavenumber for Rectified Flow (blue), GenCFD (red), and the ground-truth target (black dashed), with inset log-MSE error annotations. The leftmost snapshot corresponds to the spectrum of the initial data $u_{0}$ upon which the model is conditioned. Subsequent panels track the spectral evolution of the model predictions, as diffusion time progresses.
Figure D.8: Energy spectrum evolution for Sample 1. ReFlow (blue) outperforms GenCFD (red) in capturing high-wavenumber energy.
Figure D.9: Energy spectrum evolution for Sample 2. ReFlow (blue) remains closer to the ground truth (dashed) than GenCFD (red).
Figure D.10: Energy spectrum evolution for Sample 3. The inset log-MSE annotations quantify ReFlow’s advantage at each $t$.
Figure D.11: Energy spectrum evolution for Sample 4. ReFlow’s spectrum (blue) gets closer to the target (dashed) as $\tau\to 1.0$.
Figure D.12: Energy spectrum evolution for Sample 5. ReFlow’s superior high-$k$ fidelity is evident throughout.
Figure D.13: Energy spectrum evolution for Sample 6. ReFlow better captures the ground-truth spectrum’s slope across all timesteps.
Appendix E Computational Considerations
[…]
Table E.1: Aggregate hardware utilization for all “usual-setup” runs.
Metric Range Typical
GPU power draw (W) 200–260 230
GPU utilization (%) 30–80 55
GPU memory allocated (GB) 12–17 15
GPU temperature (°C) 40–60 50
Process GPU memory usage (%) 60–80 70
Process GPU memory access time (%) 10–50 30
Process CPU threads in use 7 7
System CPU utilization (per core, %) 60–100 80
Process memory in use (GB) up to 50 25
Disk I/O read (MB total) up to $1.2\times 10^{6}$ $5\times 10^{5}$
Disk utilization (%) 70–75 72
Network traffic (bytes total) $\approx 3\times 10^{13}$ $3\times 10^{13}$
Since all datasets are paired with models of identical size and the hardware envelope (GPU power, memory, temperature, CPU threads, I/O) is effectively the same across runs, any observed performance differences are attributable to the sampling strategy (noise schedule, ODE steps) rather than to model-size or hardware variations (the network traffic is due to Wandb usage).
All experiments were run on a single NVIDIA Quadro RTX 6000 GPU (Driver 570.124.06, CUDA 12.8), operating in persistence mode.
Each model was trained for 100 000 iterations. We explored architectures ranging from approximately 5 million to 10 million trainable parameters; across this range we observed no significant differences i n final predictive performance or convergence behavior. All other hyperparameters (learning rate schedules, batch sizes, noise schedules, etc.) were held fixed when evaluated against each other, ensuring that differences in accuracy and runtime reflect only the conditioning and sampling strategies under study.
Appendix F Additional Visualizations
Test-time ensemble perturbation (evaluation only).
Following the protocol of GenCFD [21], we construct small “micro-ensembles” around each test initial condition $\bar{u}$ by sampling $M_{\mathrm{micro}}$ perturbed fields uniformly in an $\varepsilon$–ball centered at $\bar{u}$. Each perturbed member is then advanced to the evaluation time $t$ with a high-fidelity PDE solver to produce a reference ensemble $\{u^{(j)}_{\rm ref}(t)\}_{j=1}^{M_{\mathrm{micro}}}$. Our model generates a corresponding ensemble $\{u^{(j)}_{\rm model}(t)\}_{j=1}^{M_{\mathrm{micro}}}$ from the same perturbed inputs, but without ever observing ensembles during training (which only uses independent pairs $(u_{0},u_{1})$). We compare these two ensembles via mean-field error $e_{\mu}$, standard-deviation error $e_{\sigma}$, and spatially averaged 1-Wasserstein distance $\overline{W}_{1}$ (see SM §D).
Out-of-distribution testing (cloud–shock and shear-flow). Because the micro-perturbations render each initial field slightly OOD, this evaluation probes the model’s ability to generalize under small chaotic perturbations. In these tasks, our Rectified Flow ensemble closely tracks the reference uncertainty and multiscale statistics, whereas deterministic baselines collapse to a single trajectory and fail to capture the conditional variability.
In-distribution scaling (Richtmyer–Meshkov). For the Richtmyer–Meshkov dataset, we additionally study the effect of training set size using a held-out test set from the same distribution. Here, even deterministic approaches improve with more data, but still underperform on the micro-ensemble test: they require significantly larger training sets to match ReFlow’s accuracy and exhibit poor calibration under small perturbations.
In this appendix we provide a comprehensive set of qualitative comparisons between the baselines (GenCFD variants, UViT, FNO) models and our Rectified Flow (ReFlow) approach across three representative 2D fluid-flow benchmarks (Richtmyer–Meshkov, shear layer roll-up, and cloud-shock interaction). For each task we show:
  • Mean fields Ensemble averages reveal how each deterministic model diffuses fine-scale features (UVit/FNO) versus perserving sharp interfaces (ReFlow, GenCFD variants).
  • Uncertainty maps Per-pixel standard deviations illustrate whether uncertainty localizes in physically meaningful regions (e.g. shocks, shear interfaces) or remains overly diffuse.
  • Random samples To qualitatively assess generative fidelity, we include representative random samples from ReFlow alongside outputs from selected baselines, specifically the deterministic FNO and the hybrid GenCFD $\boldsymbol{\circ}$ FNO. The latter combines FNO’s low-frequency structural prediction with GenCFD’s generative modeling of fine-scale turbulence. This comparison highlights ReFlow’s ability to produce realistic, high-resolution flow fields that preserve physical coherence across different scales.
  • Spectral & pointwise comparisons (Fig. F.4: Richtmyer–Meshkov global $E(k)$, local spectra, and pointwise density histograms; Fig. F.7: Shear-layer global $u_{x}$ spectra and pointwise $u_{x}$ distributions; Fig. F.11: Cloud–shock global $m_{y}$ spectra and pointwise energy distributions) — combined Fourier-space and histogram panels quantify multiscale fidelity and calibration of the predictive ensembles.
Together, these visualizations underscore ReFlow’s ability to maintain sharper physical structures, produce better-calibrated uncertainties, and faithfully reproduce multiscale spectral statistics.
Figure F.1: Richtmyer–Meshkov Mean Field Comparison: Mean velocity fields for the Richtmyer–Meshkov dataset, shown across all physical channels and models. In many cases, predictions closely resemble the ground truth, with differences often imperceptible, underscoring the overall high quality and consistency of the generated mean fields.
Figure F.2: Richtmyer–Meshkov Uncertainty Comparison: Per-pixel standard deviation fields from the Richtmyer–Meshkov dataset across all models. Stochastic generative models capture spatially coherent uncertainty patterns, outperforming the deterministic baselines (FNO and UViT), which exhibit limited variability and poor uncertainty localization.
Figure F.3: Richtmyer–Meshkov Individual Sample Comparison: A representative random test sample from the Richtmyer–Meshkov dataset, visualized across all four physical channels ($\rho$, $m_{x}$, $m_{y}$, $p$). All models produce physically plausible results, with only little variation in fidelity to fine-scale structures. ReFlow offers the closest visual match to the ground truth, preserving coherent details and sharp interfaces. Notably, the GenCFD $\boldsymbol{\circ}$ FNO hybrid enhances spectral richness by reconstructing high-frequency features based on a low-frequency FNO prior, illustrating its capacity to recover sharper modes even when starting from a coarse prediction.
Figure F.4: Richtmyer–Meshkov: spectral and pointwise distribution comparisons between UVit and Rectified Flow.
Figure F.5: Shear-Layer Mean and Standard Deviation Comparison: Visualization of mean-field ($\mu$) and standard-deviation ($\sigma$) predictions for the Shear-Layer dataset across both velocity components ($u_{x}$, $u_{y}$). For each model, the left half shows the predicted mean fields and the right half shows per-pixel standard deviations. The top row contains the reference ground truth. Despite the reduced channel complexity of this dataset, clear differences emerge in how well models capture coherent structures and localized uncertainties. ReFlow preserves sharper features and more accurately reflects the spatial distribution of uncertainty.
Figure F.6: Shear-Layer Individual Sample Comparison. A representative random test sample from the Shear-Layer dataset, shown across both physical channels ($u_{x},u_{y}$). All models generate physically plausible results, though with varying degrees of detail and sharpness. ReFlow and GenCFD $\boldsymbol{\circ}$FNO most closely resemble the ground truth, preserving fine-scale structures and coherent interfaces. GenCFD $\boldsymbol{\circ}$FNO enhances high-frequency detail by leveraging the low-frequency prior from the initial FNO prediction.
Figure F.7: Shear-layer: global $u_{x}$ spectra and pointwise $u_{x}$ distributions.
Figure F.8: Cloud-Shock Mean Field Comparison: Mean velocity fields for the Cloud-Shock dataset, shown across all physical channels and models.
Figure F.9: Cloud–Shock Uncertainty Comparison: Per-pixel standard deviation visualizations for the Cloud–Shock dataset across all models.
Figure F.10: Cloud–Shock Individual Sample Comparison: A representative random test sample from the Cloud–Shock dataset, displayed across all physical channels. While all models produce plausible results, ReFlow and the hybrid GenCFD $\boldsymbol{\circ}$ FNO most closely align with the ground truth, particularly in preserving fine-scale detail. The spectral reconstruction effect of the hybrid model is especially evident, recovering high-frequency structures from the coarse FNO prediction and enhancing the overall visual fidelity.
Figure F.11: Cloud-shock: spectral and pointwise energy comparisons.