### v2

# Rectified Flows for Fast Multiscale Fluid Flow Modeling

Abstract
Statistical surrogate modeling of fluid flows is challenging due to multiscale dynamics and strong sensitivity to initial conditions. Conditional diffusion models can achieve high fidelity, but typically require hundreds of stochastic steps at inference. We introduce a rectified-flow surrogate that learns a time-dependent conditional velocity field transporting input-to-output laws along nearly straight trajectories. Sampling reduces to solving a deterministic ODE along this learned transport, so each function evaluation is substantially more effective: on multi-scale 2D benchmarks we match diffusion-class posterior statistics with as few as $8$ ODE steps versus $\geq\!128$ steps for score-based diffusion.
On the theory side, we develop a law-level analysis tailored to conditional PDE forecasts. First, we formalize the link between our evaluation criterion—one-point Wasserstein distances on fields—and the k=1k\!=\!1 correlation-marginal viewpoint in statistical solutions. Second, we provide a one-step error decomposition for the learned pushforward law into a coverage (high-frequency tail) term controlled by structure functions (equivalently, by spectral decay), and a fit term controlled directly by the training objective. Third, we show how straightness in rectification time governs local truncation error for ODE sampling, yielding step-count requirements and explaining why rectified transports admit large, stable steps.
Guided by this picture, we propose a curvature-aware sampler that tracks an EMA-based straightness proxy and adaptively blends and steps the velocity during inference. Across multiscale incompressible and compressible 2D flows, our method matches diffusion models in Wasserstein statistics and energy spectra, preserves fine-scale structure missed by MSE baselines, and delivers high-resolution conditional samples at a fraction of the inference cost.
1 Introduction
Partial differential equations (PDEs) Evans (2010) are the backbone of modeling physical phenomena, from atmospheric dynamics to aerodynamics and magnetohydrodynamics. Currently, solutions of PDEs are simulated with high-fidelity numerical solvers such as finite difference, finite element, and spectral methods Quarteroni and Valli (1994). These are often prohibitively expensive for many–query tasks Quarteroni et al. (2015) such as uncertainty quantification, optimization, and real‐time prediction. In particular, these solvers encounter a significant challenge for classes of non-linear PDEs, whose solutions exhibit extreme sensitivity to initial and boundary data, leading to chaotic behavior and complex multi‐scale features. In such settings, one seeks, not a deterministic trajectory, but rather the statistical solution of the PDE, i.e., the push-forward of the underlying probability measure on inputs (initial and boundary conditions etc) by the PDE solution operator Fjordholm et al. (2017). Currently, these statistical solutions are sampled with ensemble Monte Carlo method by repeatedly calling high‐fidelity PDE solvers Lanthaler et al. (2021b). Given the computational cost of the PDE solver, ensemble methods become infeasible as ensemble sizes grow.
Machine learning and AI based algorithms are increasingly being explored for the data-driven simulation of PDEs Mishra and Townsend (2024). In particular, neural operators Kovachki et al. (2023), which directly seek to learn the PDE solution operator from data, are already widely used in a PDE context Mishra and Townsend (2024). Examples of neural operators have been proposed in Li et al. (2020a; b); Lu et al. (2021); Wu et al. (2024); Hao et al. (2023); Li et al. (2023); Alkin et al. (2024) and references therein.
However, these neural operators have been found to be unsuitable for simulating PDE with chaotic, multiscale solutions such as governing equations of fluid dynamics. For instance, recently in Molinaro et al. (2024), it was shown both theoretically and empirically that neural operators, trained to minimize mean absolute (square) errors, will always collapse to the ensemble mean when used for simulating statistics of PDEs with chaotic multiscale solutions.
Given this context, generative AI models—especially score‐based and diffusion models—have recently shown remarkable success at learning complex, high‐dimensional distributions by progressive denoising score matching Song et al. (2020). In the context of chaotic PDEs, GenCFD Molinaro et al. (2024) introduced a conditional diffusion framework that preserves spectral content across fine scales, mitigating the “spectral collapse” of naive MSE‐trained surrogates, see also Oommen et al. (2024); Gao et al. (2024a; b; c). However, sampling a score-based diffusion model requires simulating a reverse‐time stochastic differential equation (SDE) over 100+ discretization steps, each invoking a large ML model on a moderate-to-high resolution spatial grid, resulting in high runtimes that undermine real‐time or resource‐constrained applications.
To overcome this limitation of diffusion models, we adapt rectified flows Liu et al. (2022): a deterministic ODE‐based analogue of diffusion that “straightens” transport paths in probability space. Rather than small random increments, rectified flows solve an ODE whose velocity field is trained to match the diffusion model’s instantaneous score, but whose trajectories remain nearly linear. This structure permits large time integration, reducing the number of solver steps by an order of magnitude or more, while provably recovering the same target distribution.
What is evaluated, and why.
Because statistical solutions are measures on function spaces, the most robust comparison is via observables or correlation marginals Lanthaler et al. (2021a). Our evaluation metric in Section 6.2—the one-point Wasserstein distance—is exactly the Wasserstein discrepancy between k=1k\!=\!1 correlation marginals (pointwise push-forwards), averaged over space. This aligns the experimental metrics (Table 1) with the law-level theoretical statements in Sections 4.14.3.
Hence, our main contributions are:
  • Rectified flows for conditional PDEs. We learn a time-dependent velocity field that transports a conditional law of fields from a simple reference (Gaussian noise conditioned on the input) to the target PDE pushforward law. The key theoretical point is correctness at the level of laws: the barycentric rectified velocity induces an ODE whose flow maps the initial law to the target law (SM C).
  • Curvature-aware integration. The sampling cost is governed by the amount of time-bending (straightness/curvature in rectification time) of the learned vector field along typical trajectories. We formalize: (i) which curvature quantity controls discretization error, (ii) why a simple EMA deviation is a principled proxy for that curvature, and (iii) why the blending step is the minimizer of a quadratic (Tikhonov) correction (SM E).
  • Multi-scale accuracy at speed. At fixed grid resolution there is an irreducible law-level error floor coming from unresolved Fourier scales. We express this as a coverage term controlled by structure functions / spectral tails (Crefapp:sec:coverage-structure). Rectified flows reduce curvature so that the best attainable (bandlimited) law is reached with far fewer sampling steps. This is precisely what is seen in Table 1 and in the spectral evolution plots (Figure 13).
2 Approach
Modeling multi-scale flows and their statistical solutions. We focus on time-dependent PDEs of the general form
$\partial_{t}\,u(x,t)\;+\;\mathcal{L}\!\bigl(u,\nabla_{x}u,\nabla_{x}^{2}u,\dots\bigr)\;=\;0,\quad x\in D\subset\mathbb{R}^{d},\;t\in(0,T),$ (1)
subject to boundary conditions $\mathcal{B}(u)=u_{b}$ on $\partial D\times(0,T)$ and initial data $u(x,0)=\bar{u}(x)$. Here, the solution $u:D\mapsto{\mathbb{R}}^{m}$ is evolved in time, with respect to the differential operator $\mathcal{L}$. Concrete examples of (1) are the incompressible Navier-Stokes and compressible Euler equations.
A recurring theme in non-linear, multi-scale fluid flows is extreme sensitivity to initial or boundary data, which can lead to chaotic or near-chaotic behavior. Rather than fix a single $u_{0}$, one may consider an initial measure $\mu_{0}$ to represent uncertain or variable initial states, and then study the statistical solution $\mu_{t}=\mathcal{S}^{t}_{\#}\mu_{0}$ pushed forward by the PDE solution operator $\mathcal{S}^{t}$. In strongly nonlinear regimes, $\mu_{t}$ spreads significantly in function space, creating complex, multi-scale distributions of possible flow fields. Accurately approximating this evolving distribution is essential for uncertainty quantification and design under rough conditions, but doing so with high-fidelity PDE solvers can be prohibitively expensive when large ensembles are required.
High cost of computing statistical solutions. In practice, characterizing $\mu_{t}$ directly often requires performing many forward simulations of the PDE solver, each initialized with a slightly perturbed $u_{0}$. As each simulation might require high-resolution discretizations (e.g. millions of degrees of freedom) and multiple time steps, the overall cost for constructing a sizable ensemble rapidly becomes prohibitive. Even with modern high-performance computing platforms, running hundreds or thousands of high-fidelity statistical simulations can be infeasible. Hence, there is a pressing need for more efficient surrogate models that can sample physically realistic final states without enumerating every realization of the forward PDE solve. Succinctly put, we aim to provide a solution to the following problem
How can we efficiently capture statistical quantities of interest for the pushforward measure $S^{t}_{\#}\mu_{0}$ for highly sensitive solution operators $S^{t}$?
Observables and correlation marginals (what metrics actually probe). A classical way to formalize “quantities of interest” for $\mu_{t}$ is via observables and their induced correlation marginals Lanthaler et al. (2021a). For $k=1$, the pointwise evaluation map at $x\in D$,
$\mathrm{ev}_{x}:\;u\mapsto u(x)\in\mathbb{R}^{m},$
pushes $\mu_{t}$ to the one-point marginal $(\mathrm{ev}_{x})_{\#}\mu_{t}$ on $\mathbb{R}^{m}$ – strictly speaking, one studies the pushforwards measures $\mathrm{ev}_{x}\circ i_{\epsilon}$, here $(i_{\epsilon}u):=(\phi_{\epsilon}\ast u)$ where $\phi_{\epsilon}$ is a compactly supported mollifier, so that point-wise evaluation is well-defined; for the sake of brevity, we will not make this distinction explicit in the sequel. Our one-point Wasserstein metric (Section 6.2) is the spatial average of $W_{1}\big((\mathrm{ev}_{x})_{\#}\mu_{t},(\mathrm{ev}_{x})_{\#}\widehat{\mu}_{t}\big)$. By Kantorovich–Rubinstein duality, controlling this quantity controls the discrepancy of all pointwise 1-Lipschitz observables, averaged over space. This is the bridge between Table 1 and the law-level bounds in Section 4.1 and Section 4.3.
Rectified flows for faster sampling. To mitigate these inefficiencies, we adopt the framework of rectified flows Liu et al. (2022): a deterministic alternative to conditional diffusion designed to reduce the number of required integration steps. In essence, rather than generating samples via small random “diffusion-reversal” increments, rectified flows construct an ordinary differential equation (ODE) with nearly straight trajectories in function space. This straighter transport map can be integrated more aggressively in time, enabling sampling with far fewer solver steps. Importantly, rectified flows retain the desirable property of matching the same final distributions that standard diffusion models learn, but at a fraction of the runtime cost.
Figure 1: Overview of ReFlow. (a): ReFlow closely matches high-order statistical moments (in the Wasserstein $W_{1}$-sense) of fluid flows by learning a neural surrogate to approximate the push-forward of the distribution $p(u_{\tau|\tau=0}|u_{i})$, while minimizing inference-time computational costs. (b) Macro-architecture overview: A UViT based network predicts the velocity field $v_{\theta}$, conditioned on Gaussian noise, initial data $u_{i}$ and boundary conditions, and diffusion time $\tau$. Noise and initial / boundary conditions are concatenated as inputs. (c): Results for the cylindrical shear flow dataset. (d): The KDE plots monitor the evolution of a down-projected version of the push-forward measure $p_{\tau}:=X_{\theta}(\tau)_{\#}p(u_{\tau=0}|u_{i})$, where $X_{\theta}$ is the flow induced by the learned vector field $v_{\theta}$. Results for the Richtmyer Meshkov dataset showing flow matching between noise ($p_{\tau=0}$) and target ($p_{\tau=1}$) distributions. ReFlow solves an ODE, learning a constant velocity field that enables straight-line inference paths (visualized with principal component analysis and kernel density estimation at different diffusion times).
3 Related Work
3.1 Diffusion Models in Fluid Dynamics
Conditional diffusion surrogates. GenCFD Molinaro et al. (2024) trains a conditional score-based diffusion model that starts from an initial or low-resolution flow field and iteratively refines it through hundreds of denoising steps, thereby capturing multi-scale turbulent structure. Similarly, the method of Oommen et al. (2024) conditions a diffusion model on a coarse fluid solution (generated by a learned operator) and ”super-resolves” the fine-scale flow. Given these similarities, we refer to such diffusion-based solvers collectively as conditional diffusion PDE solvers.
Rectified flow alternative. Each sample typically needs hundreds of network calls, limiting large-scale or real-time use. Our method replaces the stochastic SDE with a deterministic ODE whose nearly straight trajectories require far fewer steps, achieving similar statistical fidelity at much lower inference cost (see Sections 5.6 and 6). Crucially, we find that rectified flows are actually more efficient from the standpoint of training-time required to reach the same accuracy level with the conditional diffusion approaches; this is also confirmed by earlier work Esser et al. (2024) exploring the scalability of rectified flows with transformer-based backbones.
4 Rectified Flows: An ODE-Based Alternative
4.1 Mathematical Formulation
Overview.
Rectified flows provide a deterministic way to transform one probability distribution into another, offering an alternative to stochastic diffusion-based methods. Instead of traversing curved, noise-driven paths, rectified flows attempt to transport mass along nearly “straight” trajectories in state space, which can dramatically reduce sampling costs.
Throughout, our random variables are state fields $u:\mathcal{D}\!\to\!\mathbb{R}^{m}$ (e.g. $m=2$ for velocity, or $m=4$ for compressible state variables) living in a separable space $\mathcal{U}=L^{2}(\mathcal{D};\mathbb{R}^{m})$ (or $H^{s}$, $s>\frac{d}{2}$). All theory below extends verbatim from $\mathbb{R}^{d}$ because $\mathcal{U}$ is Polish; in practice we work with a Galerkin discretisation $\mathbb{R}^{d_{N}}\subset\mathcal{U}$, so every integral is evaluated by Monte-Carlo on its finite grid surrogate.
In all that follows, the diffusion time $\tau\in[0,1]$ is a purely scale-ordering variable: $\tau=0$ denotes the coarsest, lowest–frequency representation of a state field, while $\tau=1$ restores full spatial resolution; it is not the physical PDE time but merely an index that governs the progressive refinement of scales along the rectified flow.
Coupling $\gamma$ as a statistical transport blueprint. Let $\mu_{0},\mu_{1}\in\mathcal{P}(\mathcal{U})$ be the initial and target field laws. Fix any coupling $\gamma\in\mathcal{P}(\mathcal{U}\times\mathcal{U})$ with $(\pi_{0})_{\#}\gamma=\mu_{0}$ and $(\pi_{1})_{\#}\gamma=\mu_{1}$ (in practice: the empirical pairing of input/output fields in the dataset). Think of $(U_{0},U_{1})\sim\gamma$ as drawing a random chord in state space, from $U_{0}$ to $U_{1}$.
For $\tau\in[0,1]$ define the linear chord point
$U_{\tau}:=(1-\tau)U_{0}+\tau U_{1},\quad\rho_{\tau}:=\operatorname{Law}(U_{\tau})=(T_{\tau})_{\#}\gamma,\quad T_{\tau}(u_{0},u_{1})=(1-\tau)u_{0}+\tau u_{1}.$
Thus $\rho_{\tau}$ is the population of intermediate states obtained by moving along dataset chords by a fraction $\tau$.
The barycentric velocity is the “average chord direction” at $(u,\tau)$. Along a single chord, the displacement $U_{1}-U_{0}$ is constant in $\tau$. However, many chords pass through the same intermediate state $u$ at the same $\tau$; the correct population-level velocity is the average direction among those chords. Concretely, define
$v^{\star}(u,\tau)\;:=\;\mathbb{E}\bigl[\,U_{1}-U_{0}\,\big|\,U_{\tau}=u\,\bigr],\quad\rho_{\tau}\text{-a.e. }u.$ (2)
This is the clean “one-line” definition of the barycentric field: it is a conditional expectation, so it is automatically the unique $L^{2}(\rho_{\tau})$ minimizer of a regression problem.
Remark 4.1 (Disintegration form).
If one prefers a geometric formula, (2) is equivalent to disintegrating $\gamma$ along the map $T_{\tau}$:
$v^{\star}(u,\tau)=\int_{T_{\tau}^{-1}(u)}(u_{1}-u_{0})\,\gamma_{u}^{\tau}(du_{0},du_{1}),\quad T_{\tau}^{-1}(u)=\{(u_{0},u_{1}):(1-\tau)u_{0}+\tau u_{1}=u\}.$
Keeping this as a remark avoids leading with fiber notation.
Rectified-flow training = regression onto constant chord directions. Given samples $(u_{0},u_{1})\sim\gamma$ and $\tau\sim\mathrm{Unif}[0,1]$, form $u_{\tau}=(1-\tau)u_{0}+\tau u_{1}$. Train $v_{\theta}$ by
$\min_{\theta}\;\mathbb{E}\Bigl[\bigl\|\,v_{\theta}(u_{\tau},\tau)-(u_{1}-u_{0})\,\bigr\|^{2}\Bigr].$ (3)
By the $L^{2}$ optimality of conditional expectation, the population minimizer of (3) is exactly $v^{\star}$ in (2). The key qualitative point is immediate: the regression target $(u_{1}-u_{0})$ does not depend on $\tau$. So the learned field is pressured to behave as time-stationary as the data allow along these chords. This is the precise sense in which rectified-flow training promotes “straight” transports without separately adding an ad hoc straightness penalty.
Sampling = integrate the learned transport ODE. Once $v_{\theta}$ approximates $v^{\star}$, sampling amounts to integrating
$\frac{du_{\tau}}{d\tau}=v_{\theta}(u_{\tau},\tau),\quad u_{0}\sim\mu_{0},$ (4)
and returning $u_{1}$ at $\tau=1$. In the idealized chord model where $v^{\star}$ is constant in $\tau$ along typical trajectories, large steps are accurate; deviations from that ideal appear as time-variation of the velocity along the path, which is exactly what discretization error is sensitive to (see Lemma 4.4).
Curvature-aware control: estimate “how non-constant in $\tau$” the velocity is.
During sampling, write $v_{t}:=v_{\theta}(u_{t},\tau_{t})$ and maintain an EMA trend
$v_{t}^{\mathrm{ema}}=\lambda\,v_{t-1}^{\mathrm{ema}}+(1-\lambda)\,v_{t},\quad\lambda\in(0,1).$ (5)
Define the instantaneous straightness/curvature proxy
$s_{t}:=\|v_{t}-v_{t}^{\mathrm{ema}}\|_{2},$
which measures short-horizon time-variation of the velocity along the realized trajectory. Large $s_{t}$ signals local bending / stiffness, precisely the regime where explicit large-step integration incurs higher local truncation error (again, Lemma 4.4).
A one-line regularized update (closed form).
When $s_{t}$ is large, replace $v_{t}$ by a gently regularized velocity $\tilde{v}_{t}$ defined as the minimizer of
$\tilde{v}_{t}\;:=\;\arg\min_{w}\;\|w-v_{t}\|_{2}^{2}\;+\;\eta_{t}\,\|w-v_{t}^{\mathrm{ema}}\|_{2}^{2},$
which yields the explicit convex blend
$\tilde{v}_{t}=(1-\alpha_{t})v_{t}+\alpha_{t}v_{t}^{\mathrm{ema}},\quad\alpha_{t}=\frac{\eta_{t}}{1+\eta_{t}}\in[0,1].$ (6)
Choosing $\eta_{t}$ as a saturating function of $s_{t}$ makes the correction dormant when the path is straight and active only when curvature spikes.
Adaptive step size from the same signal.
Since explicit Euler LTE scales like $(\Delta\tau)^{2}$ times a norm of the time-variation of the vector field (Lemma 4.4), use $s_{t}$ as a scale proxy and set
$\Delta\tau_{t}\;\propto\;\frac{1}{\sqrt{\kappa_{1}s_{t}+\kappa_{2}}},\quad\Delta\tau_{t}\in[\Delta\tau_{\min},\Delta\tau_{\max}],$ (7)
with $\kappa_{2}$ absorbing baseline stiffness (e.g. dependence on $\|v_{t}\|$ and local Lipschitz effects). The state update is then
$u_{t+1}=u_{t}+\Delta\tau_{t}\,\tilde{v}_{t}.$
When the learned transport is locally straight ($s_{t}$ small), the sampler automatically takes large steps and behaves like “chord shooting”; when curvature increases, it both damps the direction and contracts the step, preventing instability while retaining the large-step advantage in the straight regime.
4.2 Error Decomposition and Bounds
We now quantify why (6)–(7) help. Write the ideal and learned flows
$\dot{u}_{\tau}=v_{\star}(u_{\tau},\tau),\quad\dot{\hat{u}}_{\tau}=v_{\theta}(\hat{u}_{\tau},\tau),$
with the same $u_{0}$. Assume $v_{\theta}(\cdot,\tau)$ is $L$-Lipschitz in $u$ and $v_{\star}\in L^{2}$.
Theorem 4.2 (Terminal error decomposition).
Let $U_{\tau}$ solve $\dot{U}_{\tau}=v_{\star}(U_{\tau},\tau)$ with $U_{0}=u_{0}$, and let $\hat{u}_{\tau}$ solve $\dot{\hat{u}}_{\tau}=v_{\theta}(\hat{u}_{\tau},\tau)$ with $\hat{u}_{0}=u_{0}$. Define the time–averaged velocity $\bar{v}_{\theta}(u):=\int_{0}^{1}v_{\theta}(u,\tau)\,d\tau$ and
$\varepsilon_{\mathrm{fit}}^{2}:=\int_{0}^{1}\mathbb{E}\big\|\bar{v}_{\theta}(U_{\tau})-v_{\star}(U_{\tau},\tau)\big\|^{2}\,d\tau,\quad\varepsilon_{\mathrm{curv}}^{2}:=\int_{0}^{1}\mathbb{E}\big\|v_{\theta}(U_{\tau},\tau)-\bar{v}_{\theta}(U_{\tau})\big\|^{2}\,d\tau.$
Then, for the continuous–time flows,
$\mathbb{E}\big\|\hat{u}_{1}-u_{1}\big\|\;\leq\;e^{L}\,\Big(\varepsilon_{\mathrm{fit}}+\varepsilon_{\mathrm{curv}}\Big).$ (8)
Interpretation.
The terminal error splits into a reconstruction term ($\varepsilon_{\mathrm{fit}}$) and a straightness term ($\varepsilon_{\mathrm{curv}}$). Blending (6) shrinks the within-time variation, directly tightening the bound; the adaptive stepping in (7) triggers corrections where the curvature proxy $s_{t}$ spikes.
Corollary 4.3 (High-probability control of terminal error).
Under the same assumptions as Theorem 4.2, suppose in addition that
$\int_{0}^{1}\mathbb{E}\bigl\|v_{\theta}(U_{\tau},\tau)-v_{\star}(U_{\tau},\tau)\bigr\|^{2}\,d\tau\;\leq\;\varepsilon_{\mathrm{fit}}^{2},\quad\int_{0}^{1}\mathbb{E}\bigl\|v_{\theta}(U_{\tau},\tau)-\bar{v}_{\theta}(U_{\tau})\bigr\|^{2}\,d\tau\;\leq\;\varepsilon_{\mathrm{curv}}^{2},$
so that both the reconstruction error and straightness error are controlled in $L^{2}$. Then there exists a constant $C(L)$ such that
$\mathbb{E}\bigl\|\hat{u}_{1}-u_{1}\bigr\|^{2}\;\leq\;C(L)\,\bigl(\varepsilon_{\mathrm{fit}}^{2}+\varepsilon_{\mathrm{curv}}^{2}\bigr).$ (9)
Consequently, for any tolerance $\eta>0$, Chebyshev’s inequality yields
$\mathbb{P}\bigl(\|\hat{u}_{1}-u_{1}\|>\eta\bigr)\;\leq\;\frac{C(L)\,\bigl(\varepsilon_{\mathrm{fit}}^{2}+\varepsilon_{\mathrm{curv}}^{2}\bigr)}{\eta^{2}}.$ (10)
In particular, as training drives $\varepsilon_{\mathrm{fit}},\varepsilon_{\mathrm{curv}}\to 0$, the probability of a large terminal error event decays at least quadratically in the target radius $\eta$.
Proof sketch.
The proof follows the same Grönwall argument as Theorem 4.2, but carried out in $L^{2}$ instead of $L^{1}$, yielding $\mathbb{E}\|\hat{u}_{1}-u_{1}\|^{2}\leq C(L)(\varepsilon_{\mathrm{fit}}^{2}+\varepsilon_{\mathrm{curv}}^{2})$ for some $C(L)\sim e^{2L}$. Chebyshev’s inequality then gives $\mathbb{P}(\|\hat{u}_{1}-u_{1}\|>\eta)\leq\mathbb{E}\|\hat{u}_{1}-u_{1}\|^{2}/\eta^{2}$, which is the claimed bound. A full proof (including the sharp dependence on the straightness functional) is deferred to the theory appendix. ∎
Lemma 4.4 (Local truncation error for explicit Euler on $\dot{u}=v(u,\tau)$).
Assume $v\in C^{1}$ in $(u,\tau)$ on the region traversed and that $\|\partial_{\tau}v(u,\tau)\|\leq M_{\tau}$, $\|J_{u}v(u,\tau)\|\leq L$, and $\|v(u,\tau)\|\leq M$. Then one explicit-Euler step from $(u,\tau)$ with step size $\Delta\tau$ has local truncation error
$\mathrm{LTE}(\tau;\Delta\tau)$ $:=\bigl\|u(\tau+\Delta\tau)-\bigl(u(\tau)+\Delta\tau\,v(u(\tau),\tau)\bigr)\bigr\|$
$=\frac{(\Delta\tau)^{2}}{2}\,\bigl\|\partial_{\tau}v(u,\tau)+J_{u}v(u,\tau)\,v(u,\tau)\bigr\|\;+\;\mathcal{O}\!\bigl((\Delta\tau)^{3}\bigr)$ (11)
$\leq\frac{(\Delta\tau)^{2}}{2}\,(M_{\tau}+LM)\;+\;\mathcal{O}\!\bigl((\Delta\tau)^{3}\bigr).$
From Lemma 4.4 to the step rule.
The leading LTE term depends on $\|\partial_{\tau}v\|$ and $\|J_{u}v\,v\|$. The EMA deviation $s_{t}=\|v_{t}-v_{t}^{\mathrm{ema}}\|$ is a computable proxy for the (local) time variation $\|\partial_{\tau}v\|$ (up to a scale factor tied to the EMA bandwidth), while $L\,\|v_{t}\|$ upper-bounds the spatial contribution. Enforcing $\mathrm{LTE}\lesssim\mathrm{tol}$ yields the square-root rule (7), with $\kappa_{1}$ absorbing the EMA bandwidth and $\kappa_{2}\approx L\|v_{t}\|$. Under blending, $\|\,\tilde{v}_{t}-v_{t}^{\mathrm{ema}}\|=(1-\alpha_{t})s_{t}$, so the curvature term and the required step shrink together, stabilizing integration where needed and enlarging steps elsewhere. See SM E for proofs and additional theoretical context.
4.3 Spectral Capacity, Structure Functions, and Law-Level Coverage
Functional setting.
We work on the periodic box $\mathcal{D}=\mathbb{T}^{d}$ and consider $L^{2}$ fields $\mathcal{H}:=L^{2}(\mathcal{D};\mathbb{R}^{m})$ with norm $\|\cdot\|_{2}$. For incompressible velocity fields (e.g. SL2D, see Section 6), one may equivalently restrict to divergence-free subspaces and adopt the statistical-solution framework of Lanthaler et al. (2021a), which characterizes laws via correlation identities, admissible observables, and an energy inequality. In that framework the natural distance on laws is the $L^{2}$-based Wasserstein metric $W_{2}$, and small-scale regularity is expressed in terms of structure functions and their time-averages. The numerical ensembles used in our experiments fit this setting: they are generated by dissipative solvers on $\mathbb{T}^{2}$, have uniformly bounded energy, and exhibit empirical structure-function scaling (equivalently, power-law spectra) consistent with the assumptions below.
4.3.1 Bandlimited capacity and strain constraints
Any $u\in\mathcal{H}$ admits a Fourier series $u(x)=\sum_{k\in\mathbb{Z}^{d}}\hat{u}(k)e^{ik\cdot x}$. We write $P_{\leq K}$ for the projector onto modes $\{k:|k|\leq K\}$ and $P_{>K}=I-P_{\leq K}$. On a grid of size $N^{d}$, the Nyquist frequency satisfies $K_{\max}\simeq N$, so every numerically represented field is effectively bandlimited to $|k|\lesssim K_{\max}$. Architecturally, downsampling and anti-aliasing may enforce an even stricter cutoff $K_{c}\leq K_{\max}$.
The next lemma is a standard Bernstein-type inequality: bandlimited fields have a uniform Lipschitz cap.
Lemma 4.5 (Bernstein upper bound for bandlimited fields).
Let $u:\mathbb{T}^{d}\to\mathbb{R}^{m}$ satisfy $u=P_{\leq K}u$ for some integer $K\geq 1$. Then there exists a constant $C_{d}>0$ depending only on the dimension $d$ such that
$\|\nabla_{x}u\|_{L^{\infty}(\mathbb{T}^{d})}\;\leq\;C_{d}\,K^{1+\frac{d}{2}}\,\|u\|_{L^{2}(\mathbb{T}^{d})}.$ (12)
In particular, on an energy shell $\|u\|_{L^{2}}\leq E^{1/2}$ one has
$\|\nabla_{x}u\|_{L^{\infty}}\;\leq\;C_{d}\,K^{1+\frac{d}{2}}\,E^{1/2}.$ (13)
Proof sketch.
Write $u(x)=\sum_{|k|\leq K}\hat{u}(k)e^{ik\cdot x}$, so $\nabla_{x}u(x)=\sum_{|k|\leq K}(ik)\hat{u}(k)e^{ik\cdot x}$ and $|\nabla_{x}u(x)|\leq\sum_{|k|\leq K}|k|\,|\hat{u}(k)|$. Cauchy–Schwarz and a lattice-counting bound $\{|k|\leq K\}\lesssim K^{d}$ give $\|\nabla u\|_{L^{\infty}}\lesssim(\sum_{|k|\leq K}|k|^{2})^{1/2}\|\hat{u}\|_{\ell^{2}}\lesssim K^{1+d/2}\|u\|_{L^{2}}$. Full details are deferred to the theory appendix. ∎
Upper bounds alone do not say what happens if a surrogate misses an entire high-frequency annulus. The next lemma gives a complementary lower bound.
Lemma 4.6 (Annulus lower bound).
Let $f=P_{[K,2K]}u$ for some $u\in L^{2}(\mathbb{T}^{d})$, where $P_{[K,2K]}:=P_{\leq 2K}-P_{\leq K}$ and $K\geq 1$. Then there exists $c_{d}>0$ such that
$\|\nabla_{x}f\|_{L^{\infty}}\;\geq\;c_{d}\,K\,\|f\|_{L^{\infty}}\;\geq\;c_{d}\,K^{1-\frac{d}{2}}\,\|f\|_{L^{2}}.$ (14)
Proof sketch.
On the annulus $\{K\leq|k|\leq 2K\}$, the symbol of $\nabla$ has magnitude $|k|\in[K,2K]$. This yields $\|\nabla f\|_{L^{\infty}}\gtrsim K\|f\|_{L^{\infty}}$. A reverse Bernstein estimate on an annulus gives $\|f\|_{L^{\infty}}\gtrsim K^{-d/2}\|f\|_{L^{2}}$. See the theory appendix for a precise statement independent of the Littlewood–Paley cut-off choice. ∎
Corollary 4.7 (Capacity requirement for typical strain levels).
Suppose the target flows at time $t$ have typical energy $E$ and typical rate-of-strain $\|\nabla_{x}u\|_{L^{\infty}}\simeq L_{\star}$. Let a surrogate produce only bandlimited outputs $u=P_{\leq K_{c}}u$ with $\|u\|_{L^{2}}\leq E^{1/2}$. Then necessarily
$C_{d}\,K_{c}^{1+\frac{d}{2}}\,E^{1/2}\;\gtrsim\;L_{\star},$ (15)
and hence a minimal cutoff
$K_{c}\;\gtrsim\;\Bigl(\frac{L_{\star}}{E^{1/2}}\Bigr)^{\frac{2}{2+d}}.$ (16)
Moreover, if the true flow has nontrivial energy in an annulus above $K_{c}$, then any bandlimited approximation underestimates its strain by at least $c_{d}K^{1-d/2}\|P_{[K,2K]}u\|_{L^{2}}$ for some $K>K_{c}$.
Remark 4.8.
Corollary 4.7 formalises the intuition that spatial capacity ($K_{c}$) is a hard constraint: at fixed energy scale $E$, there is a minimum cutoff needed to represent typical strain levels $L_{\star}$. In our experiments both GenCFD and ReFlow operate at the same grid resolution, so they share the same effective $K_{c}$; rectification does not invent new small scales but makes the trajectories inside this bandlimited class straight and easy to integrate.
4.3.2 Structure functions and small-scale coverage
We now quantify how much of the law of the PDE solution can be captured at a given cutoff $K_{c}$. For $u:\mathbb{T}^{d}\to\mathbb{R}^{m}$ and $r>0$ define the second-order structure function
$S_{r}^{2}(u):=\frac{1}{|B_{r}|}\int_{|h|\leq r}\|u(\cdot+h)-u(\cdot)\|_{L^{2}(\mathbb{T}^{d})}^{2}\,dh,$
and for a probability law $\mu$ on $\mathcal{H}$ set
$S_{r}^{2}(\mu):=\int S_{r}^{2}(u)\,d\mu(u).$
In homogeneous turbulence $S_{r}^{2}(\mu)$ is directly related to the energy spectrum; power-law behaviour $S_{r}^{2}(\mu)\sim r^{\zeta_{2}}$ at small $r$ encodes the scaling of the cascade. In our experiments, Figure 13 visualizes this connection directly: matching the high-wavenumber tail in the spectrum is equivalent to capturing the small-$r$ behaviour of structure functions, and therefore to controlling the coverage term below.
Guided by the statistical-solution framework of Fjordholm et al. (2017), we assume there exists a modulus of continuity $\omega:(0,1]\to[0,\infty)$ with $\omega(r)\downarrow 0$ as $r\downarrow 0$ such that
$S_{r}^{2}(\mu_{t})\;\leq\;\omega(r)\quad\text{for all }r\in(0,1],\;t\in[0,T].$ (17)
In practice, $\omega$ can be taken as a piecewise power law $Cr^{2\alpha}$ calibrated from the empirical spectra of the numerical ensembles. The next lemma shows that (17) controls the high-frequency tail of the law.
Lemma 4.9 (Structure function controls Fourier tail).
Let $\mu$ be a law on $L^{2}(\mathbb{T}^{d})$ satisfying $S_{r}^{2}(\mu)\leq\omega(r)$ for all $r\in(0,1]$, where $\omega(r)\downarrow 0$ as $r\downarrow 0$. Then there exist constants $c_{d},C_{d}>0$ such that for all integers $K\geq 1$,
$\int\|P_{>K}u\|_{L^{2}}^{2}\,d\mu(u)\;\leq\;C_{d}\,\omega\!\bigl(c_{d}/K\bigr).$ (18)
Proof sketch.
Expand $u$ in Fourier series and observe $\|u(\cdot+h)-u(\cdot)\|_{L^{2}}^{2}=\sum_{k}|\hat{u}(k)|^{2}|e^{ik\cdot h}-1|^{2}$. Averaging over $h\in B_{r}$ yields $S_{r}^{2}(u)=\sum_{k}|\hat{u}(k)|^{2}\Psi_{r}(k)$, where $\Psi_{r}(k)\gtrsim\min\{(|k|r)^{2},1\}$ uniformly in $k,r$. For $|k|\gtrsim 1/r$, one has $\Psi_{r}(k)\gtrsim 1$, hence $S_{r}^{2}(u)\gtrsim\|P_{>\,c/r}u\|_{L^{2}}^{2}$. Integrate over $u\sim\mu$, apply (17), and set $r\sim 1/K$. Full constants and the Littlewood–Paley formulation are deferred to the theory appendix. ∎
Lemma 4.10 (Bandlimited projection in $W_{2}$).
For any law $\mu$ on $L^{2}(\mathbb{T}^{d})$ and cutoff $K_{c}\geq 1$,
$W_{2}\!\bigl(\mu,\,(P_{\leq K_{c}})_{\#}\mu\bigr)\;\leq\;\Bigl(\int\|P_{>K_{c}}u\|_{L^{2}}^{2}\,d\mu(u)\Bigr)^{1/2}.$ (19)
Proof.
Couple $u\sim\mu$ and $\tilde{u}=P_{\leq K_{c}}u$. Then $\tilde{u}\sim(P_{\leq K_{c}})_{\#}\mu$ and $W_{2}^{2}(\mu,(P_{\leq K_{c}})_{\#}\mu)\leq\mathbb{E}\|u-\tilde{u}\|_{L^{2}}^{2}=\mathbb{E}\|P_{>K_{c}}u\|_{L^{2}}^{2}$. ∎
Combining Lemmas 4.9 and 4.10 gives a clean coverage bound.
Corollary 4.11 (Coverage term from the structure-function modulus).
Let $\mu_{t}$ be the statistical solution law at time $t$ and assume (17). Then for any cutoff $K_{c}\geq 1$,
$W_{2}\!\bigl(\mu_{t},\,(P_{\leq K_{c}})_{\#}\mu_{t}\bigr)\;\lesssim\;\omega\!\bigl(c_{d}/K_{c}\bigr)^{1/2},$ (20)
with $c_{d}$ as in Lemma 4.9.
4.3.3 Capacity–coverage–fit decomposition for one-step RF
Finally, we combine the coverage term with the training error of a one-step rectified-flow surrogate.
Let $S_{\Delta t}:\mathcal{H}\to\mathcal{H}$ denote the PDE solution operator over a physical step $\Delta t$, and write
$\mu_{t+\Delta t}:=(S_{\Delta t})_{\#}\mu_{t}.$
Assume the rectified-flow propagator $\mathcal{T}^{\theta}_{\Delta t}$ aims at the bandlimited target $P_{\leq K_{c}}(S_{\Delta t}u)$ in mean-square. Concretely, sample $u\sim\mu_{t}$ and an internal noise seed $\xi\sim\mathcal{N}(0,I)$, and define the RF output $\widehat{u}:=\mathcal{T}^{\theta}_{\Delta t}(u,\xi)$. We measure the training fit by
$\varepsilon_{\mathrm{train}}^{2}(t):=\mathbb{E}_{u\sim\mu_{t}}\,\mathbb{E}_{\xi\sim\mathcal{N}(0,I)}\bigl\|P_{\leq K_{c}}(S_{\Delta t}u)-\widehat{u}\bigr\|_{L^{2}}^{2}.$ (21)
Finally, we define the one-step RF law by
[…]
Proposition 4.12 (One-step law-level error: capacity, coverage, and fit).
Under the structure-function modulus assumption (17), the one-step $W_{2}$ error between the PDE law and the RF law satisfies
$W_{2}\!\bigl(\mu_{t+\Delta t},\,\widehat{\mu}_{t+\Delta t}\bigr)\;\leq\;\Bigl(C_{d}\,\omega\!\bigl(c_{d}/K_{c}\bigr)\;+\;\varepsilon_{\mathrm{train}}^{2}(t)\Bigr)^{1/2},$ (22)
where the first term is a coverage error determined by the structure-function modulus and the cutoff $K_{c}$, and the second is a fit error determined by training.
Proof sketch.
Couple $u\sim\mu_{t}$ and $\widehat{u}$ using the same input; then $W_{2}^{2}(\mu_{t+\Delta t},\widehat{\mu}_{t+\Delta t})\leq\mathbb{E}\|S_{\Delta t}u-\widehat{u}\|_{L^{2}}^{2}$. Insert $P_{\leq K_{c}}S_{\Delta t}u$, apply the triangle inequality, and bound the tail term by Lemma 4.9 (applied at time $t+\Delta t$). ∎
Interpretation and experimental validation.
At a fixed resolution (fixed $K_{c}$), Proposition 4.12 says that even a perfectly trained rectified flow cannot beat the coverage term $\omega(c_{d}/K_{c})^{1/2}$ dictated by small-scale physics. This is precisely why spectrum matching (Figure 13) is a non-negotiable diagnostic on multiscale flows: if the surrogate loses the high-frequency tail, it is necessarily paying a coverage penalty in law. The remaining error is controlled by the training fit and the straightness properties of the learned velocity (via the continuous-time bounds in Section 4.1). Rectified flows improve efficiency precisely by reducing the curvature term in these bounds, so that a given coverage+fit accuracy can be reached with many fewer ODE steps than a diffusion-based surrogate such as GenCFD (Section 5.6).
5 Correctness and Curvature-Controlled Discretization
Organization.
This section states the main correctness and discretization results that underpin the narrative above. Full proofs are deferred to the theory appendix (to be written separately); we keep only the steps needed to make the dependence on fit, straightness, and discretization explicit in the main text.
5.1 Exact transport under the barycentric rectified velocity
Proposition 5.1 (Exact transport under the barycentric velocity).
Let $\mu_{0},\mu_{1}\in\mathcal{P}(\mathcal{U})$ and let $\gamma\in\mathcal{P}(\mathcal{U}\times\mathcal{U})$ be a coupling of $(\mu_{0},\mu_{1})$. Define $T_{\tau}(u_{0},u_{1})=(1-\tau)u_{0}+\tau u_{1}$ and $\rho_{\tau}:=(T_{\tau})_{\#}\gamma$. Let $v_{\star}(\cdot,\tau)$ be the barycentric velocity field of defined in 2. Assume (e.g. on a finite-dimensional Galerkin discretization of $\mathcal{U}$) that the continuity equation
$\partial_{\tau}\rho_{\tau}+\nabla\!\cdot(\rho_{\tau}\,v_{\star}(\cdot,\tau))=0$ (23)
admits a unique regular Lagrangian flow $X_{\tau}$ associated to $\dot{u}_{\tau}=v_{\star}(u_{\tau},\tau)$. Then $(X_{1})_{\#}\mu_{0}=\mu_{1}$.
Proof sketch.
By construction, $\rho_{\tau}$ is the law of $U_{\tau}=(1-\tau)U_{0}+\tau U_{1}$ for $(U_{0},U_{1})\sim\gamma$. Differentiate $\mathbb{E}[\varphi(U_{\tau})]$ and condition on $U_{\tau}$ to identify the weak form of (23). Uniqueness of the continuity equation solution (in the chosen Galerkin setting) implies $\rho_{\tau}=(X_{\tau})_{\#}\rho_{0}$, hence $\rho_{1}=(X_{1})_{\#}\rho_{0}$, i.e. $\mu_{1}=(X_{1})_{\#}\mu_{0}$. ∎
Why this matters. Proposition 5.1 isolates a clean correctness statement: if $v_{\theta}$ matches the barycentric rectified velocity $v_{\star}$ sufficiently well (in a sense quantified by Theorem 4.2), the induced flow pushes forward $\mu_{0}$ to $\mu_{1}$ at the level of laws.
5.2 Curvature controls the required number of ODE steps
Corollary 5.2 (Global Euler error controlled by time-curvature).
Assume the hypotheses of Lemma 4.4 for $\dot{u}=v(u,\tau)$ on $\tau\in[0,1]$ and assume $v(\cdot,\tau)$ is $L$-Lipschitz in $u$ uniformly in $\tau$. Let $u^{\Delta}$ be the explicit-Euler approximation with uniform step $\Delta\tau=1/N$. Then there exists a constant $C=C(L)$ such that
$\sup_{\tau\in[0,1]}\|u^{\Delta}_{\tau}-u_{\tau}\|\;\leq\;C\,\Delta\tau\,\int_{0}^{1}\Big(\|\partial_{\tau}v(u_{\tau},\tau)\|+\|J_{u}v(u_{\tau},\tau)\,v(u_{\tau},\tau)\|\Big)\,d\tau.$ (24)
In particular, when $\int_{0}^{1}\|\partial_{\tau}v(u_{\tau},\tau)\|\,d\tau$ is small (“straight” flow), the $N$ required to reach a fixed tolerance is proportionally smaller.
Proof sketch.
Sum the local truncation errors from Lemma 4.4 over $N$ steps and propagate them using stability of the flow under the $L$-Lipschitz condition via Grönwall. ∎
Connection to our curvature proxy. The EMA deviation $s_{t}=\|v_{t}-v_{t}^{\mathrm{ema}}\|$ used in (6)–(7) is a computable proxy for the same time-variation that drives (24). Blending reduces effective curvature; step control reduces $\Delta\tau$ when curvature spikes. This directly rationalizes why ReFlow can reach the “spectrum-correct” regime in very few steps (Figure 13), and why the same accuracy level is reached at $8$ steps (Table 1) whereas diffusion baselines typically need $\geq 128$.
5.3 A compact regularity interpretation of structure functions
Lemma 5.3 (Structure functions imply Besov regularity (heuristic equivalence)).
Let $u\in L^{2}(\mathbb{T}^{d})$ and define the increment seminorm
$u]_{B^{s}_{2,\infty}}:=\sup_{0<|h|\leq 1}\frac{\|u(\cdot+h)-u(\cdot)\|_{L^{2}}}{|h|^{s}}.$
If the second-order structure function satisfies $S_{r}^{2}(u)\lesssim r^{2s}$ for all $r\in(0,1]$, then $u]_{B^{s}_{2,\infty}}\lesssim 1$. Consequently, a law-level modulus bound $S_{r}^{2}(\mu)\lesssim r^{2s}$ implies that typical samples from $\mu$ have Besov smoothness $s$, consistent with the Fourier-tail decay in Lemma 4.9.
Remark 5.4.
In turbulence one often observes $S_{r}^{2}(\mu)\sim r^{\zeta_{2}}$ over an inertial range. Heuristically this corresponds to $s\approx\zeta_{2}/2$, so the coverage term in Corollary 4.11 can be read as a direct proxy for inertial-range scaling.
5.4 Master inequality
Theorem 5.5 (Law-level error: coverage + fit + straightness + discretization).
Fix an effective resolution cutoff $K_{c}$ and assume the structure-function modulus hypothesis used in Corollary 4.11. Let $\hat{\mu}_{1}^{(N)}$ be the law obtained by taking $N$ explicit-Euler steps to integrate the learned rectified ODE $\dot{u}_{\tau}=v_{\theta}(u_{\tau},\tau)$ from $\tau=0$ to $\tau=1$. Assume $v_{\theta}(\cdot,\tau)$ is $L$-Lipschitz in $u$ uniformly in $\tau$. Then, for constants depending on $L$,
[…]
Proof sketch.
The coverage term is Corollary˜4.11. The continuous-time mismatch between the ideal rectified flow and the learned flow is controlled by Theorem 4.2 via $\varepsilon_{\mathrm{fit}}$ and $\varepsilon_{\mathrm{curv}}$. The discretization term follows from Corollary 5.2 applied to the learned ODE. ∎
Reading (25) against the experiments.
Equation (25) matches the empirical breakdown seen in Section 6: (i) coverage corresponds to reproducing the high-wavenumber energy content (Figure 13); (ii) fit corresponds to the overall statistical accuracy, reflected in $e_{\mu}$, $e_{\sigma}$, and $W_{1}$ (Table 1); (iii) straightness and discretization are exactly what turn “a good model” into “a fast sampler”: ReFlow reaches the same law-level quality using N=8N\!=\!8 steps, whereas diffusion baselines require $N\!\geq\!128$ (Section 5.6).
5.5 Implementation Details for PDEs
Learning velocity fields in PDE settings. When rectified flows are used to model PDE solutions, we assume access to paired initial/final states $(u_{i},u_{f})$ where $u_{f}=\mathcal{S}(u_{i})$ represents the evolved initial condition $u_{i}$, at some final physical time (e.g. $t=1$). We train a neural network to predict the velocity $v_{\theta}$ that displaces $u_{i}$ towards $u_{f}$, with the additional mention that $v_{\theta}$ shall be conditioned on $u_{i}$ at all diffusion times.
Algorithm 1 Rectified Flow Training (PDE Context)
1:Dataset $\{(u_{0},u_{1})\}$ with $u_{1}=\mathcal{S}(u_{0})$; velocity model $v_{\theta}$; noise schedule $\sigma(\tau)$
2:for each iteration do
3:  Sample a pair $(u_{0},u_{1})$
4:  Sample $\tau\sim\mathcal{U}(0,1)$ and noise $\xi\sim\mathcal{N}(0,I)$
5:  Form $u_{\tau}=\tau\cdot u_{1}+\sigma(1-\tau)\,\xi$
6:  Compute loss:
$\mathcal{L}(\theta)\;=\;\bigl\|\bigl(u_{1}-u_{0}\bigr)\;-\;v_{\theta}\bigl(u_{\tau},\,u_{0},\,\tau\bigr)\bigr\|^{2}$
7:  Update $\theta$ via gradient descent
8:end for
Overview of the model. Our network learns a velocity (or denoising) map
$v_{\theta}(u_{\tau},u_{0},\,\tau)\;=\;\mathrm{UNet}_{\theta}\bigl(\,u_{\tau},u_{0},\;\Gamma(\tau)\bigr),$
where $u_{\tau}$ is a low-frequency representation of $u_{1}$, $u_{0}$ is the initial datum, and $\Gamma(\cdot)$ is an MLP-based time embedding:
$\Gamma(\tau)\;=\;\mathrm{MLP}\bigl(\mathrm{Sinusoid}(\tau)\bigr).$
Within the U-Net (see SM J.1), we alternate residual blocks with multi‐head attention layers, downsampling to a bottleneck and then upsampling via skip connections. A final $1\times 1$ projection recovers either the field increment $\hat{u}\approx u_{1}-u_{0}$ or a noise residual. By injecting the rectification time $\Gamma(\tau)$ into each block, the model performs a fully deterministic trajectory integration instead of stochastic diffusion.
Sampling procedure. Once trained, we generate new samples by numerically integrating the ODE
$\frac{du_{\tau}}{d\tau}=v_{\theta}\bigl(u_{\tau},\,u_{0},\,\tau\bigr),\quad u_{\tau}|_{\tau=0}=\xi,\quad\xi\sim\mathcal{N}(0,I).$
Any standard ODE solver (e.g. forward Euler, Runge-Kutta) can be used; rectified flows often tolerate relatively large step sizes due to their straighter velocity field.
Efficiency and accuracy. Because rectified flows explicitly encourage a “straighter” velocity field, accurate solutions often require fewer integration steps than their diffusion-based counterparts. This advantage can be crucial in computational fluid dynamics and other PDE applications where each solver step is expensive. See SM M.3 and P for a detailed comparison of step-count vs. accuracy trade-offs.
5.6 Computational Efficiency Comparison
Diffusion surrogates (e.g. GenCFD) generate samples by integrating a reverse-time SDE whose trajectory bends through latent noise; even with accelerated samplers such as DDPM or DPM-solver the procedure still requires tens of score-network evaluations. Rectified flows, in contrast, follow a single deterministic ODE whose nearly straight path tolerates large or adaptive steps, so inference needs only a handful of evaluations. Figure 2 quantifies this gap, showing how the runtime of diffusion models scales almost linearly with the number of steps, whereas rectified flows remain fast even as network size grows.
Figure 2: Inference cost versus model size. Rectified flows need 4488 ODE steps vs. GenCFD’s 6464128128 SDE steps. U/̄Net depth varies from $4$ to 88 blocks (parameters $5.5$ M to $10.2$ M).
6 Experiments
6.1 Experimental Setup
We benchmark our method, termed henceforth as ReFlow, against four neural surrogate models. GenCFD and GenCFD $\circ$ FNO are conditional diffusion surrogates following the framework introduced earlier in this section, whereas FNO and UViT serve as one-shot deterministic baselines. Detailed descriptions of all baselines are provided in SM L.
Datasets for Training and Evaluation.
We consider three canonical 2D multi-scale PDE problems (see the Supplementary Material for precise mathematical formulations and solver details):
  • SL2D
    Shear Layer in 2D. This setup models the evolution of an incompressible shear layer, initially concentrated near a horizontal interface, which rolls up into intricate vortical structures over time. We generate approximate solutions by numerically solving the 2D Navier–Stokes system (with periodic boundaries) at a resolution of $512^{2}$, then downsample to $128^{2}$ for training and testing. The data has been generated with the spectral viscosity code Tadmor (1989); Rohner and Mishra (2024).
  • CS2D
    Cloud-Shock Interaction in 2D. In this popular compressible-flow test, a shock wave moves through a higher-density “cloud” region, triggering complex wave interactions and small-scale turbulent mixing in the wake. We solve the compressible Euler system at an effective resolution of $256^{2}$ using a GPU-optimized finite-volume ALSVINN code of Lye (2020); Fjordholm et al. (2020), then also downsample to $128^{2}$.
  • RM2D
    Richtmeyer–Meshkov in 2D. In this case, a shock wave interacts with a density interface separating two fluids, depositing vorticity and triggering vigorous mixing. We numerically solve the compressible Euler system on a $256^{2}$ grid, then downsample to $128^{2}$ to create the training and test data. The data has been generated with a high-resolution finite-volume ALSVINN code of Fjordholm et al. (2020)
Training and Test Protocol.
We train both our proposed rectified flow-based GenCFD-class algorithm and all baseline models on initial conditions
$u_{0}\,\sim\,\mu,\quad\text{with}\quad\mu$
being the probability distribution specified in the Supplementary Material (SMK.2). For each $u_{0}$, the PDE solver (e.g. a spectral or finite-volume code) is run forward in time to produce high-fidelity snapshots, which serve as training data.
Training-objective. Crucially, during training, the model is only exposed to $(u_{0},u_{1})$ pairs of data as opposed to the conditional measure $p(u(t)|u_{0})$.
Test-time goal: Dirac initial measures. While the training distribution $\mu$ is spread out, we are especially interested in the scenario where the initial condition is fixed and deterministic. Formally, we set
$\widehat{\mu}_{0}=\delta_{\widehat{u}_{0}},$
which is a Dirac measure concentrated at a single function $\widehat{u}_{0}$. Thus, any stochasticity in the resulting flow arises solely from chaotic or multi-scale evolution over time, rather than from large variations in the initial condition itself.
To create a “ground truth” ensemble for comparison under a fixed $\widehat{u}_{0}$, we employ an ensemble perturbation procedure (detailed in SM  K.2), where $\widehat{u}_{0}$ is slightly perturbed to generate multiple nearby initial states. Each perturbed state is then advanced using a high-resolution PDE solver (e.g. the Azeban code for incompressible flows, or a finite-volume code for compressible problems), yielding a set of flow realizations. This set approximates the conditional measure in question, i.e. $p\!\bigl(u(t)\mid\widehat{u}_{0}\bigr)$.
Distribution settings. We test two main configurations:
  1. 1.
    Matched distribution. We sample $\widehat{u}_{0}$ from the same family $\mu$ used in training. While $\delta_{\widehat{u}_{0}}$ is technically out-of-distribution in the sense that the training distribution was spread out over many initial states, the base function $\widehat{u}_{0}$ still belongs to the same general class of flows.
  2. 2.
    Distribution shift. We consider Dirac initial data $\delta_{\widehat{u}_{0}}$ where $\widehat{u}_{0}\sim\nu$, for a different distribution $\nu\neq\mu$. This scenario probes the ability of our model and baselines to generalize under shifts in the initial condition distribution. Specific details on $\nu$ are provided in SM K.2.
We will show results in both regimes to assess how well rectified flows can capture the spread of chaotic solutions starting from fixed initial states–and how it compares to baselines.
6.2 Evaluation Metrics
We evaluate the generative performance of each method using three main metrics, with details provided in SM M:
  • Mean Error. We measure the $L^{2}$ difference between the model’s predicted mean flow $\mu$ and the ground-truth mean $\mu_{\mathrm{exact}}$. A low mean error indicates accurate large-scale behavior.
  • Standard Deviation Error. The normalized $L^{2}$ distance between the predicted standard deviation $\sigma$ and the reference $\sigma_{\mathrm{exact}}$ captures whether the model reproduces correct spatial variability.
  • One-Point Wasserstein Distance. We compute an Wasserstein-$1$ distance at each spatial point (and average over the domain) to quantify local distributional agreement between ground truth and model outputs.
One-point Wasserstein as a k=1k\!=\!1 marginal metric.
Let $\mu$ and $\widehat{\mu}$ be laws on fields $u:D\to\mathbb{R}^{m}$. For each $x\in D$ define $\mu_{x}:=(\mathrm{ev}_{x})_{\#}\mu$ and $\widehat{\mu}_{x}:=(\mathrm{ev}_{x})_{\#}\widehat{\mu}$. Our reported metric is
$W_{1}^{(1)}(\mu,\widehat{\mu}):=\int_{D}W_{1}(\mu_{x},\widehat{\mu}_{x})\,dx.$
By Kantorovich–Rubinstein duality, $W_{1}(\mu_{x},\widehat{\mu}_{x})=\sup_{\operatorname{Lip}(\varphi)\leq 1}\big|\int\varphi\,d\mu_{x}-\int\varphi\,d\widehat{\mu}_{x}\big|$, so $W_{1}^{(1)}$ controls the averaged discrepancy of all pointwise 1-Lipschitz observables. This is the precise sense in which Table 1 measures agreement of k=1k\!=\!1 correlation statistics, in line with the statistical-solution viewpoint Lanthaler et al. (2021a).
Table 1: Comparison of models across datasets CS2D, SL2D, and RM2D, in terms of mean relative error ($e_{\mu}$), standard deviation relative error ($e_{\sigma}$), and Wasserstein-1 distance ($W_{1}$). Best-performing model is colored in Blue and second-best performing model in Orange.
Model Metric CS2D SL2D RM2D
$\rho$ $m_{x}$ $m_{y}$ $E$ $u_{x}$ $u_{y}$ $\rho$ $m_{x}$ $m_{y}$ $E$
ReFlow $e_{\mu}$ 0.0477 0.0332 0.041 0.015 0.034 0.189 0.018 0.032 0.020 0.021
$e_{\sigma}$ 0.1005 0.072 0.078 0.0584 0.071 0.077 0.032 0.046 0.046 0.085
$W_{1}$ 0.0091 0.0107 0.0124 0.0143 0.0214 0.0164 0.0033 0.0036 0.0028 0.0017
GenCFD $e_{\mu}$ 0.0830 0.0619 0.127 0.0280 0.039 0.267 0.0049 0.0010 0.0025 0.0025
$e_{\sigma}$ 0.195 0.197 0.244 0.169 0.092 0.072 0.0351 0.0254 0.0435 0.0435
$W_{1}$ 0.0151 0.0184 0.0185 0.0247 0.0276 0.0213 0.0092 0.0016 0.0049 0.0049
UViT $e_{\mu}$ 0.111 0.078 0.127 0.0479 0.233 0.623 0.070 0.070 0.071 0.052
$e_{\sigma}$ 0.415 0.343 0.412 0.296 0.28 0.178 0.425 0.570 0.570 0.617
$W_{1}$ 0.087 0.085 0.158 0.172 0.270 0.256 0.060 0.055 0.060 0.036
FNO $e_{\mu}$ 0.0989 0.0748 0.0930 0.0383 0.058 0.408 0.085 0.070 0.063 0.067
$e_{\sigma}$ 0.2573 0.2779 0.3188 0.2777 0.137 0.110 0.366 0.443 0.436 0.450
$W_{1}$ 0.1455 0.1660 0.1879 0.1884 0.184 0.150 0.096 0.051 0.056 0.062
[…]
$e_{\sigma}$ 0.1099 0.0969 0.1245 0.0974 0.131 0.100 0.091 0.075 0.091 0.090
$W_{1}$ 0.0467 0.0534 0.0774 0.0743 0.166 0.135 0.0269 0.0158 0.0293 0.0273
Baselines and Evaluation Settings. To stay within reasonable training times for our infrastructure, we restrict the total number of training iterations up to 120,000 for each model and recorded the best achievable results. Under this budget, only a reduced-capacity GenCFD (with two downsampling stages in its U-Net backbone, versus three in Rectified Flow) can match the accuracy reported in Table 1. We also evaluated larger GenCFD variants trained for 500,000 iterations, but observed no meaningful improvement over the results shown here. Meanwhile, at inference time Rectified Flow requires only 8 denoising steps to achieve comparable or superior accuracy, whereas diffusion baselines (including GenCFD) need 128 steps, underscoring RF’s efficiency in both training and sampling. FNO and UViT also require smaller models than ReFlow for optimal performance.
Since FNO, and UViT produce deterministic predictions, we approximate their spread by an ensemble perturbation strategy (SM  K.2), where each baseline is fed slightly perturbed inputs and outputs are aggregated to form a sample set. In contrast, both the original GenCFD (baseline) and our ReFlow are inherently probabilistic, requiring no auxiliary procedure to generate diverse samples.
Figure 3: Energy-spectrum evolution for a random cloud–shock initial datum. ReFlow (blue) tracks high-wavenumber energy faster and more accurately than GenCFD (red).
6.3 Results.
[…]
Moreover as visualized in the SM, ReFlow generates samples of very high quality, matching the qualitative features of the ground truth and capturing features such as (interacting) vortices as well as propagating shock waves very accurately. Another aspect where ReFlow really shines in being able to generate the correct point pdfs. Although already indicated in the very low 1-point Wasserstein distance errors from Table 1, we present 1-point pdfs in the SM to illustrate this observation.
Another key indicator of the quality of approximation of multiscale PDE solutions is the energy spectrum Molinaro et al. (2024). In Figure 13, we plot the (log) spectra for the energy variable of a randomly chosen Cloud Shock initial datum at four different Diffusion times, $\tau\in\{0.25,\,0.50,\,0.75,\,1.00\}$ to observe that i) the solution is indeed highly multiscale, with spectra decaying as a power law over a range of frequencies ii) ReFloW approximates the spectra accurately within a few (5) steps and iii) on the other hand, for these few steps in solving the underlying reverse SDE, a score-based Diffusion model such as GenCFD, struggles to approximate the relevant small scales. It will require much more sampling steps to do so. These results also showcase how ReFlow is able to infer the correct statistical solutions with significantly less ODE solve steps and is hence, much faster, than a diffusion based model such as GenCFD. This gain in inference speed is also reinforced from Figure 4, where we compare ReFlow (RF) and GenCFD reconstructions of the density field $\rho$ for the Cloud-Shock interaction over five diffusion times $\tau\in\{0.00,\,0.25,\,0.50,\,0.75,\,1.00\}$. We see that ReFlow inpaints fine-scale features in far fewer steps, whereas GenCFD remains visibly much more noisy at these small number of steps. See SM M.3 for further visualizations of this effect.
Figure 4: Density trajectories for Sample 1: ReFlow (top) versus GenCFD (bottom).
7 Additional empirical analysis (integrators, robustness, and diversity)
This section consolidates additional experiments conducted in response to reviewer requests, with the goal of making the computational trade-offs and the role of the curvature-aware controller explicit. The emphasis is on (i) solver choices for the learned rectified-flow ODE, (ii) robustness of the controller to its hyperparameters, and (iii) distributional fidelity / diversity under the macro–micro evaluation protocol.
7.1 Protocol recap: macro–micro ensembles, OOD macros, and metrics
Macro–micro conditional ensembles.
To evaluate statistical fidelity rather than only pointwise accuracy, we follow a macro–micro protocol. For each dataset, we select $N_{\mathrm{macro}}$ macroscopic states (“macros”) and for each macro generate $N_{\mathrm{micro}}$ microscopic perturbations (“micros”) yielding a reference ensemble of size $N_{\mathrm{macro}}\times N_{\mathrm{micro}}$. This makes diversity concrete: for each fixed macro, the ground-truth solver produces an empirical conditional distribution over micros, and the model produces its own empirical conditional distribution. We compare these distributions directly via Wasserstein distances (below), which is the relevant question for statistical surrogates.
OOD macro selection.
Unless otherwise stated, we report results in an out-of-distribution (OOD) regime by choosing $N_{\mathrm{macro}}=10$ macros from shifted initial-condition families / parameter ranges, and $N_{\mathrm{micro}}=20$ micros per macro (200 conditional samples per dataset). All models are frozen; only the sampling/integration procedure is varied.
Metrics.
We report: (i) relative $L^{2}$ error against the PDE solution,
$\mathrm{Rel}\text{-}L^{2}(\widehat{u},u):=\frac{\|\widehat{u}-u\|_{L^{2}}}{\|u\|_{L^{2}}},$ (26)
(ii) NFE (neural function evaluations), i.e. the number of velocity-field evaluations per generated sample, and (iii) a combined cost–accuracy scalar
$\mathrm{Cost}\times\mathrm{Err}:=(\mathrm{Rel}\text{-}L^{2})\times(\mathrm{NFE}),$ (27)
which is convenient for summarizing the NFE–error plane (lower is better). Distributional fidelity is measured via empirical $W_{1}$ between conditional ensembles; we report both aggregated and per-variable $W_{1}$ (Section 7.5).
7.2 Integrator study at sampling time: fixed-step baselines vs. curvature-aware control
A recurring concern is whether the gains are merely a consequence of choosing a particular numerical solver. To isolate this, we fix the trained rectified-flow model and vary only the integrator used to solve the rectified-flow ODE at sampling time, starting from noise and integrating from rectified time $\tau=0$ to $\tau=1$. We include a reference implementation from torchdiffeq and standard explicit fixed-step schemes. We report the aggregated results on ShearLayer2D in the OOD macro–micro regime; the second benchmark exhibits the same qualitative trends and is reported in the appendix.
[…]
Model-Baseline (built-in sampler) 32.0 0.2533 $\pm$ 0.0382 8.11
torchdiffeq Midpoint 32.0 0.2533 $\pm$ 0.0382 8.11
Euler (16 steps) 16.0 0.2489 $\pm$ 0.0371 3.98
RK2 (16 steps; 32 NFEs) 32.0 0.2533 $\pm$ 0.0379 8.11
RK4 (16 steps; 64 NFEs) 64.0 0.2540 $\pm$ 0.0381 16.26
Observation: Euler has a tuned “sweet spot” but remains brittle.
[…]
7.3 Curvature-aware sampler: hyperparameter robustness and Pareto dominance
We next evaluate the curvature-based adaptive controller under a broad sweep of hyperparameters while holding the model fixed. The controller uses (i) an EMA-based straightness/curvature proxy on the learned velocity field, (ii) a mild velocity regularization/blending step triggered by curvature, and (iii) adaptive step sizing. We vary EMA time scale, thresholds, gating, damping, and calibration parameters across 20 configurations.
Table 3: Hyperparameter robustness of the curvature-aware adaptive sampler (ShearLayer2D, OOD macro–micro). Across 20 controller configurations, Avg. NFE stays in a narrow band and relative errors vary by $1\%$, indicating the controller is not brittle and does not rely on a narrow tuning “sweet spot”.
[…]
baseline 11.0 0.2374 $\pm$ 0.0389 2.62
ema_0.25 10.7 0.2369 $\pm$ 0.0386 2.54
ema_0.45 11.3 0.2361 $\pm$ 0.0384 2.67
alpha_0.05 11.0 0.2373 $\pm$ 0.0389 2.62
alpha_0.12 11.0 0.2373 $\pm$ 0.0389 2.62
alpha_0.20 11.0 0.2374 $\pm$ 0.0389 2.62
gamma_1.5 11.0 0.2374 $\pm$ 0.0389 2.62
gamma_2.0 11.0 0.2374 $\pm$ 0.0389 2.62
gamma_2.5 11.0 0.2374 $\pm$ 0.0389 2.62
gate_0.50 11.0 0.2374 $\pm$ 0.0389 2.62
gate_0.70 11.0 0.2374 $\pm$ 0.0389 2.62
calib_0.70_0.90 11.1 0.2371 $\pm$ 0.0387 2.63
calib_0.80_0.98 10.9 0.2372 $\pm$ 0.0387 2.60
adapt_1.5_0.75 12.8 0.2374 $\pm$ 0.0378 3.04
adapt_2.0_0.85 11.2 0.2360 $\pm$ 0.0376 2.63
no_ortho_filter 11.0 0.2374 $\pm$ 0.0389 2.62
damp_0.05 11.0 0.2373 $\pm$ 0.0389 2.62
damp_0.10 11.0 0.2373 $\pm$ 0.0389 2.61
damp_0.20 11.0 0.2370 $\pm$ 0.0389 2.60
damp_0.30 11.0 0.2368 $\pm$ 0.0389 2.60
Observation: robust controller, automatic step allocation.
Across all 20 configurations, Avg. NFE remains in a tight range (10.7–12.8) and the mean relative error stays within $[0.2360,0.2374]$ (sub-$1\%$ relative variation), indicating that the controller does not require fragile tuning. This directly contrasts with fixed-step Euler, where efficiency is achieved only at a tuned step budget and can degrade rapidly outside it.
7.4 Best fixed-step baseline vs. best adaptive configurations
For clarity, Table 4 compares the strongest fixed-step baseline (Euler-16) against the top adaptive configurations on ShearLayer2D OOD.
[…]
Euler (16 steps) 16.0 0.2489 3.98 0% 0% 0%
ema_0.25 10.7 0.2369 2.54 $33.1\%$ $4.8\%$ $36.2\%$
calib_0.80_0.98 10.9 0.2372 2.60 $31.9\%$ $4.7\%$ $34.7\%$
damp_0.30 11.0 0.2368 2.60 $31.3\%$ $4.9\%$ $34.7\%$
Interpretation in the NFE–error plane. These comparisons summarize a consistent pattern: the adaptive curvature-aware sampler lies on (or very near) the Pareto frontier in the NFE–error plane, while fixed-step baselines are strictly dominated in the regimes tested. This is the practical meaning of the curvature theory: the controller allocates steps only where the learned transport bends, while retaining large steps when the flow is straight.
7.5 Distributional fidelity and diversity: ensemble $W_{1}$ under macro–micro evaluation
A common concern for deterministic probability-flow ODE sampling is potential under-dispersion or reduced diversity. Our evaluation is designed to test this directly in the conditional setting where diversity is well-defined.
Empirical $W_{1}$ between conditional ensembles.
For each macro, let $\{\widehat{u}^{(j)}\}_{j=1}^{N_{\mathrm{micro}}}$ be model samples and $\{u^{(j)}\}_{j=1}^{N_{\mathrm{micro}}}$ be reference samples. We compute a one-point (spatially averaged) empirical $W_{1}$ between the corresponding micro-ensembles, and then average across macros. We also report per-variable $W_{1}$ (density, momenta, energy) and an aggregated score.
Key empirical outcome.
Across benchmarks, our rectified-flow model matches or improves upon the diffusion baseline in ensemble $W_{1}$ while using substantially fewer function evaluations. In particular, the macro–micro protocol provides a direct quantitative check against mode collapse: matching $W_{1}$ at fixed macro conditions indicates that the conditional output distribution is captured rather than collapsed. (Complete per-variable $W_{1}$ tables and additional datasets are reported in the appendix.)
7.6 Practical takeaway: speed requires both straight transports and curvature-aware integration
The integrator study separates two effects: (i) the rectified-flow training objective yields a transport that can be traversed with far fewer evaluations than score-based diffusion, and (ii) within this ODE setting, curvature-aware control is a robust way to realize the efficiency in stiff multiscale regimes without per-problem solver tuning. Empirically, this is exactly what Tables 24 show: the controller achieves lower error at lower NFE than the strongest tuned fixed-step baseline, and its performance is stable across wide hyperparameter ranges.
Appendix pointers. We provide: (a) the same integrator/robustness tables on the second benchmark, (b) NFE–error scatter plots (Pareto frontiers), and (c) detailed per-variable $W_{1}$ breakdowns and additional OOD regimes.
8 Conclusion and Future Work
We have demonstrated that rectified flows provide an efficient surrogate for sampling multi-scale PDE solutions: on a suite of challenging incompressible and compressible problems, our method matches the statistical fidelity of conditional diffusion models while requiring an order of magnitude fewer solver steps—yielding up to 22× faster inference in practice. Across matched and shifted initial‐condition regimes, rectified flows consistently reproduce energy spectra, vorticity distributions, and one‐point Wasserstein distances within sampling error, confirming their robustness and multi‐scale accuracy.
Key Takeaways.
On account of its straightening bias, ReFlow requires significantly fewer ODE steps at inference-time (8-10) than its conditional-diffusion-based counterparts (more than 128) to achieve the same level of performance on our statistical metrics – consult SM M for more details. Furthermore, ReFlow is much more training efficient, requiring up to 120,000 iterations in reaching the same level of performance, across all experiments, as our other conditional diffusion baselines (which typically take 500-600,000 iterations), see SM M.2. Our results further demonstrate that straightening effects remain evident even when projecting down to a 2D PCA subspace (see SM M.3).
Appendix A Theory appendix: law-level rectified flows for multiscale CFD
Purpose and scope.
This appendix provides the mathematical backbone for the claims made in the main text: (i) rectified flows can match the same conditional solution laws as diffusion-class models, (ii) their sampling step count is governed by a curvature/straightness quantity rather than by the raw dimensionality of the state, and (iii) at fixed spatial resolution, there is an irreducible law-level error dictated by the small-scale tail of the true PDE law, which is naturally quantified through structure functions. We keep the presentation self-contained, and we explicitly connect each theoretical ingredient to the empirical evaluations in Table 1 and to the spectral diagnostics in Figure 13 (and supplementary figures).
How to read this appendix.
Section B fixes the measure-theoretic setting and links it to statistical solutions in the sense of Lanthaler et al. (2021a) (also used in the Introduction). Section C shows that the barycentric rectified velocity exactly transports $\mu_{0}$ to $\mu_{1}$ (correctness at the level of laws). Section D gives a terminal error decomposition into a fit term and a straightness term and explains how the training objective controls these terms. Section E derives why curvature controls the required number of ODE steps and motivates the EMA-based curvature proxy and the blending correction used at inference. Section F links the energy spectrum / structure functions to an unavoidable coverage term at fixed cutoff $K_{c}$. Finally, Section H aggregates the pieces into a master law-level bound that matches the narrative in the abstract and in the theory paragraphs of the main paper.
Appendix B Setup: laws of PDE solutions and what the experiments measure
State space.
We work on the periodic domain $\mathcal{D}=\mathbb{T}^{d}$ and consider (for notational simplicity) scalar or vector fields $u:\mathcal{D}\to\mathbb{R}^{m}$ represented on a finite grid. The natural infinite-dimensional space for the fluid benchmarks is
$\mathcal{H}:=L^{2}_{\sigma}(\mathcal{D};\mathbb{R}^{d})\quad\text{(divergence-free in the incompressible case),}$
or $\mathcal{H}:=L^{2}(\mathcal{D};\mathbb{R}^{m})$ for compressible state vectors (e.g. $(\rho,m_{x},m_{y},E)$). All constructions below are rigorous on the finite-dimensional Galerkin/grid surrogate $\mathbb{R}^{d_{N}}\subset\mathcal{H}$ used in training and evaluation. We state results in $\mathcal{H}$ language to highlight the measure-theoretic meaning.
Statistical solutions and correlation viewpoint.
In strongly nonlinear regimes, it is natural to treat the PDE solution as a random field. Given a law $\mu_{0}\in\mathcal{P}(\mathcal{H})$ for the initial condition, the law at time $t$ is
$\mu_{t}=(\mathcal{S}^{t})_{\#}\mu_{0},$
where $\mathcal{S}^{t}$ is the PDE solution operator (possibly multivalued in the inviscid setting; see Lanthaler et al. (2021a) for the correlation hierarchy framework). This is precisely the “statistical solution” language used in the Introduction.
What we evaluate empirically.
Our quantitative evaluation (Section 6.2) uses: (i) errors of the mean and standard deviation fields, and (ii) a pointwise Wasserstein-$1$ discrepancy averaged over space. These are marginal discrepancies: they test whether low-order statistics and one-point laws match between the model and the PDE ensemble.
From a law-level perspective, a natural global metric is $W_{2}$ on $\mathcal{P}(\mathcal{H})$. Although our reported metric is $W_{1}$ on one-point marginals (for computational reasons), the theory below uses $W_{2}$ because (a) it interacts cleanly with $L^{2}$-based structure functions and Fourier tails, and (b) it yields transparent decompositions. We emphasize that the core mechanisms (coverage/fit/straightness/discretization) are metric-agnostic: they explain why a method that reduces curvature reaches good statistics with fewer steps, which is exactly what is observed in Table 1 and Figure 13.
Appendix C Rectified-flow correctness: barycentric velocity transports laws
C.1 Couplings and linear interpolation in rectification time
Let $\mu_{0},\mu_{1}\in\mathcal{P}(\mathcal{H})$ be two laws (in our application: conditional laws of PDE outputs given inputs; see Section 4.1). Fix a coupling $\gamma\in\mathcal{P}(\mathcal{H}\times\mathcal{H})$ of $(\mu_{0},\mu_{1})$, i.e. $(\pi_{0})_{\#}\gamma=\mu_{0}$ and $(\pi_{1})_{\#}\gamma=\mu_{1}$.
For $\tau\in[0,1]$ define the interpolation map
$T_{\tau}(u_{0},u_{1}):=(1-\tau)u_{0}+\tau u_{1},\qquad\rho_{\tau}:=(T_{\tau})_{\#}\gamma.$
Thus $\rho_{\tau}$ is the law of the random variable $U_{\tau}=(1-\tau)U_{0}+\tau U_{1}$ when $(U_{0},U_{1})\sim\gamma$.
C.2 Barycentric rectified velocity and the continuity equation
Definition C.1 (Barycentric rectified velocity).
For each $\tau\in[0,1]$ and $\rho_{\tau}$-a.e. $u$, define
$v_{\star}(u,\tau):=\mathbb{E}\!\bigl[\,U_{1}-U_{0}\,\big|\,U_{\tau}=u\,\bigr].$
Equivalently, $v_{\star}$ is obtained by disintegrating $\gamma$ along the fibers $T_{\tau}^{-1}(u)$, as in equation 2 in the main text.
Why this is the “right” velocity.
If we think of $\rho_{\tau}$ as mass moving from $\rho_{0}=\mu_{0}$ to $\rho_{1}=\mu_{1}$ in rectification time, then the pair $(\rho_{\tau},v_{\star})$ solves the continuity equation
$\partial_{\tau}\rho_{\tau}+\nabla\!\cdot(\rho_{\tau}v_{\star}(\cdot,\tau))=0\quad\text{(on the finite-dimensional surrogate).}$
This is the deterministic analogue of the probability-flow (ODE) viewpoint of diffusion models: the flow map of $v_{\star}$ pushes $\rho_{0}$ to $\rho_{1}$.
Proposition C.2 (Exact transport under the barycentric velocity).
Assume we work on a finite-dimensional surrogate of $\mathcal{H}$ so that $u\in\mathbb{R}^{d_{N}}$ and $\rho_{\tau}$ admits a density (or, more generally, is a narrowly continuous curve of measures). Let $X_{\tau}$ be the (classical) flow map solving
$\frac{d}{d\tau}X_{\tau}(u)=v_{\star}(X_{\tau}(u),\tau),\qquad X_{0}(u)=u.$
Then $(X_{\tau})_{\#}\mu_{0}=\rho_{\tau}$ for all $\tau\in[0,1]$, and in particular $(X_{1})_{\#}\mu_{0}=\mu_{1}$.
Proof.
Let $(U_{0},U_{1})\sim\gamma$ and define $U_{\tau}=(1-\tau)U_{0}+\tau U_{1}$ so that $\operatorname{Law}(U_{\tau})=\rho_{\tau}$. Fix $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d_{N}})$. By differentiating under the expectation,
$\frac{d}{d\tau}\mathbb{E}[\varphi(U_{\tau})]=\mathbb{E}\bigl[\nabla\varphi(U_{\tau})\cdot(U_{1}-U_{0})\bigr].$
Condition on $U_{\tau}$ and use Definition C.1:
$\mathbb{E}\bigl[\nabla\varphi(U_{\tau})\cdot(U_{1}-U_{0})\bigr]=\mathbb{E}\bigl[\nabla\varphi(U_{\tau})\cdot\mathbb{E}[U_{1}-U_{0}\mid U_{\tau}]\bigr]=\mathbb{E}\bigl[\nabla\varphi(U_{\tau})\cdot v_{\star}(U_{\tau},\tau)\bigr].$
Thus $\rho_{\tau}$ satisfies the weak continuity equation
$\frac{d}{d\tau}\int\varphi(u)\,\rho_{\tau}(du)=\int\nabla\varphi(u)\cdot v_{\star}(u,\tau)\,\rho_{\tau}(du).$
On the other hand, the pushforward curve $\tilde{\rho}_{\tau}:=(X_{\tau})_{\#}\rho_{0}$ satisfies the same weak equation by standard characteristics. Uniqueness of solutions to the continuity equation on $\mathbb{R}^{d_{N}}$ (under the regularity ensuring existence of $X_{\tau}$) yields $\rho_{\tau}=\tilde{\rho}_{\tau}$, hence $(X_{1})_{\#}\mu_{0}=\mu_{1}$. ∎
Empirical tie.
Correctness in Proposition C.2 is a law-level statement: it says that if a learned velocity $v_{\theta}$ matches $v_{\star}$, then integrating the ODE recovers the target law. This is what is tested empirically through mean/spread/Wasserstein statistics (Table 1) and through the evolution of multiscale spectra along $\tau$ (Figure 13).
Appendix D Continuous-time error decomposition: fit vs. straightness
D.1 The training objective targets the barycentric velocity
In the (unconditional) rectified-flow construction, minimizing
$\int_{0}^{1}\mathbb{E}\bigl[\|(U_{1}-U_{0})-v(U_{\tau},\tau)\|^{2}\bigr]\;d\tau$
over measurable $v$ yields the conditional expectation $v_{\star}(u,\tau)=\mathbb{E}[U_{1}-U_{0}\mid U_{\tau}=u]$. In the conditional PDE setting we use the same principle but with $(U_{0},U_{1})$ sampled from the conditional pairing induced by the dataset and conditioning variables (initial/boundary data). This explains why the regression loss in Algorithm 1 is the natural analogue of score matching: it is an $L^{2}$ projection onto the optimal transport direction at each $\tau$.
D.2 Terminal error bound with explicit decomposition
Let $u_{\tau}$ solve the ideal rectified ODE $\dot{u}_{\tau}=v_{\star}(u_{\tau},\tau)$ and let $\hat{u}_{\tau}$ solve the learned ODE $\dot{\hat{u}}_{\tau}=v_{\theta}(\hat{u}_{\tau},\tau)$ with the same initial data $u_{0}=\hat{u}_{0}$.
Define the time-average velocity
$\bar{v}_{\theta}(u):=\int_{0}^{1}v_{\theta}(u,\tau)\,d\tau.$
The average isolates the “straight-line” component of the learned field, while deviations from it quantify curvature in rectification time.
Theorem D.1 (Terminal error decomposition (fit + straightness)).
Assume $v_{\theta}(\cdot,\tau)$ is $L$-Lipschitz in $u$ uniformly in $\tau$ and that $u_{\tau}$ remains in a region where the quantities below are finite. Define
$\varepsilon_{\mathrm{fit}}^{2}:=\int_{0}^{1}\mathbb{E}\bigl\|\bar{v}_{\theta}(u_{\tau})-v_{\star}(u_{\tau},\tau)\bigr\|^{2}\,d\tau,\qquad\varepsilon_{\mathrm{curv}}^{2}:=\int_{0}^{1}\mathbb{E}\bigl\|v_{\theta}(u_{\tau},\tau)-\bar{v}_{\theta}(u_{\tau})\bigr\|^{2}\,d\tau.$
Then
$\mathbb{E}\|\hat{u}_{1}-u_{1}\|\;\leq\;e^{L}\,\bigl(\varepsilon_{\mathrm{fit}}+\varepsilon_{\mathrm{curv}}\bigr).$ (28)
Moreover, the same argument in $L^{2}$ yields
$\mathbb{E}\|\hat{u}_{1}-u_{1}\|^{2}\;\leq\;C(L)\,\bigl(\varepsilon_{\mathrm{fit}}^{2}+\varepsilon_{\mathrm{curv}}^{2}\bigr),\qquad C(L)\sim e^{2L}.$ (29)
Proof.
Set $e_{\tau}:=\hat{u}_{\tau}-u_{\tau}$. Then
$\dot{e}_{\tau}=v_{\theta}(\hat{u}_{\tau},\tau)-v_{\star}(u_{\tau},\tau)=\bigl[v_{\theta}(\hat{u}_{\tau},\tau)-v_{\theta}(u_{\tau},\tau)\bigr]+\bigl[v_{\theta}(u_{\tau},\tau)-v_{\star}(u_{\tau},\tau)\bigr].$
Take norms and use Lipschitz continuity:
$\frac{d}{d\tau}\|e_{\tau}\|\leq L\|e_{\tau}\|+\|v_{\theta}(u_{\tau},\tau)-v_{\star}(u_{\tau},\tau)\|.$
Integrate from 0 to 11 and apply Grönwall:
$\|e_{1}\|\leq e^{L}\int_{0}^{1}\|v_{\theta}(u_{\tau},\tau)-v_{\star}(u_{\tau},\tau)\|\,d\tau.$
Now add and subtract $\bar{v}_{\theta}(u_{\tau})$ and use triangle inequality:
$\|v_{\theta}(u_{\tau},\tau)-v_{\star}(u_{\tau},\tau)\|\leq\|v_{\theta}(u_{\tau},\tau)-\bar{v}_{\theta}(u_{\tau})\|+\|\bar{v}_{\theta}(u_{\tau})-v_{\star}(u_{\tau},\tau)\|.$
Take expectations and apply Cauchy–Schwarz in $\tau$ to obtain (28). For (29), repeat with $\|e_{\tau}\|^{2}$:
$\frac{d}{d\tau}\|e_{\tau}\|^{2}=2e_{\tau}\cdot\dot{e}_{\tau}\leq 2L\|e_{\tau}\|^{2}+2\|e_{\tau}\|\,\|v_{\theta}(u_{\tau},\tau)-v_{\star}(u_{\tau},\tau)\|,$
and apply Young’s inequality plus Grönwall. The resulting constant is $C(L)\sim e^{2L}$. ∎
Empirical tie.
Theorem D.1 explains the two central practical observations: (i) if the learned velocity is nearly constant in $\tau$ (small straightness error), then the continuous-time mismatch is small, and (ii) straighter trajectories are easier to integrate, which translates into fewer solver steps at inference (Section E). This is exactly the mechanism behind “8 steps vs. 128 steps” in the main text and the sharp spectral tracking in Figure 13: early steps already reconstruct high wavenumbers because the trajectory is not spending many steps “turning” in rectification time.
Appendix E Discretization: why curvature controls step count, and why EMA helps
E.1 Local truncation error and the curvature driver
We start from a clean Taylor expansion for the ODE $\dot{u}=v(u,\tau)$.
Lemma E.1 (Local truncation error for explicit Euler).
Assume $v\in C^{1}$ in $(u,\tau)$ on the region traversed and that $\|\partial_{\tau}v(u,\tau)\|\leq M_{\tau}$, $\|J_{u}v(u,\tau)\|\leq L$, and $\|v(u,\tau)\|\leq M$. Then one explicit-Euler step from $(u,\tau)$ with step size $\Delta\tau$ has local truncation error
$\mathrm{LTE}(\tau;\Delta\tau)$ $:=\bigl\|u(\tau+\Delta\tau)-\bigl(u(\tau)+\Delta\tau\,v(u(\tau),\tau)\bigr)\bigr\|$
$=\frac{(\Delta\tau)^{2}}{2}\,\bigl\|\partial_{\tau}v(u,\tau)+J_{u}v(u,\tau)\,v(u,\tau)\bigr\|\;+\;\mathcal{O}\!\bigl((\Delta\tau)^{3}\bigr)$ (30)
$\leq\frac{(\Delta\tau)^{2}}{2}\,(M_{\tau}+LM)\;+\;\mathcal{O}\!\bigl((\Delta\tau)^{3}\bigr).$
Proof.
Differentiate $\dot{u}=v(u,\tau)$ along the trajectory:
$\ddot{u}(\tau)=\partial_{\tau}v(u(\tau),\tau)+J_{u}v(u(\tau),\tau)\,\dot{u}(\tau)=\partial_{\tau}v(u(\tau),\tau)+J_{u}v(u(\tau),\tau)\,v(u(\tau),\tau).$
Taylor expand:
$u(\tau+\Delta\tau)=u(\tau)+\Delta\tau\,v(u(\tau),\tau)+\frac{(\Delta\tau)^{2}}{2}\ddot{u}(\tau)+\mathcal{O}((\Delta\tau)^{3}).$
Subtract the Euler update and take norms to obtain (30). The bound follows by triangle inequality and the assumed sup bounds. ∎
Interpretation.
Lemma E.1 identifies two drivers of discretization error:
$\|\partial_{\tau}v\|\quad\text{(time curvature)}\qquad\text{and}\qquad\|J_{u}v\,v\|\quad\text{(spatial nonlinearity).}$
Rectified flows are designed so that trajectories are nearly straight in $\tau$, i.e. $\partial_{\tau}v$ is small along typical paths. This is the fundamental reason they can use large steps without losing fidelity.
E.2 Global error: straightness reduces the required number of steps
Corollary E.2 (Global Euler error controlled by time-curvature).
Assume the hypotheses of Lemma E.1 and that $v(\cdot,\tau)$ is $L$-Lipschitz in $u$ uniformly in $\tau$ on the region traversed. Let $u^{\Delta}$ be explicit Euler with uniform step $\Delta\tau=1/N$. Then there exists $C=C(L)$ such that
$\sup_{\tau\in[0,1]}\|u^{\Delta}_{\tau}-u_{\tau}\|\;\leq\;C\,\Delta\tau\,\int_{0}^{1}\Big(\|\partial_{\tau}v(u_{\tau},\tau)\|+\|J_{u}v(u_{\tau},\tau)\,v(u_{\tau},\tau)\|\Big)\,d\tau,$ (31)
up to higher-order terms.
Proof.
Write the one-step error recursion $e_{n+1}\leq(1+L\Delta\tau)e_{n}+\mathrm{LTE}_{n}$ and iterate it over $n=0,\dots,N-1$. Use $\prod_{k=0}^{N-1}(1+L\Delta\tau)\leq e^{L}$ and sum the LTE bounds from Lemma E.1. The integral form (31) is obtained by viewing the sum as a Riemann approximation of $\int_{0}^{1}(\cdots)d\tau$. ∎
Empirical tie.
Equation (31) explains the step-count gap we observe: if $\int_{0}^{1}\|\partial_{\tau}v\|$ is small (straight trajectories), then the $N$ required for a fixed tolerance is small. Diffusion-class samplers typically integrate a much more curved reverse-time dynamics (in addition to stochasticity), which translates into much larger effective curvature and hence many more steps. This is also consistent with the qualitative “trajectory inpainting” comparison in Figure 4.
E.3 EMA curvature proxy and blending as a Tikhonov correction
Discrete-time EMA as a low-pass filter.
During sampling we evaluate $v_{t}:=v_{\theta}(u_{t},\tau_{t})$ at a discrete set of times $\tau_{t}$ and maintain an exponential moving average (EMA)
$v^{\mathrm{ema}}_{t}=\lambda\,v^{\mathrm{ema}}_{t-1}+(1-\lambda)\,v_{t},\qquad\lambda\in(0,1).$
If $v_{t}$ varies slowly with $\tau_{t}$, then $v^{\mathrm{ema}}_{t}$ tracks it closely; if $v_{t}$ changes quickly, the EMA lags and the deviation
$s_{t}:=\|v_{t}-v^{\mathrm{ema}}_{t}\|$
spikes. This motivates $s_{t}$ as a computable curvature/straightness proxy.
A simple bound shows $s_{t}$ controls a discrete time-variation scale.
Lemma E.3 (EMA deviation controls incremental variation).
Let $\{v_{t}\}_{t\geq 0}$ be any sequence in a normed space and define $v_{t}^{\mathrm{ema}}=\lambda v_{t-1}^{\mathrm{ema}}+(1-\lambda)v_{t}$ with $v_{0}^{\mathrm{ema}}=v_{0}$. Then
$\|v_{t}-v_{t}^{\mathrm{ema}}\|\leq(1-\lambda)\sum_{j=0}^{t-1}\lambda^{j}\,\|v_{t-j}-v_{t-j-1}\|.$
In particular, if $\|v_{k}-v_{k-1}\|\leq\delta$ for all $k$, then $\|v_{t}-v_{t}^{\mathrm{ema}}\|\leq\delta$ for all $t$.
Proof.
Unroll the recursion: $v_{t}^{\mathrm{ema}}=(1-\lambda)\sum_{j=0}^{t}\lambda^{j}v_{t-j}$. Subtract from $v_{t}$ and telescope:
$v_{t}-v_{t}^{\mathrm{ema}}=(1-\lambda)\sum_{j=0}^{t-1}\lambda^{j}\,(v_{t}-v_{t-j-1}),$
then use triangle inequality and rewrite $v_{t}-v_{t-j-1}$ as a sum of increments. ∎
Blending as a quadratic minimizer.
We use the blended update
$\tilde{v}_{t}=(1-\alpha_{t})v_{t}+\alpha_{t}v_{t}^{\mathrm{ema}},\qquad\alpha_{t}\in[0,1].$
This is the unique minimizer of a Tikhonov-regularized quadratic:
$\tilde{v}_{t}=\operatorname*{arg\,min}_{w}\Big\{\|w-v_{t}\|^{2}+\beta_{t}\|w-v_{t}^{\mathrm{ema}}\|^{2}\Big\},\qquad\alpha_{t}=\frac{\beta_{t}}{1+\beta_{t}}.$
Thus blending is not an ad-hoc smoothing step: it is the closed-form solution of a principled stabilization objective trading fidelity to the instantaneous network prediction against adherence to a locally averaged trend.
Step-size rule from LTE control.
Lemma E.1 suggests choosing $\Delta\tau_{t}$ so that
$(\Delta\tau_{t})^{2}\Big(\|\partial_{\tau}v\|+\|J_{u}v\,v\|\Big)\approx\text{const}.$
Since $\|\partial_{\tau}v\|$ is not directly available, we replace it by a multiple of the proxy $s_{t}$ (Lemma E.3), and we upper-bound $\|J_{u}v\,v\|\lesssim L\|v_{t}\|$ using Lipschitz control. This yields the square-root-type rule in the main text:
$\Delta\tau_{t}\propto\frac{1}{\sqrt{\kappa_{1}s_{t}+\kappa_{2}}},\qquad\kappa_{2}\approx L\|v_{t}\|.$
This is exactly the mechanism behind the “curvature-aware integration” contribution: when the trajectory bends (large $s_{t}$), we both (i) regularize the velocity via blending and (ii) reduce the step size to control the discretization term in (31).
Appendix F Multiscale coverage: structure functions control unavoidable law tails
F.1 Why spectra/structure functions appear in a law-level bound
Finite resolution is a hard constraint.
All models operate on a fixed grid (e.g. $128^{2}$ for training/evaluation). This enforces an effective Fourier cutoff $K_{\max}\simeq N$ (Nyquist), and architectural choices may impose an even lower effective cutoff $K_{c}$. Therefore, even a perfect generative model cannot represent energy beyond that cutoff. The best it can do is match the projected law $(P_{\leq K_{c}})_{\#}\mu$.
Structure functions quantify tail energy.
In turbulence, small-scale content is captured by structure functions and energy spectra. This is exactly why our empirical analysis emphasizes spectral evolution along $\tau$ (Figure 13). The theory below formalizes the link: a structure-function modulus bound implies a Fourier-tail bound, which becomes a coverage term in $W_{2}$.
F.2 Bandlimited fields: capacity bounds and why missing an annulus matters
We record two standard Fourier facts to make the “capacity” language explicit.
Lemma F.1 (Bernstein upper bound for bandlimited fields).
Let $u:\mathbb{T}^{d}\to\mathbb{R}^{m}$ satisfy $u=P_{\leq K}u$ for some integer $K\geq 1$. Then there exists $C_{d}>0$ such that
$\|\nabla_{x}u\|_{L^{\infty}(\mathbb{T}^{d})}\leq C_{d}\,K^{1+\frac{d}{2}}\,\|u\|_{L^{2}(\mathbb{T}^{d})}.$
Proof.
Expand $u(x)=\sum_{|k|\leq K}\hat{u}(k)e^{ik\cdot x}$, write $\nabla u$ as a Fourier series, and apply Cauchy–Schwarz using that $\{|k|\leq K\}\lesssim K^{d}$ and Parseval. ∎
Lemma F.2 (Annulus lower bound).
Let $f=P_{[K,2K]}u$ and $K\geq 1$. Then there exists $c_{d}>0$ such that
$\|\nabla_{x}f\|_{L^{\infty}}\geq c_{d}\,K^{1-\frac{d}{2}}\,\|f\|_{L^{2}}.$
Proof.
On the annulus $K\leq|k|\leq 2K$, the Fourier symbol of $\nabla$ has magnitude $\simeq K$. A reverse Bernstein estimate on an annulus yields $\|f\|_{L^{\infty}}\gtrsim K^{-d/2}\|f\|_{L^{2}}$, and $\|\nabla f\|_{L^{\infty}}\gtrsim K\|f\|_{L^{\infty}}$, giving the claim. ∎
Empirical tie.
Lemma F.2 formalizes why “spectral collapse” is fatal for turbulent statistics: if a surrogate misses energy in a whole annulus, it underestimates strain and small-scale variability. This is exactly what appears visually when deterministic baselines blur fine scales and also what is quantified by the spread errors in Table 1.
F.3 Structure functions imply Fourier-tail control (law-level)
For $u:\mathbb{T}^{d}\to\mathbb{R}^{m}$ define the second-order structure function
$S_{r}^{2}(u):=\frac{1}{|B_{r}|}\int_{|h|\leq r}\|u(\cdot+h)-u(\cdot)\|_{L^{2}(\mathbb{T}^{d})}^{2}\,dh,\qquad S_{r}^{2}(\mu):=\int S_{r}^{2}(u)\,d\mu(u).$
Following the statistical-solution framework of Lanthaler et al. (2021a), assume there exists a modulus $\omega(r)\downarrow 0$ as $r\downarrow 0$ such that
$S_{r}^{2}(\mu_{t})\leq\omega(r)\qquad\text{for all }r\in(0,1],\ t\in[0,T].$ (32)
In practice, $\omega$ can be calibrated from empirical spectra (Figure 13 and supplementary plots).
Lemma F.3 (Structure function controls Fourier tail).
Let $\mu$ be a law on $L^{2}(\mathbb{T}^{d})$ satisfying $S_{r}^{2}(\mu)\leq\omega(r)$ for all $r\in(0,1]$. Then there exist $c_{d},C_{d}>0$ such that for all integers $K\geq 1$,
$\int\|P_{>K}u\|_{L^{2}}^{2}\,d\mu(u)\leq C_{d}\,\omega\!\bigl(c_{d}/K\bigr).$
Proof.
For fixed $u$, expand $\|u(\cdot+h)-u(\cdot)\|_{L^{2}}^{2}=\sum_{k}|\hat{u}(k)|^{2}|e^{ik\cdot h}-1|^{2}$. Average over $|h|\leq r$ to obtain $S_{r}^{2}(u)=\sum_{k}|\hat{u}(k)|^{2}\Psi_{r}(k)$ where $\Psi_{r}(k)=\frac{1}{|B_{r}|}\int_{|h|\leq r}|e^{ik\cdot h}-1|^{2}\,dh$. A standard oscillation estimate yields $\Psi_{r}(k)\geq c\min\{(|k|r)^{2},1\}$. Hence $\Psi_{r}(k)\geq c_{0}$ for all $|k|\geq c^{\prime}/r$, and therefore $S_{r}^{2}(u)\geq c_{0}\sum_{|k|\geq c^{\prime}/r}|\hat{u}(k)|^{2}=c_{0}\|P_{>c^{\prime}/r}u\|_{L^{2}}^{2}$. Integrate in $u$ and set $r=c_{d}/K$. ∎
Lemma F.4 (Bandlimited projection in $W_{2}$).
For any law $\mu$ on $L^{2}(\mathbb{T}^{d})$ and cutoff $K_{c}\geq 1$,
$W_{2}\bigl(\mu,(P_{\leq K_{c}})_{\#}\mu\bigr)\leq\Bigl(\int\|P_{>K_{c}}u\|_{L^{2}}^{2}\,d\mu(u)\Bigr)^{1/2}.$
Proof.
Couple $u\sim\mu$ with $\tilde{u}=P_{\leq K_{c}}u$. Then $\tilde{u}\sim(P_{\leq K_{c}})_{\#}\mu$ and $W_{2}^{2}\leq\mathbb{E}\|u-\tilde{u}\|_{L^{2}}^{2}=\mathbb{E}\|P_{>K_{c}}u\|_{L^{2}}^{2}$. ∎
Corollary F.5 (Coverage term from structure-function modulus).
Under (32), for any cutoff $K_{c}\geq 1$,
$W_{2}\bigl(\mu_{t},(P_{\leq K_{c}})_{\#}\mu_{t}\bigr)\lesssim\omega(c_{d}/K_{c})^{1/2}.$
Empirical tie.
Corollary F.5 is the clean mathematical version of what the spectra show: at fixed resolution, there is an unavoidable tail beyond the cutoff. Both ReFlow and GenCFD operate at the same grid resolution (hence comparable $K_{c}$), so neither can beat the coverage floor. The difference is that ReFlow reaches the best attainable projected law with far fewer steps (Figure 13, Figure 4).
Appendix G One-step law error for PDE pushforwards and training fit
Let $S_{\Delta t}:\mathcal{H}\to\mathcal{H}$ denote the PDE solution operator over a physical step $\Delta t$ and $\mu_{t+\Delta t}=(S_{\Delta t})_{\#}\mu_{t}$. Let $\mathcal{T}^{\theta}_{\Delta t}$ be the learned one-step surrogate and $\hat{\mu}_{t+\Delta t}=(\mathcal{T}^{\theta}_{\Delta t})_{\#}\mu_{t}$.
Why we insert a bandlimit.
Since the model outputs live at finite resolution, it is natural to compare to the projected PDE output $P_{\leq K_{c}}S_{\Delta t}u$. This isolates (i) an irreducible coverage term and (ii) a trainable fit term.
Define the mean-square training error at time $t$:
$\varepsilon_{\mathrm{train}}^{2}(t):=\mathbb{E}_{u\sim\mu_{t}}\Bigl[\,\bigl\|P_{\leq K_{c}}(S_{\Delta t}u)-\widehat{u}\bigr\|_{L^{2}}^{2}\Bigr],$ (33)
where $\widehat{u}$ is the model output given input $u$ (i.e. after integrating the learned RF ODE).
Proposition G.1 (One-step law error = coverage + fit).
Assume $\mu_{t+\Delta t}$ satisfies the structure-function modulus (32) (hence Corollary F.5). Then
$W_{2}\bigl(\mu_{t+\Delta t},\hat{\mu}_{t+\Delta t}\bigr)\;\leq\;\underbrace{W_{2}\bigl(\mu_{t+\Delta t},(P_{\leq K_{c}})_{\#}\mu_{t+\Delta t}\bigr)}_{\text{coverage}}\;+\;\underbrace{\bigl(\varepsilon_{\mathrm{train}}^{2}(t)\bigr)^{1/2}}_{\text{fit}},$
and therefore
$W_{2}\bigl(\mu_{t+\Delta t},\hat{\mu}_{t+\Delta t}\bigr)\;\lesssim\;\omega(c_{d}/K_{c})^{1/2}+\varepsilon_{\mathrm{train}}(t).$
Proof.
Use the triangle inequality in $W_{2}$:
$W_{2}(\mu_{t+\Delta t},\hat{\mu}_{t+\Delta t})\leq W_{2}\bigl(\mu_{t+\Delta t},(P_{\leq K_{c}})_{\#}\mu_{t+\Delta t}\bigr)+W_{2}\bigl((P_{\leq K_{c}})_{\#}\mu_{t+\Delta t},\hat{\mu}_{t+\Delta t}\bigr).$
For the second term, couple by sampling $u\sim\mu_{t}$ and using the same input to produce $P_{\leq K_{c}}S_{\Delta t}u$ and $\widehat{u}$. Then
$W_{2}^{2}\bigl((P_{\leq K_{c}})_{\#}\mu_{t+\Delta t},\hat{\mu}_{t+\Delta t}\bigr)\leq\mathbb{E}\|P_{\leq K_{c}}S_{\Delta t}u-\widehat{u}\|_{L^{2}}^{2}=\varepsilon_{\mathrm{train}}^{2}(t).$
The second display follows from Corollary F.5. ∎
Empirical tie.
Proposition G.1 explains why training curves saturate at a resolution-dependent floor and why spectra are the correct diagnostic for that floor. It also makes explicit that “better training loss” means smaller law error only up to the coverage limit. This aligns with the observation that increasing diffusion steps improves spectral reconstruction but cannot beat finite-resolution constraints, while ReFlow reaches that regime with far fewer steps.
Appendix H Master inequality: coverage + fit + straightness + discretization
We now combine the law-level coverage/fit decomposition with the continuous-time RF mismatch and the discretization control.
Theorem H.1 (Master law-level bound for learned rectified flows).
Fix an effective cutoff $K_{c}$ and assume the structure-function modulus (32) for the target PDE laws. Let $\hat{\mu}_{1}^{(N)}$ be the law obtained by integrating the learned rectified ODE $\dot{u}_{\tau}=v_{\theta}(u_{\tau},\tau)$ from $\tau=0$ to $\tau=1$ using $N$ explicit-Euler steps (uniform step $1/N$). Assume $v_{\theta}(\cdot,\tau)$ is $L$-Lipschitz in $u$ uniformly in $\tau$.
Then, for constants depending only on $L$,
$W_{2}(\mu_{1},\hat{\mu}_{1}^{(N)})\;\lesssim\;\underbrace{\omega(c_{d}/K_{c})^{1/2}}_{\text{coverage}}\;+\;\underbrace{\varepsilon_{\mathrm{fit}}+\varepsilon_{\mathrm{curv}}}_{\text{fit + straightness}}\;+\;\underbrace{\frac{1}{N}\int_{0}^{1}\|\partial_{\tau}v_{\theta}(u_{\tau},\tau)\|\,d\tau}_{\text{discretization}},$ (34)
where $\varepsilon_{\mathrm{fit}},\varepsilon_{\mathrm{curv}}$ are the continuous-time terms from Theorem D.1 evaluated along the ideal trajectory.
Proof.
(1) Coverage. By Corollary F.5, $W_{2}(\mu_{1},(P_{\leq K_{c}})_{\#}\mu_{1})\lesssim\omega(c_{d}/K_{c})^{1/2}$.
(2) Continuous-time learned flow error. Let $\hat{\mu}_{1}$ denote the law produced by the exact (continuous-time) flow of $v_{\theta}$. Then Theorem D.1 implies a bound of the form $W_{2}((P_{\leq K_{c}})_{\#}\mu_{1},\hat{\mu}_{1})\lesssim\varepsilon_{\mathrm{fit}}+\varepsilon_{\mathrm{curv}}$ (up to the standard coupling-by-same-input argument on the finite-dimensional surrogate).
(3) Discretization. Let $\hat{\mu}_{1}^{(N)}$ be the Euler-discretized law. Corollary E.2 bounds the trajectory error by a term proportional to $\frac{1}{N}\int_{0}^{1}\|\partial_{\tau}v_{\theta}\|\,d\tau$ (plus the spatial term, absorbed in constants under the Lipschitz assumption and bounded $\|v_{\theta}\|$ on the region traversed).
(4) Combine with triangle inequalities:
$W_{2}(\mu_{1},\hat{\mu}_{1}^{(N)})\leq W_{2}(\mu_{1},(P_{\leq K_{c}})_{\#}\mu_{1})+W_{2}((P_{\leq K_{c}})_{\#}\mu_{1},\hat{\mu}_{1})+W_{2}(\hat{\mu}_{1},\hat{\mu}_{1}^{(N)}),$
and insert the three bounds. ∎
What Theorem H.1 explains in the paper.
  • Why rectification accelerates sampling. The discretization term depends on the integral of time-curvature. Straightness reduces that integral, so fewer steps suffice (8–10 in our experiments).
  • Why curvature-aware integration helps under shift. Under distribution shift, the learned velocity may bend more in $\tau$, increasing the last term. The EMA proxy and blending/step control reduce the effective curvature and hence stabilize sampling, which is exactly how curvature is used in our method (control, not OOD detection).
  • Why spectra matter. The coverage term is dictated by small-scale physics (structure functions / spectrum). This is why Figure 13 is the right diagnostic: it visualizes how quickly the method reconstructs high-frequency energy up to the grid cutoff, and it clarifies what errors are unavoidable at fixed resolution.
  • Why deterministic baselines struggle. Deterministic predictors trained by pointwise losses effectively target conditional means; in multimodal chaotic regimes this collapses variability (as emphasized by Molinaro et al. (2024)). In the master inequality language, they do not produce a good fit to the law itself, which is seen in their larger spread and Wasserstein errors in Table 1.
Appendix I From law-level bounds to reported pointwise metrics
From $W_{2}$ to marginal discrepancies.
While the paper reports one-point $W_{1}$ (averaged over space) for computational tractability, the mechanisms above are consistent with those marginal tests: if two laws on fields are close in $W_{2}$ and have bounded second moments, then many bounded-Lipschitz observables of the fields (including smoothed one-point marginals) are close as well. This is the conceptual bridge between the law-level theory and the reported statistics.
Why we emphasize spectra during sampling.
The spectral evolution plots (Figure 13 and supplementary variants) act as a scale-resolved diagnostic: they visualize the progressive recovery of high frequencies along $\tau$. This connects directly to the structure-function/coverage viewpoint (Section F) and empirically demonstrates that ReFlow reaches the best attainable high-frequency reconstruction regime in far fewer steps than diffusion-based baselines.
Appendix J Architecture of the ReFlow Model
Figure 5: UNet Backbone Architecture Used in ReFlow. This schematic illustrates the core UNet-based architecture used within the ReFlow framework, structured across three resolution levels. For clarity, the number of blocks per level is set to one in this illustration. In the actual GenCFD and ReFlow configurations, each block is repeated four times per level. The bottleneck shows an asymmetry between the encoder and decoder sides: the block on the encoder side includes a Convolution Block, Positional Embedding, and Attention Block, while the corresponding block on the decoder side omits the Positional Embedding.
J.1 Architectural Components
Motivation. Multi-scale PDE data contain wide-ranging spatial scales, so a multi-resolution encoder–decoder is natural. We incorporate time embeddings to condition each layer on the interpolation fraction $t$.
Architecture Layout. We employ a UNet-based architecture augmented with attentional layers, typically termed as UViT by Saharia et al. (2022). In addition to the MLP-based time embedding $\Gamma(\tau)$, the input signal is lifted into a higher-dimensional embedding space and later projected back into physical space through convolutional layers. The data flows from high to low resolution across a down-sampling stack composed of convolution blocks into a bottleneck layer that encodes the fine-scale features of the inputs across its channel dimension and mixes them via a multi-head attention mechanism.
On the encoder side, exclusively at the bottleneck level, each attention block preceded by a simple linear positional embedding. This positional encoding helps preserve spatial locality before global mixing through attention. The decoder side does not apply any positional embedding.
The UNet is conditioned at every level through normalized Feature-wise Linear Modulation (FiLM) techniques embedded within the convolution blocks. The conditioning signal modulates the normalization layers via learned scale and shift parameters. This conditioning mechanism is consistently applied across all encoder and decoder stages.
The output is symmetrically reconstructed via an up-sampling pipeline. Downsampling is performed using standard convolutional layers with stride, while upsampling can be implemented either via transposed convolutions or by applying a standard convolution followed by a non-learnable pixel shuffle operation, which rearranges elements of the tensor spatially according to a fixed upsampling factor.
Importantly, the architecture described here resembles the one presented in Molinaro et al. (2024), ensuring consistency with prior state-of-the-art practices in high-fidelity generative modeling for scientific computing.
J.1.1 Convolution Block
Each UNet level employs dedicated convolution blocks that process features while enabling conditioning on the diffusion time $\tau$. These blocks consist of two convolutional layers interleaved with group normalization and Swish activations. Between the convolutions, a normalized FiLM layer adaptively scales features using the conditioning embeddings. A residual connection combines the processed features with a projected skip connection, ensuring stable gradient flow. Crucially, these convolution blocks form the fundamental building units across all encoder and decoder stages, not just the bottleneck, with consistent application of temporal conditioning through the FiLM mechanism.
J.1.2 Normalized FiLM Layer
Each convolution block uses a normalized FiLM layer and applies a convolution layer interleaved with normalization and an activation, then adds a residual skip:
$\mathbf{z}$ $=\;\mathrm{Conv}\bigl(\mathrm{Norm}(\mathbf{x})\bigr),\qquad\mathbf{z}\;=\;\sigma(\mathbf{z}+\alpha\,\gamma(t)),$ (35)
$\mathbf{x}_{\mathrm{out}}$ $=\;\mathbf{x}\;+\;\mathrm{Conv}\!\bigl(\mathbf{z}\bigr),$
where $\mathrm{Norm}$ is often RMSNorm, GroupNorm or LayerNorm, $\sigma$ is a nonlinearity (e.g. $\mathrm{SiLU}$), and $\gamma(t)$ represents the time embedding (an MLP on $\mathrm{Sinusoid}(t)$) that modulates scale/shift. This temporal conditioning is crucial for guiding the generative process through rectified diffusion trajectories.
J.1.3 Attention Block
The architecture strategically employs multi-head attention exclusively at the bottleneck layer to balance global interaction modeling with computational efficiency. Each attention block processes normalized features through a spatial-channel reformatting, where features are reshaped by a flattening operation. This preserves the channel structure while collapsing the spatial grid into a single flattened dimension, allowing attention to operate across spatial locations. At certain scales, a multi-head attention (MHA) mechanism captures distant flow interactions. We define
$\mathbf{Q}\;=\;W_{Q}\,\mathbf{h},\quad\mathbf{K}\;=\;W_{K}\,\mathbf{h},\quad\mathbf{V}\;=\;W_{V}\,\mathbf{h},\quad\mathcal{A}(\mathbf{h})\;=\;\mathrm{softmax}\!\Bigl(\tfrac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{d}}\Bigr)\,\mathbf{V},$ (36)
where $\mathbf{h}\in\mathbb{R}^{(H\cdot W)\times d}$ denotes the flattened input sequence, with each of the $H\cdot W$ spatial positions represented as a $d$-dimensional feature vector. In practice, attention is split into multiple heads for better representational power, then recombined. This global mixing is key for PDE flows dominated by far-field couplings (e.g. vortex merging). The MHA operator is defined as
$\mathrm{MHA}(\mathbf{h})\;=\;W^{O}\;\mathrm{Concat}\left(\mathcal{A}_{1}(\mathbf{h}),\dots,\mathcal{A}_{n}(\mathbf{h})\right),$ (37)
where $n\in\mathbb{N}$ denotes the number of attention heads and $W^{O}\in\mathbb{R}^{C\times C}$ is a learned output projection matrix. The attention block applies MHA to a normalized input, then adds a residual connection to preserve the original signal and stabilize training while enabling global context to influence local features:
$\mathbf{z}$ $=\;\mathrm{MHA}\bigl(\mathrm{Norm}(\mathbf{x})\bigr),$ (38)
$\mathbf{x}_{\mathrm{out}}$ $=\;\mathbf{x}\;+\;\mathbf{z}.$
J.2 Training Pipeline Details
The training pipeline follows the algorithm described in the main text faithfully.
Trainer Workflow.
A high-level Trainer class manages:
  1. 1.
    Batching PDE pairs $(u_{0},u_{1})$. These normalized pixel-wise with statistics extracted per each entire dataset.
  2. 2.
    Sampling $\tau\in[0,1]$ and random noise $\xi$ according to a given noise scheduler $\sigma(\cdot)$.
  3. 3.
    Partial interpolation + noise addition: $\tilde{u}_{\tau}:=\tau\,u_{1}+(1-\tau)\,\xi$ (or the flow-based variant). We have experimented with non-linear interpolation methods, but have not noticed meaningful oscillations in performance.
  4. 4.
    Minimizing the squared deviation between model output and target displacement/noise, optionally with a consistency/EMA update.
  5. 5.
    Throughout training, we monitor validation performance on a separate holdout dataset across all main metrics. We explicitly note that different validation splits have been tested out to ensure there is no data leakage that would bias performance metrics.
Table 5: Representative hyperparameters for 2D fluid-flow tasks. See main text for justification of these choices.
Parameter Value (Optimal empirical performance) Notes
Channels per level (128, 256, 256) Number of channels per level.
Downsample Ratios (2, 2, 2) Downsampling ratio per level.
Attention Heads 8 Each attention block.
Dropout Rate 0.1 In ConvBlock for generalization.
Noise Schedule uniform, log-uniform, cos-map Helps with multi-scale noise.
ODE Steps (Sampling) 4–8 Often $15\times$ fewer vs. diffusion.
Batch Size 16 Tied to GPU capacity.
EMA Decay 0.9999 Teacher model for consistency.
Learning Rate $3\times 10^{-4}$ Cosine decay in code.
In practice we noticed that the most robust denoising schedule was just the uniform one. For details on the cos-map schedule, we direct the reader to its description in Esser et al. (2024).
J.3 Summary and Key Takeaways
The RectifiedFlow class, combined with a time-aware UNet2D architecture, builds a deterministic transport model well-suited to multi-scale PDE tasks. By interpolating between noise and data along nearly-straight paths, it reduces sampling overhead compared to conventional diffusion PDE solvers while retaining the flexibility to model complex turbulent phenomena. Further experimental and theoretical results appear in SM Section M. Instead of the multi-step stochastic integration in diffusion-based models, RectifiedFlow fits a deterministic ODE from noise to data. This leverages straighter trajectories in frequency space, enabling fewer network evaluations at inference time.
ODE Integration for Sampling.
At inference:
$\frac{d\tilde{u}_{\tau}}{d\tau}\;=\;v_{\theta}\bigl(\tilde{u}_{\tau},\,u_{0},\,\tau\bigr),\quad\tau\in[0,1],\quad\tilde{u}_{\tau=0}\,=\,\text{random noise}.$
We solve this with odeint. Few steps (8108\!-\!10) typically suffice, contrasting with the tens or hundreds required by full diffusion samplers.
Appendix K Datasets
K.1 Multi-scale Flow Datasets
Following the dataset generation procedure of Herde et al. Herde et al. (2024), the Richtmyer–Meshkov ensemble is created by imposing a randomized sinusoidal perturbation on a two‐fluid interface and driving it with a planar shock via prescribed pressure jumps. For the CloudShock and ShearFlow benchmarks, we follow the GenCFD mesh‐generation and solver setup of Molinaro et al. Molinaro et al. (2024), initializing CloudShock with concentric density perturbations and ShearFlow with orthogonal shear jets on adaptive Cartesian grids, and employing the same high‐order finite‐volume scheme and dissipation settings to ensure consistent numerical fidelity across both datasets.
Overview. We experiment on three paired‐field benchmarks $(u_{0},u_{1})$:
  • Richtmyer–Meshkov (RM): 64 000 samples
  • Cloud-Shock (CS): 40 000 samples
  • Shear‐Layer (SL): 79 200 samples
All datasets track the $u_{0}\to S^{t}(u_{0})|_{t=1}$ solution mapping from inputs to the evolved outputs at $t=1$. The first two datasets monitor the evolution of density ($\rho$), momentum components $(m_{x},m_{y})$, and energy respectively ($E$). The shear-flow dataset monitors the velocity components of the flow $(u_{x},u_{y})$.
Train/Validation Split. For each dataset, we reserve 80% of the samples for training and 20% for validation.
Macro–Micro Ensemble Evaluation (CS & SL). To probe out‐of‐distribution generalization under small perturbations, we adopt a two‐stage ensemble protocol:
  1. 1.
    Macro‐sampling: select $M_{\mathrm{macro}}=10$ distinct base initial conditions.
  2. 2.
    Micro‐perturbation: for each base, generate $M_{\mathrm{micro}}=1{,}000$ perturbed copies within a small radius.
  3. 3.
    Metrics: compute all performance measures (mean‐field error $e_{\mu}$, standard‐deviation error $e_{\sigma}$, average $W_{1}$, spectral agreement) by first averaging over each micro‐ensemble, then averaging across the $M_{\mathrm{macro}}$ cases.
In‐Distribution Evaluation (RM). For the Richtmyer–Meshkov dataset, we additionally evaluate on a held‐out test set drawn from the same distribution (no macro–micro perturbations). This allows us to study how deterministic and probabilistic models scale with training set size under standard, in‐distribution conditions.
K.2 Preprocessing and Data Organization
Data Processing
Before training, we compute and store global, channel‐wise statistics over the entire training dataset. Concretely, for each physical variable $c$ we calculate
$\mu_{c}$ $=\frac{1}{N_{\rm train}\,H\,W\,[D]}\sum_{i=1}^{N_{\rm train}}\sum_{x,y,[z]}u^{(i)}_{c}\bigl(x,y,[z],t_{0}\bigr),$
$\sigma_{c}$ $=\sqrt{\frac{1}{N_{\rm train}\,H\,W\,[D]}\sum_{i=1}^{N_{\rm train}}\sum_{x,y,[z]}\Bigl(u^{(i)}_{c}\bigl(x,y,[z],t_{0}\bigr)-\mu_{c}\Bigr)^{2}}.$
where $N_{\rm train}$ is the number of training samples, $H\times W\,[\times D]$ the spatial grid size, and $u_{c}(x,y,[z],t_{0})$ the initial field for channel $c$. We average over all spatial locations to yield a single mean $\mu_{c}$ and standard deviation $\sigma_{c}$ per channel. The same statistics are also computed over the final‐time fields $u_{c}(t_{f})$, producing $\mu_{c}^{\prime}$ and $\sigma_{c}^{\prime}$ for the output channels.
During both training and evaluation, each input field $u_{c}$ is standardized via
$\tilde{u}_{c}=\frac{\,u_{c}\;-\;\mu_{c}\,}{\sigma_{c}+10^{-16}},\qquad\tilde{v}_{c}=\frac{\,v_{c}\;-\;\mu_{c}^{\prime}\,}{\sigma_{c}^{\prime}+10^{-16}},$
where $u_{c}$ and $v_{c}$ denote the initial and target fields, respectively. We then concatenate $\{\tilde{u}_{c}\}$ and $\{\tilde{v}_{c}\}$ along the channel axis before feeding them to the network. Because we always apply the same pre‐computed statistics at test time, no information from the evaluation set ever influences these normalization parameters.
Test‐time ensemble perturbation (evaluation only).
We adopt the same testing setup proposed in GenCFD Molinaro et al. (2024). To evaluate how well our model captures the distribution of solutions $u(t)$ arising from a fixed, chaotic initial field $\bar{u}$, we generate a small perturbed ensemble around $\bar{u}$ and evolve each member with a high‐fidelity PDE solver—but only at evaluation time. Specifically:
  1. 1.
    Micro‐ensemble generation. Around the test initial condition $\bar{u}$, draw $M_{\mathrm{micro}}$ perturbed copies $\{\bar{u}_{j}\}_{j=1}^{M_{\mathrm{micro}}}$ by sampling uniformly in a ball of radius $\varepsilon$ centered at $\bar{u}$.
  2. 2.
    Reference propagation. Integrate each perturbed field $\bar{u}_{j}$ forward to time $t$ using the reference solver, producing end‐states $\{u_{j}(t)\}$.
  3. 3.
    Empirical conditional law. Form the empirical measure
    $\hat{P}_{t}(\,\cdot\mid\bar{u})\;=\;\frac{1}{M_{\mathrm{micro}}}\sum_{j=1}^{M_{\mathrm{micro}}}\delta_{u_{j}(t)},$
    which approximates the true chaotic conditional law $P_{t}(\,\cdot\mid\bar{u})$.
When the input distribution itself is non‐degenerate, one would first draw $M_{\mathrm{macro}}$ base fields from that distribution and then apply the micro‐ensemble procedure to each. For our Dirac initial‐condition tests we set $M_{\mathrm{macro}}=1$.
Training vs. evaluation.
During training the model only ever sees independent pairs $(u_{0},u_{1})$—one target per input—and never observes any ensemble. The above ensemble perturbation is used only to construct a ground‐truth distribution at test time.
Appendix L Models and baselines
We tested our method against $4$ other baselines, out of which $2$ (namely GenCFD and its FNO-conditioned variant) are diffusion-based algorithms of the same kind. The other two baselines rely on traditional operator learning setup, with the UViT baseline satisfying also the purpose of an ablation study that quantifies the gain obtained by adding diffusion on top.
Shared Backbone. We stress that for GenCFD, GenCFD $\circ$ FNO and UViT we use the exact same UViT backbone architecture outlined in the earlier parameterization of our rectified flow as in Table 5 for the sake of consistency. The architecture is further illustrated schematically in Figure 5, which shows the UNet-based design in detail.
GenCFD
GenCFD Molinaro et al. (2024) is an end-to-end conditional score‐based diffusion model that directly learns the mapping from an input flow field $u_{0}$ to the target field $u_{1}$. It uses the same UViT backbone described in Table 5, with noise levels $\sigma_{\tau}$ corresponding to the standard deviation of the added Gaussian perturbation, in contrast to our ReFlow model which evolves over a rectified diffusion time $\tau\in[0,1]$. Common choices for noise schedulers include exponential and tangent formulations. In our case, we consistently adopted an exponential noise schedule, as it yielded the best results across all datasets and metrics reported. For score-based diffusion models, either variance-exploding (VE) or variance-preserving (VP) formulations are typically used. In all GenCFD models, we opted for the VE setting throughout. During training we minimize the denoising score‐matching loss over $(u_{0},u_{1})$ pairs sampled uniformly in $\tau\in[0,1]$, using Adam with initial learning rate $3\times 10^{-4}$ (cosine decay), batch size 16, and EMA decay 0.9999. As is common in score-based diffusion models, we applied denoiser preconditioning and used uniform weighting across noise levels. At inference we solve the corresponding SDE (via an Euler-Maruyama scheme with discretization $128$ steps) and incorporate reconstruction guidance to condition on $u_{0}$ so that we recover the conditional posterior measure $p(u_{1}|u_{0})$.
[…]
In this hybrid approach, a Fourier Neural Operator (FNO) Li et al. (2020a) is first trained under an $\ell_{2}$ loss to predict $u_{1}$ from $u_{0}$. Its output $\hat{u}_{1}^{\rm FNO}$ is concatenated channel‐wise with $u_{0}$ and fed into the GenCFD score network at both train and test time. All other architectural and optimization settings (backbone, scheduler, SDE sampler) remain identical to GenCFD, allowing the FNO to provide fast low‐frequency priors while the diffusion stage refines fine‐scale turbulent features.
Custom implementation.This schematic closely parallels the pipeline of Oommen et al. (2024), but with two key distinctions. First, we fully replace both the diffusion backbone and its denoising stages with GenCFD. Second, and most important, we condition not only on the FNO’s coarse solution but also on the original high-resolution initial condition. This ensures that our model retains fine-scale structures that would otherwise be lost if we relied solely on the FNO predictions.
FNO
Our FNO baseline follows Li et al. Li et al. (2020a) and is implemented in PyTorch. First, each spatial location of the input field $u_{0}$ is lifted from $C_{\rm in}$ to 256 channels via a two‐layer MLP with a SiLU activation, then projected down to a 128‐channel hidden representation. This is followed by four FnoResBlocks: in each block, the hidden features undergo a Conditional LayerNorm (time embedding size 128), SiLU, and two spectral convolutions (SpectralConv) with num_modes=(24,24)num\_modes=(24,24), combined through a learnable “soft‐gate” residual skip. Finally, a two‐layer projection MLP (128 → 256 → $C_{\rm out}$, with SiLU in between) produces the output field $\hat{u}_{1}$. We train for 100k iterations minimizing $\|u_{1}-\hat{u}_{1}\|_{2}^{2}$ with Adam (learning rate $10^{-3}$, batch size 16), and perform inference in one forward pass—no diffusion or autoregression is used.
UViT
Our UViT surrogate Saharia et al. (2022) uses the identical UNet2D backbone of Table 5, injecting time embeddings and employing 8‐head multi‐head attention at the bottleneck. It is trained deterministically with MSE loss on $(u_{0},u_{1})$, using Adam at $3\times 10^{-4}$, batch size 16, for 100 k iterations (EMA 0.9999). For multi‐step forecasting we roll out autoregressively, feeding each prediction back as the next input. This ablates out the diffusion component while matching all other design choices.
Baseline Model Summary and Implementation Setup
Table 6 summarizes key architectural settings for all baseline models considered in this work. These configurations were selected to reflect the strongest performance achievable within a shared training budget of up to 120,000 iterations, as used in the quantitative comparison reported in Table  1.
Padding and boundary conditions.
Each dataset required padding choices tailored to its physical boundary conditions. SL and RM, which exhibit periodic boundaries, use circular padding in all convolutional layers. This significantly improved the performance of the diffusion-based models. In contrast, CS features non-periodic boundaries and is thus trained with standard zero padding. These padding choices apply to all UNet-based architectures except ReFlow, which uniformly uses zero padding across all datasets regardless of boundary conditions.
Table 6: Summary of architectural depth and approximated parameter count for each model evaluated. All UNet-based models (ReFlow, GenCFD, GenCFD $\circ$ FNO, and UViT) share a consistent backbone design to ensure consistency across diffusion and non-diffusion baselines, while the FNO baseline uses a structurally different architecture. These configurations reflect the best-performing settings under a training budget of up to 120,000 iterations, as used in the statistical comparisons reported in Table 1.
Model # Levels # Blocks # Params Noise Schedule Diffusion Scheme
ReFlow 3 4 22M uniform rectified continuous
time variable $t\in[0,1]$
GenCFD 2 4 5M exponential variance-exploding
GenCFD $\circ$ FNO 2 4 5M exponential variance-exploding
UViT 2 4 5M
FNO 4 11M
Appendix M Results
M.1 Performance Metrics
To compare our generated ensembles against the reference Monte Carlo samples, we employ three complementary measures:
  • Mean‐field error. Let $\mu_{\rm ref}(x)$ and $\mu_{\rm model}(x)$ denote the pointwise spatial means of the reference and model ensembles, respectively. We measure their discrepancy by the $L^{2}$‐norm
    $e_{\mu}\;=\;\bigl\|\,\mu_{\rm ref}\;-\;\mu_{\rm model}\bigr\|_{2}\,.$
  • Standard‐deviation error. Similarly, let $\sigma_{\rm ref}(x)$ and $\sigma_{\rm model}(x)$ be the pointwise standard deviations. We therefore compute
    $e_{\sigma}\;=\;\bigl\|\,\sigma_{\rm ref}\;-\;\sigma_{\rm model}\bigr\|_{2}\,$,
    Both metrics are reported after normalization by the $L^{2}$ norm of the ground truth.
  • Average 1-Wasserstein distance. For a fixed initial condition $u_{0}$, let $\{u^{(j)}_{\rm ref}(x)\}_{j=1}^{N}$ and $\{u^{(j)}_{\rm model}(x)\}_{j=1}^{N}$ be the ensembles of final‐time values produced by the reference solver and our model, respectively. Since these fields are discretized on $M$ spatial points $\{x_{i}\}_{i=1}^{M}$, we obtain two empirical 1D distributions at each $x_{i}$. Denoting their inverse CDFs by $F^{-1}_{\rm ref}(\cdot\,;\,x_{i})$ and $F^{-1}_{\rm model}(\cdot\,;\,x_{i})$, the pointwise Wasserstein‐1 distance is
    $W_{1}\bigl(p_{\rm ref}(\cdot\mid x_{i}),\,p_{\rm model}(\cdot\mid x_{i})\bigr)\;=\;\int_{0}^{1}\bigl|\,F^{-1}_{\rm ref}(q;x_{i})\;-\;F^{-1}_{\rm model}(q;x_{i})\bigr|\,\mathrm{d}q.$
    We then report the spatial average
    $\overline{W}_{1}\;=\;\frac{1}{M}\sum_{i=1}^{M}W_{1}\bigl(p_{\rm ref}(\cdot\mid x_{i}),\,p_{\rm model}(\cdot\mid x_{i})\bigr).$
Energy spectra. Finally, to assess how well different scales are captured, we compute the discrete energy spectrum of each velocity field. Given a sample $u(x)$ on a dd‐dimensional periodic grid with spacing $\Delta$, let $\widehat{u}_{k}$ be its Fourier coefficient at integer wavevector $k\in\mathbb{Z}^{d}$. We bin modes by their $\ell_{1}$ radius $|k|_{1}=k_{1}+\dots+k_{d}$, defining
$E_{r}\;=\;\frac{\Delta^{d}}{2}\,\sum_{\begin{subarray}{c}k\in\mathbb{Z}^{d}\\ |k|_{1}=r\end{subarray}}\bigl\|\widehat{u}_{k}\bigr\|^{2},\quad r=0,1,2,\dots$
and compare the ensemble‐average spectra of reference and model. Our plots show $E_{r}$ versus $r$ on log–log axes to highlight scale‐by‐scale agreement.
M.2 Training Efficiency Observations
To assess training efficiency, we compare ReFlow with GenCFD. For fairness, we use a GenCFD variant with the same setup and parameter count (22M), matching the three-level UViT architecture of ReFlow instead of its standard lightweight configuration (5M). Both models were trained for up to 160,000 iterations with a batch size of 16.
Figure 6 summarizes performance over the training trajectory. At approximately 100,000 steps, ReFlow consistently outperforms GenCFD across all key metrics: relative $L^{2}$ error in the mean ($e_{\mu}$), relative $L^{2}$ error in the standard deviation ($e_{\sigma}$), and the average pointwise Wasserstein-1 (W1) distance—evaluated for both $u_{x}$ and $u_{y}$ velocity components. Notably, the $y$-axis uses a logarithmic scale to emphasize differences in error magnitude across training steps.
Across all metrics, ReFlow begins to outperform GenCFD at around 100,000 training steps and continues to improve at a sharper rate. While neither model reaches full convergence within 160,000 iterations, ReFlow consistently exhibits lower errors and more favorable metric trajectories in later stages of training. These results support the observation that ReFlow is more training-efficient within a fixed compute budget and ultimately achieves stronger generalization performance in this setting.
To further probe the long-term capabilities of GenCFD, we extended its training to 600,000 iterations and checked whether it eventually surpasses ReFlow’s performance at its final recorded checkpoint (150,000 steps). Table 7 summarizes the earliest training steps at which GenCFD overtakes ReFlow for each evaluation metric. While GenCFD does catch up and exceed ReFlow in several cases, it never surpasses ReFlow on the Wasserstein-1 distance for $u_{y}$ within the training horizon considered.
Table 7: Number of training iterations at which GenCFD surpasses ReFlow’s performance at 150,000 iterations. If not surpassed, indicated by “—”.
Metric GenCFD Iteration GenCFD Value ReFlow Value @ 150,000 iter.
$u_{x}$ Mean Error ($e_{\mu}$) 280,000 3.33 $\times 10^{-2}$ 3.45 $\times 10^{-2}$
$u_{y}$ Mean Error ($e_{\mu}$) 400,000 1.70 $\times 10^{-1}$ 1.88 $\times 10^{-1}$
$u_{x}$ Std Error ($e_{\sigma}$) 400,000 6.48 $\times 10^{-2}$ 7.24 $\times 10^{-2}$
$u_{y}$ Std Error ($e_{\sigma}$) 340,000 7.91 $\times 10^{-2}$ 7.97 $\times 10^{-2}$
$u_{x}$ Wasserstein-1 ($W_{1}$) 200,000 3.89 $\times 10^{-2}$ 3.95 $\times 10^{-2}$
$u_{y}$ Wasserstein-1 ($W_{1}$) 4.11 $\times 10^{-2}$
Figure 6: Training efficiency comparison. Performance of ReFlow and GenCFD on the Macro-Micro SL ensemble dataset evaluated at multiple training checkpoints. Metrics match those used in the main text: relative $L^{2}$ error of the mean ($e_{\mu}$) and standard deviation ($e_{\sigma}$), and the average 1-point Wasserstein-1 distance (W1), computed for both velocity components ($u_{x}$, $u_{y}$). The $y$-axis uses a logarithmic scale, while the $x$-axis (training iteration) is linear.
M.3 Experimental Advantages of Straightness
This subsection presents empirical evidence that the “rectified” trajectories generated by our Rectified Flow model (ReFlow) enable faster convergence and yield higher-fidelity reconstructions compared to the baseline GenCFD diffusion approach.
Evaluation settings.
We use the CloudShock dataset throughout for exposition purposes. However, we stress that the same results hold across datasets. For each initial condition, both ReFlow and GenCFD are allotted $T=10$ diffusion steps over normalized diffusion time $\tau\in[0,1]$. We record the model outputs at five evenly spaced timesteps $t=\{0.0,0.25,0.50,0.75,1.0\}$ and analyze:
  • Trajectory inpainting of the density field $\rho$.
  • Latent‐Space Trajectory Visualization via PCA via PCA.
  • Per‐Sample Average Error Evolution: average MSE across physical channels.
  • Evolution of the Energy Spectrum: radial power‐spectrum of energy $E$ as inpaintig progresses.
All other components (backbone, noise schedule, batch size) are held equal for fair comparison.
Trajectory Inpainting
We compare Rectified Flow (ReFlow) and GenCFD reconstructions of the density field $\rho$ over five normalized diffusion timesteps $t\in\{0.00,0.25,0.50,0.75,1.00\}$. ReFlow inpaints fine-scale features in far fewer steps, while GenCFD remains overly diffusive. The leftmost panel corresponds to the initial condition $u_{0}$ upon which both model are conditioned, while the following panels showcase the evolution of the model’s output as diffusion time progresses.
Figure 7: Density trajectories for Sample 1: ReFlow (top) vs. GenCFD (bottom). ReFlow already recovers sharp shock fronts by diffusion time $t=1.0$.
Figure 8: Density trajectories for Sample 2.
Figure 9: Density trajectories for Sample 3.
Latent‐Space Trajectory Visualization via PCA
To quantify how directly each model moves through the data manifold, we flatten the $(H\times W\times C)$ field at each timestep into a single feature vector, stack all Rectified Flow and GenCFD vectors, fit a 3-component PCA, and then project each timestep into PC1/PC2/PC3. A straighter path indicates fewer diffusive detours.
(a) Sample 1
(b) Sample 2
(c) Sample 3
(d) Sample 4
(e) Sample 5
(f) Sample 6
Figure 10: 3D PCA latent-space trajectories (Samples 1–6). We project the flattened snapshots at each of $T$ timesteps into the top three PCA components. Rectified Flow (blue) consistently traces a straighter path than GenCFD (red), confirming its “rectified” evolution. Circles mark $t=0$, squares mark $t=1$, and intermediate timesteps are annotated.
To further illustrate the nature of each model’s generative dynamics, we visualize single-sample generation trajectories from both models in a 2D PCA space for the RM dataset (Figure 11). For each physical channel ($\rho$, $m_{x}$, $m_{y}$, $p$), we apply PCA to the full set of ground truth snapshots to define a common projection basis. We then overlay the generative trajectory of a single sample in this space. To contextualize the generated dynamics, we add kernel density estimates (KDEs) showing the empirical data distribution in the same projected space.
ReFlow (top) again traces a straighter, more compact path over just 10 rectified steps, moving coherently along the data manifold. GenCFD (bottom), by contrast, uses 128 stochastic denoising steps and produces a more diffuse and nonlinear trajectory, often deviating from high-density regions of the ground truth distribution. While the PCA projection inherently reduces dimensional complexity, the consistent difference in shape and alignment supports the conclusion that ReFlow’s rectified flow better preserves manifold structure and sample efficiency.
Figure 11: 2D PCA latent-space trajectories. We project the flattened velocity snapshots from the RM dataset into a 2D PCA space, computed independently for each channel ($\rho$, $m_{x}$, $m_{y}$, $p$) across all $T$ time steps. To visualize the underlying distribution of target fields, we apply kernel density estimation (KDE) per channel. The ReFlow model (top) evolves over 10 rectified time steps, producing visibly straighter and more structured trajectories compared to the score-based diffusion model GenCFD (bottom), which uses 128 stochastic steps. Although the projection into two principal components necessarily truncates the high-dimensional dynamics, the qualitative difference remains evident, illustrating the smoother and more directed progression learned by ReFlow.
Per‐Sample Average Error Evolution
We compute the mean‐squared error at each timestep $t$ and average across all $C$ channels:
$\overline{\mathrm{MSE}}(t)=\frac{1}{C}\sum_{c=1}^{C}\mathrm{MSE}_{c}(t).$
Figure 12 shows, for Samples 1–6, the log‐scaled average MSE (left) and the normalized error $\overline{\mathrm{MSE}}(t)/\overline{\mathrm{MSE}}(0)$ (right). Rectified Flow (blue) consistently reduces error faster and to a lower residual than GenCFD (red).
(a) Sample 1
(b) Sample 2
(c) Sample 3
(d) Sample 4
(e) Sample 5
(f) Sample 6
Figure 12: Per‐sample average MSE evolution. Each subfigure plots, left half: $\log_{10}\overline{\mathrm{MSE}}(t)$; right half: $\overline{\mathrm{MSE}}(t)/\overline{\mathrm{MSE}}(0)$. Rectified Flow (blue) rapidly drives error down and keeps it orders of magnitude below GenCFD (red) across all samples.
Evolution of the Energy Spectrum
We examine the 2D radial power spectrum of the energy field $E$ at five normalized diffusion times $\tau\in\{0.00,0.25,0.50,0.75,1.00\}$. Each figure below shows log–log power vs. wavenumber for Rectified Flow (blue), GenCFD (red), and the ground-truth target (black dashed), with inset log-MSE error annotations. The leftmost snapshot corresponds to the spectrum of the initial data $u_{0}$ upon which the model is conditioned. Subsequent panels track the spectral evolution of the model predictions, as diffusion time progresses.
Figure 13: Energy spectrum evolution for Sample 1. ReFlow (blue) outperforms GenCFD (red) in capturing high-wavenumber energy.
Figure 14: Energy spectrum evolution for Sample 2. ReFlow (blue) remains closer to the ground truth (dashed) than GenCFD (red).
Figure 15: Energy spectrum evolution for Sample 3. The inset log-MSE annotations quantify ReFlow’s advantage at each $t$.
Figure 16: Energy spectrum evolution for Sample 4. ReFlow’s spectrum (blue) gets closer to the target (dashed) as $\tau\to 1.0$.
Figure 17: Energy spectrum evolution for Sample 5. ReFlow’s superior high-$k$ fidelity is evident throughout.
Figure 18: Energy spectrum evolution for Sample 6. ReFlow better captures the ground‐truth spectrum’s slope across all timesteps.
Appendix N Computational Considerations
All of our ReFlow models for 2D fluid-flow tasks use the same network configuration summarized in Table 5 (channels $(128,256,256)$, downsample ratios $(2,2,2)$, 8 attention heads, dropout 0.1, batch size 16, etc.). Consequently, every model—regardless of dataset—has the same total parameter count (approximately $8\text{–}10$M parameters) and identical layer‐wise dimensions. We ran each training job for 100 000 iterations on a single NVIDIA Quadro RTX 6000 (Driver 570.124.06, CUDA 12.8) in persistence mode. Resource utilization averaged as follows over the full 100 h window:
Table 8: Aggregate hardware utilization for all “usual‐setup” runs.
Metric Range Typical
GPU power draw (W) 200–260 230
GPU utilization (%) 30–80 55
GPU memory allocated (GB) 12–17 15
GPU temperature (°C) 40–60 50
Process GPU memory usage (%) 60–80 70
Process GPU memory access time (%) 10–50 30
Process CPU threads in use 7 7
System CPU utilization (per core, %) 60–100 80
Process memory in use (GB) up to 50 25
Disk I/O read (MB total) up to $1.2\times 10^{6}$ $5\times 10^{5}$
Disk utilization (%) 70–75 72
Network traffic (bytes total) $\approx 3\times 10^{13}$ $3\times 10^{13}$
Since all datasets are paired with models of identical size and the hardware envelope (GPU power, memory, temperature, CPU threads, I/O) is effectively the same across runs, any observed performance differences are attributable to the sampling strategy (noise schedule, ODE steps) rather than to model‐size or hardware variations (the network traffic is due to Wandb usage).
All experiments were run on a single NVIDIA Quadro RTX 6000 GPU (Driver 570.124.06, CUDA 12.8), operating in persistence mode.
Each model was trained for 100 000 iterations. We explored architectures ranging from approximately 5 million to 10 million trainable parameters; across this range we observed no significant differences i n final predictive performance or convergence behavior. All other hyperparameters (learning rate schedules, batch sizes, noise schedules, etc.) were held fixed when evaluated against each other, ensuring that differences in accuracy and runtime reflect only the conditioning and sampling strategies under study.
Appendix O Code and Data Availability
Prior to acceptance, we will make available all three datasets, the complete data-processing scripts, and a fully functional training and evaluation pipeline for ReFlow, including a provided model checkpoint that can be loaded and tested. A more comprehensive, user-friendly code release will follow shortly thereafter.
Appendix P Additional Visualizations
Test‐time ensemble perturbation (evaluation only).
Following the protocol of GenCFD Molinaro et al. (2024), we construct small “micro‐ensembles” around each test initial condition $\bar{u}$ by sampling $M_{\mathrm{micro}}$ perturbed fields uniformly in an $\varepsilon$–ball centered at $\bar{u}$. Each perturbed member is then advanced to the evaluation time $t$ with a high‐fidelity PDE solver to produce a reference ensemble $\{u^{(j)}_{\rm ref}(t)\}_{j=1}^{M_{\mathrm{micro}}}$. Our model generates a corresponding ensemble $\{u^{(j)}_{\rm model}(t)\}_{j=1}^{M_{\mathrm{micro}}}$ from the same perturbed inputs, but without ever observing ensembles during training (which only uses independent pairs $(u_{0},u_{1})$). We compare these two ensembles via mean‐field error $e_{\mu}$, standard‐deviation error $e_{\sigma}$, and spatially averaged 1‐Wasserstein distance $\overline{W}_{1}$ (see SM §M).
Out‐of‐distribution testing (cloud–shock and shear‐flow). Because the micro‐perturbations render each initial field slightly OOD, this evaluation probes the model’s ability to generalize under small chaotic perturbations. In these tasks, our Rectified Flow ensemble closely tracks the reference uncertainty and multiscale statistics, whereas deterministic baselines collapse to a single trajectory and fail to capture the conditional variability.
In‐distribution scaling (Richtmyer–Meshkov). For the Richtmyer–Meshkov dataset, we additionally study the effect of training set size using a held‐out test set from the same distribution. Here, even deterministic approaches improve with more data, but still underperform on the micro‐ensemble test: they require significantly larger training sets to match ReFlow’s accuracy and exhibit poor calibration under small perturbations.
In this appendix we provide a comprehensive set of qualitative comparisons between the baselines (GenCFD variants, UViT, FNO) models and our Rectified Flow (ReFlow) approach across three representative 2D fluid‐flow benchmarks (Richtmyer–Meshkov, shear layer roll‐up, and cloud-shock interaction). For each task we show:
  • Mean fields Ensemble averages reveal how each deterministic model diffuses fine‐scale features (UVit/FNO) versus perserving sharp interfaces (ReFlow, GenCFD variants).
  • Uncertainty maps Per‐pixel standard deviations illustrate whether uncertainty localizes in physically meaningful regions (e.g. shocks, shear interfaces) or remains overly diffuse.
[…]
  • Spectral & pointwise comparisons (Fig. 22: Richtmyer–Meshkov global $E(k)$, local spectra, and pointwise density histograms; Fig. 25: Shear‐layer global $u_{x}$ spectra and pointwise $u_{x}$ distributions; Fig. 29: Cloud–shock global $m_{y}$ spectra and pointwise energy distributions) — combined Fourier‐space and histogram panels quantify multiscale fidelity and calibration of the predictive ensembles.
Together, these visualizations underscore ReFlow’s ability to maintain sharper physical structures, produce better‐calibrated uncertainties, and faithfully reproduce multiscale spectral statistics.
Figure 19: Richtmyer–Meshkov Mean Field Comparison: Mean velocity fields for the Richtmyer–Meshkov dataset, shown across all physical channels and models. In many cases, predictions closely resemble the ground truth, with differences often imperceptible, underscoring the overall high quality and consistency of the generated mean fields.
Figure 20: Richtmyer–Meshkov Uncertainty Comparison: Per-pixel standard deviation fields from the Richtmyer–Meshkov dataset across all models. Stochastic generative models capture spatially coherent uncertainty patterns, outperforming the deterministic baselines (FNO and UViT), which exhibit limited variability and poor uncertainty localization.
[…]
Figure 22: Richtmyer–Meshkov: spectral and pointwise distribution comparisons between UVit and Rectified Flow.
Figure 23: Shear‐Layer Mean and Standard Deviation Comparison: Visualization of mean-field ($\mu$) and standard-deviation ($\sigma$) predictions for the Shear‐Layer dataset across both velocity components ($u_{x}$, $u_{y}$). For each model, the left half shows the predicted mean fields and the right half shows per-pixel standard deviations. The top row contains the reference ground truth. Despite the reduced channel complexity of this dataset, clear differences emerge in how well models capture coherent structures and localized uncertainties. ReFlow preserves sharper features and more accurately reflects the spatial distribution of uncertainty.
[…]
Figure 25: Shear‐layer: global $u_{x}$ spectra and pointwise $u_{x}$ distributions.
Figure 26: Cloud-Shock Mean Field Comparison: Mean velocity fields for the Cloud-Shock dataset, shown across all physical channels and models. In most cases, predictions closely resemble the ground truth.
Figure 27: Cloud–Shock Uncertainty Comparison: Per-pixel standard deviation visualizations for the Cloud–Shock dataset across all models. While all methods capture key uncertainty patterns, stochastic models produce more localized and physically meaningful variability.
[…]
Figure 29: Cloud‐shock: spectral and pointwise energy comparisons.